### v3

# Characteristic Learning for Provable One Step Generation

Abstract.
We propose the characteristic generator, a novel one-step generative model that combines the efficiency of sampling in Generative Adversarial Networks (GANs) with the stable performance of flow-based models. Our model is driven by characteristics, along which the probability density transport can be described by ordinary differential equations (ODEs). Specifically, We estimate the velocity field through nonparametric regression and utilize Euler method to solve the probability flow ODE, generating a series of discrete approximations to the characteristics. We then use a deep neural network to fit these characteristics, ensuring a one-step mapping that effectively pushes the prior distribution towards the target distribution. In the theoretical aspect, we analyze the errors in velocity matching, Euler discretization, and characteristic fitting to establish a non-asymptotic convergence rate for the characteristic generator in $2$-Wasserstein distance. To the best of our knowledge, this is the first thorough analysis for simulation-free one step generative models. Additionally, our analysis refines the error analysis of flow-based generative models in prior works. We apply our method on both synthetic and real datasets, and the results demonstrate that the characteristic generator achieves high generation quality with just a single evaluation of neural network.
Key words and phrases:
flow-based generative models, one-step generation, convergence rate analysis.
* Corresponding author
1. Introduction
Generative models aim to learn and sample from an underlying target distribution, finding applications in diverse fields such as image and video generation (Radford et al., 2016; Meng et al., 2022; Ho et al., 2022), text-to-image generation (Ramesh et al., 2021, 2022; Kang et al., 2023), and speech synthesis (Kong et al., 2021; Chen et al., 2021). One of the most influential and widely-used approaches is GAN (Goodfellow et al., 2014) and its variants (Arjovsky et al., 2017). GANs offer the advantage of high sampling efficiency, as generating new samples merely entails a single evaluation of the trained generator. Despite the remarkable success in practical applications (Reed et al., 2016) and theoretical guarantee (Liang, 2021; Liu et al., 2021; Huang et al., 2022; Zhou et al., 2023), GANs have intrinsic limitations in terms of their stability (Salimans et al., 2016).
In recent years, diffusion models (Ho et al., 2020; Song et al., 2021b, c; Karras et al., 2022) and flow-based models (Liu et al., 2022; Lipman et al., 2023) have emerged as powerful generative models. These models have also laid the foundation for the development of generative AI models, such as DALL-E (Ramesh et al., 2021, 2022), Midjourney, Stable Diffusion (Esser et al., 2024), and Sora (Brooks et al., 2024). Theoretical analysis for these methods has been studied by Oko et al. (2023); Lee et al. (2022, 2023); Chen et al. (2023d, c); Benton et al. (2024a, b); Gao and Zhu (2024); Wu et al. (2024). Although diffusion or flow-based models outperform GANs in generation quality across various tasks (Dhariwal and Nichol, 2021), they require hundreds or even thousands of sequential steps involving large neural network evaluations for sampling. As a consequence, their sampling speed is much slower compared to one-step GANs.
The instability of GANs and the inefficiency of sampling in diffusion or flow-based models have emerged as significant bottlenecks in practical applications of generative models. This raises two crucial questions:
Is it possible to develop a one-step generative model that combines the efficient sampling of GANs with the stable performance of diffusion or flow-based models? If so, can we establish a rigorous error analysis for this generative model?
Several recent papers have made progress in addressing the first question using various techniques such as distillation (Luhman and Luhman, 2021; Salimans and Ho, 2022; Song et al., 2023; Zhou et al., 2024), operator learning (Zheng et al., 2023a), or trajectory models (Kim et al., 2024; Ren et al., 2024). For a more detailed discussion, please refer to Section 5.1. Despite these recent advancements, a unifying mathematical framework for designing and analyzing the one-step generative models remains largely limited (Li et al., 2024b). This paper aims to fill this gap and provide a positive answer to the aforementioned questions. Specifically, we introduce a comprehensive framework, known as the characteristic generator, aiming to streamline the sampling process using a single evaluation of the neural network. Our model merges the sampling efficiency of GANs with the promising performance of flow-based models. Furthermore, we present a rigorous error analysis for the characteristic generator. Through numerical experiments, we validate that our approach generates high-quality samples that is comparable to those generated through flow-based models, all while requiring just a single evaluation of characteristic generator.
1.1. Contributions
Our contributions are summarized as follows:
  1. (i)
    We introduce a one-step generative model called the characteristic generator, which efficiently pushes a Gaussian distribution to the target distribution without the need for simulation. We present a probability density transport equation via stochastic interpolants, resulting in probability flow ODEs along characteristics. The velocity field of the ODE is estimated through a least-squares regression. Subsequently, we generate a sequence of discrete approximations to characteristics by numerically solving the ODE. Finally, we train the characteristic generator by fitting these characteristics using a deep neural network.
  2. (ii)
    We provide a rigorous error analysis for the characteristic generator. Specifically, we derive a convergence rate ${\mathcal{O}}(n^{-\frac{1}{d+3}})$ for velocity matching (Theorem 3.9), which improves the rates in previous works such as Chen et al. (2023b). Additionally, we propose an error bound for the $2$-Wasserstein distance between the data generated by the Euler method and the target distribution (Theorem 3.11), which is of independent interest. Lastly, we present a non-asymptotic convergence rate for the characteristic generator in the $2$-Wasserstein distance, demonstrating that it achieves a rate same to that of Euler sampling (Theorem 3.13). To the best of our knowledge, this is the first analysis conducted for one-step sampling of flow-based models, providing valuable theoretical insights for distillation (Salimans and Ho, 2022; Song et al., 2023), operator learning (Zheng et al., 2023a), or trajectory model (Kim et al., 2024; Ren et al., 2024).
  3. (iii)
    We validate the generation quality and sampling efficiency of the characteristic generator using both synthetic and real data through numerical experiments. The experiment results demonstrate that our method can generate samples of high quality from noise with just one neural network evaluation. In comparison to previous one-step generative models without the aid of GANs (Salimans and Ho, 2022; Song et al., 2023; Kim et al., 2024) , our characteristic generator significantly improves the generation quality in CIFAR-10, as shown in Table 4. Furthermore, with only a few iterations, our characteristic generator achieves comparable generation performance to the state-of-the-art method (Kim et al., 2024, CTM), without the need for additional GAN training as employed by CTM.
1.2. Main Results
Let $\mu_{1}\in P_{\operatorname{ac}}({\mathbb{R}}^{d})$ be a target probability distribution with $\,\mathrm{d}\mu_{1}(x)=\rho_{1}(x)\,\mathrm{d}x$. In the context of generative learning, one only has access to data samples from $\mu_{1}$ but have no access to the density function $\rho_{1}$ itself. Let $\mu_{0}\in P_{\operatorname{ac}}({\mathbb{R}}^{d})$ be a known prior distribution. Suppose that there exists a smooth mapping $G^{*}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$, which pushes-forward the prior distribution $\mu_{0}$ onto the target distribution $\mu_{1}$, that is,
(1.1) $G_{\sharp}^{*}\mu_{0}=\mu_{1}.$
The equation (1.1) is known as the normalizing equation (Rozen et al., 2021). The goal of the generative learning, at least formally, is to find an estimator $\widehat{G}$ of the push-forward operator $G^{*}$ based on finite samples drawn from $\mu_{1}$.
In this work, we aim to construct the desired push-forward operator via a probability flow ODE
$\,\mathrm{d}x(t)=b^{*}(t,x(t))\,\mathrm{d}t,\quad x(0)=x_{0}\sim\mu_{0},$
where the velocity field $b^{*}$ is given as (2.3). Denote by $\mu_{t}$ the distribution of $x(t)$ for each $t\in(0,1)$, and define the probability flow $g_{t,s}^{*}$ as $g_{t,s}^{*}(x(t))=x(s)$ for each $10\leq t\leq s<$ 1. Let $\widehat{b}$ be the estimated velocity field obtained by velocity matching (2.8), and let $\widehat{E}_{0,K}^{\tau}$ be the numerical approximation to the ODE solution by Euler method (2.10). Denote by $\widehat{g}_{s,t}$ be the estimation of the flow $g_{t,s}^{*}$ defined as (2.12), which is referred to the characteristic generator.
Our theoretical results are established under the following assumptions on the prior and target distributions, which will be discussed in Section 3.1.
Assumption 1.
The prior distribution $\mu_{0}=N(0,I_{d})$.
Assumption 2.
The target distribution $\mu_{1}=N(0,\sigma^{2}I_{d})*\nu$, where $\nu$ is a distribution with $\operatorname{supp}(\nu)\subseteq[0,1]^{d}$.
Assumption 2 requires the target distribution to be smoothed by a Gaussian distribution. This assumption is essential as it ensures desirable properties of the probability flow ODE, such as bounded moments and the Lipschitz property of the velocity field. It is noteworthy that this assumption can be considered relatively mild, given that the smoothed distribution $\mu_{1}$ is an approximation of the original distribution $\nu$, particularly when the variance $\sigma^{2}$ of the Gaussian distribution is small. Further details can be found in Sections 3.1 and 3.2.
Our first main result gives an error bound for the velocity matching.
Theorem 1.1 (Informal version of Theorem 3.9).
Suppose that Assumptions 1 and 2 hold. Let ${\mathcal{S}}$ be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$. Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\big{[}\|b^{*}(t,X_{t})-\widehat{b}(t,X_{t})\|_{2}^{2}\big{]}\,\mathrm{d}t\Big{]}\leq C_{T}n^{-\frac{2}{d+3}}\log^{2}n,$
where $C_{T}$ is a constant depending on $d$, $\sigma$ and $T$.
We then present the $2$-Wasserstein bound for generated data by Euler method.
Theorem 1.2 (Informal version of Corollary 3.12).
Suppose that Assumptions 1 and 2 hold. Let ${\mathcal{S}}$ be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$. By setting a sufficiently large number of time steps $K$ for Euler method, the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}\Big{(}(\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big{)}\Big{]}\leq C_{T}^{1}n^{-\frac{2}{d+3}}\log^{2}n+C_{T}^{2}W_{2}^{2}\big{(}\mu_{0},\mu_{1}\big{)},$
where $C_{T}^{1}$ and $C_{T}^{2}$ are two constants depending on $d$, $\sigma$ and $T$. Further, as the stopping time $T\rightarrow 1$, the constant $C_{T}^{1}$ tends to infinity while $C_{T}^{2}$ decreases to zero.
The averaged $2$-Wasserstein bound for characteristic generator is stated as follows.
Theorem 1.3 (Informal version of Theorem 3.13).
Suppose that Assumptions 1 and 2 hold. Let ${\mathcal{S}}$ be a set of $n$ samples independently and identically drawn from the target distribution $\mu_{1}$, and let ${\mathcal{Z}}$ be the set of numerical solutions by Euler method with a sufficiently large number. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}W_{2}^{2}\Big{(}(\widehat{g}_{t,s})_{\sharp}\mu_{t},\mu_{s}\Big{)}\,\mathrm{d}s\mathrm{d}t\Big{]}\leq C_{T}n^{-\frac{2}{d+3}}\log^{2}n,$
where $C$ is a constant depending on $d$, $\sigma$ and $T$.
As a consequence of Theorem 3.13, we propose in Table 1 the convergence rate of characteristic generators induced by two special probability flow ODEs.
Table 1. Convergence rates of characteristic generator.
Probability flow ODE Convergence rate
Linear interpolants ${\mathcal{O}}(n^{-\frac{2}{3(d+3)}})$ Corollary 3.14
Föllmer flow ${\mathcal{O}}(n^{-\frac{2}{5(d+3)}})$ Corollary 3.15
Experiment results and discussions can be found in Section 4. Our code is online available at https://github.com/burning489/CharacteristicGenrator.
1.3. Preliminaries and Notations
1.3.1. Wasserstein Distance
Let $P_{\operatorname{ac}}({\mathbb{R}}^{d})$ be the space of probability measures on ${\mathbb{R}}^{d}$, which are absolutely continuous with respect to Lebesgue measure. Suppose $\mu_{0},\mu_{1}\in P_{\operatorname{ac}}({\mathbb{R}}^{d})$ with $\,\mathrm{d}\mu_{0}(x)=\rho_{0}(x)\,\mathrm{d}x$ and $\,\mathrm{d}\mu_{1}(x)=\rho_{1}(x)\,\mathrm{d}x$. The 2-Wasserstein distance (Villani, 2009, Definition 6.1) between $\mu_{0}$ and $\mu_{1}$ is defined by the formula
(1.2) $W_{2}(\mu_{0},\mu_{1})=\inf\Big{\{}{\mathbb{E}}^{1/2}\big{[}\|X_{0}-X_{1}\|_{2}^{2}\big{]}:\leavevmode\nobreak\ \operatorname{law}(X_{0})=\mu_{0},\leavevmode\nobreak\ \operatorname{law}(X_{1})=\mu_{1}\Big{\}}.$
The 2-Wasserstein distance satisfies the symmetry axiom and the triangle inequality. Further, the distance $W_{2}(\mu_{0},\mu_{1})$ is equal to zero if and only if $\mu_{0}=\mu_{1}$.
1.3.2. Deep Neural Networks
A neural network $f:{\mathbb{R}}^{N_{0}}\rightarrow{\mathbb{R}}^{N_{L+1}}$ is a function defined by
$f(x)=T_{L}(\varrho(T_{L-1}(\cdots\varrho(T_{0}(x))\cdots))),$
where the activation function $\varrho$ is applied component-wisely and $T_{\ell}(x):=A_{\ell}x+b_{\ell}$ is an affine transformation with $A_{\ell}\in{\mathbb{R}}^{N_{\ell+1}\times N_{\ell}}$ and $b_{\ell}\in{\mathbb{R}}^{N_{\ell}}$ for $\ell=0,\ldots,L$. In this paper, we consider the case where $N_{L+1}=d$. The number $L$ is called the depth of neural networks. Additionally, $S:=\sum_{\ell=0}^{L}(\|A_{\ell}\|_{0}+\|b_{\ell}\|_{0})$ represents the total number of non-zero weights within the neural network.
We denote by $N(L,S)$ the set of neural networks with depth at most $L$ and the number of non-zero weights at most $S$.
1.3.3. Notations
[…]
1.4. Organization
The remainder of this article is organized as follows. Section 2 introduces the characteristic-driven generative learning method. A thorough analysis for this method is provided in Section 3. Section 4 presents numerical studies and discussions. Section 5 discusses related work and provides additional insights. Finally, Section 6 presents the conclusion and discusses future work. The proof of theoretical results can be found in the Appendix.
2. Characteristic Generative Learning
Dating back to Moser (1965); Dacorogna and Moser (1990), researchers proposed a continuous dynamic-induced approach for solving the normalizing equation (1.1). In the field of deep generative learning, flow-based models utilize ODE-dynamics to construct probability flows, effectively pushing-forward the prior distribution towards the target distribution. This family of generative models is represented by continuous normalizing flows (CNF) (Chen et al., 2018; Grathwohl et al., 2019) and their variants (Gao et al., 2019; Rozen et al., 2021; Gao et al., 2022; Lipman et al., 2023; Neklyudov et al., 2023; Albergo and Vanden-Eijnden, 2023). The major challenges faced by flow-based models revolve around two key questions:
  • Q1.
    During the training phase, how can we estimate the velocity field of the probability flow ODE?
  • Q2.
    During the sampling phase, how can we solve the probability flow ODE efficiently?
The goal of this section is to propose the characteristic learning that has potential to address the aforementioned questions. In Section 2.1, we derive a probability flow ODE based on the concept of stochastic interpolants and the method of characteristics. Subsequently, in Section 2.2, we propose a velocity matching approach using least-squares regression, which provides an efficient solution to Q1. To tackle Q2, we first solve the probability flow ODE numerically in Section 2.3. Then Section 2.4 introduces a regression problem to fit characteristics using the obtained numerical solutions. This leads to an efficient simulation-free sampling method for flow-based generative models. Additionally, the characteristic fitting is improved by incorporating a semi-group penalty strategy. Finally, we summarizes the training and sampling algorithms in Section 2.4.
2.1. Characteristics and Probability Flow ODE
In this work, we follow the framework of stochastic interpolant (Albergo and Vanden-Eijnden, 2023; Albergo et al., 2023b, a). Let $X_{0}$ and $X_{1}$ be two independent random variables drawn from $\mu_{0}$ and $\mu_{1}$, respectively. The spatially linear stochastic interpolant $X_{t}$ is the stochastic process defined as
(2.1) $X_{t}=\alpha(t)X_{0}+\beta(t)X_{1},\quad t\in(0,1),$
where $\alpha(t)$ and $\beta(t)$ are two scalar-valued functions satisfying the following condition.
Condition 1.
The coefficient functions $\alpha(t),\beta(t)\in C([0,1])$ satisfy
  1. (i)
    $\alpha(0)=\beta(1)=1$ and $\alpha(1)=\beta(0)=0$,
  2. (ii)
    $\alpha^{2}(t)+\beta^{2}(t)>0$ for each $t\in[0,1]$,
  3. (iii)
    $\alpha(t)$ and $-\beta(t)$ are monotonically decreasing, and
  4. (iv)
    $\dot{\alpha}(t),\ddot{\alpha}(t)\in C([0,1))$, $\dot{\alpha}(t)\alpha(t)\in C([0,1])$ and $\dot{\beta}(t),\ddot{\beta}(t)\in C([0,1])$.
In this paper, we focus on two examples shown in Table 2: linear interpolants and Föllmer flow. Both of them are widely-used in generative learning, such as Nichol and Dhariwal (2021); Liu et al. (2022); Albergo and Vanden-Eijnden (2023); Albergo et al. (2023a); Lipman et al. (2023); Chang et al. (2024).
Table 2. Two examples of spatially linear interpolant.
$\alpha(t)$ $\beta(t)$
Linear interpolants (Liu et al., 2022; Lipman et al., 2023) $1-t$ $t$
Föllmer flow (Chang et al., 2024; Jiao et al., 2024) $\sqrt{1-t^{2}}$ $t$
Denote by $\mu_{t}$ the distribution of the process $X_{t}$ for each $t\in(0,1)$. The following proposition demonstrates that $\mu_{t}$ has a density $\rho_{t}$ that interpolates between $\rho_{0}$ and $\rho_{1}$. Further, the density $\rho_{t}$ satisfies the continuous equation.
Proposition 2.1 (Transport equation).
The distribution of the stochastic interpolant $X_{t}$ has a density function $\rho(t,x)$ satisfies $\rho(0,x)=\rho_{0}(x)$, $\rho(1,x)=\rho_{1}(x)$, and
$\rho(t,x)=\frac{1}{\beta_{t}}\int_{{\mathbb{R}}^{d}}\rho_{0}(x_{0})\rho_{1}\Big{(}\frac{x-\alpha_{t}x_{0}}{\beta_{t}}\Big{)}\,\mathrm{d}x_{0}=\frac{1}{\alpha_{t}}\int_{{\mathbb{R}}^{d}}\rho_{0}\Big{(}\frac{x-\beta_{t}x_{1}}{\alpha_{t}}\Big{)}\rho_{1}(x_{1})\,\mathrm{d}x_{1},$
for each time $t\in(0,1)$. Further, the density $\rho(t,x)$ solves the linear transport equation
(2.2) $\partial_{t}\rho(t,x)+\nabla\cdot(b^{*}(t,x)\rho(t,x))=0,$
where the velocity field is defined as
(2.3) $b^{*}(t,x)={\mathbb{E}}\big{[}\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}X_{1}\big{|}X_{t}=x\big{]}.$
As our primary objective is to generate samples that obey to the target distribution, we now consider the transport equation (2.2) from the lens of particles. It suffices to consider characteristics (Courant and Hilbert., 1989, Section II.2), along which the transport equation becomes an ODE
(2.4) $\,\mathrm{d}x(t)=b^{*}(t,x(t))\,\mathrm{d}t,$
where $x(t)$ is a characteristic, representing the position of particle at time $t\in(0,1)$, and $b^{*}$ is the associated velocity field that moves particles around. This characteristic ODE (Evans, 2010, Section 3.2) is known as the probability flow ODE (Song et al., 2021c). Further, the associated two-parameter continuous normalizing flow
$g_{t,s}^{*}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d},\quad x_{t}\mapsto x_{s},\quad 0\leq t\leq s\leq 1,$
where $x_{s}=x(s)$ represents the solution of (2.4) at time $s$ given $x(t)=x_{t}$. Notice that the flow $g_{t,s}^{*}$ pushes-forward the distribution $\mu_{t}$ onto $\mu_{s}$, that is,
$(g_{t,s}^{*})_{\sharp}\mu_{t}=\mu_{s},\quad 0\leq t\leq s<1.$
It is evident that the flow $g_{t,s}^{*}$ satisfies the semi-group property as follows.
Proposition 2.2 (Semi-group property).
For each $x\in{\mathbb{R}}^{d}$, it holds that
  1. (i)
    $g_{t,t}^{*}(x)-x=0$ for each $0\leq t\leq 1$, and
  2. (ii)
    $g_{t,s}^{*}(x)=g_{r,s}^{*}\circ g_{t,r}^{*}(x)$ for each $0\leq t\leq r\leq s\leq 1$.
Observe that the flow $g^{*}$ satisfies the normalizing equation (1.1), and for each fixed $x_{t}$, $\{g(t,s,x_{t})\}_{s\geq t}$ is a part of the characteristic. Consequently, the generative learning can be reduced to the problem of estimate the characteristic $g^{*}(t,s,x)$, which minimizes the following quadratic risk
(2.5) ${\mathcal{R}}(g)=\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}{\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\|Z_{s}-g(t,s,Z_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}s\mathrm{d}t,$
where $Z_{t}=g_{0,t}^{*}(Z_{0})$, $Z_{s}=g_{0,s}^{*}(Z_{0})$ and $T\in(1/2,1)$ is a pre-specified stopping time.
Given that the distributions of $Z_{t}$ and $Z_{s}$ in (2.5) are unknown, it is necessary to estimate them prior to minimizing (2.5). To achieve this, the velocity field is initially estimated in Section 2.2, followed by the utilization of the Euler method to numerically solve the probability flow ODE in Section 2.3. The resulting numerical solutions provide approximations of $(Z_{t},Z_{s})$, which are then utilized to approximate the population risk (2.5) in Section 2.4.
2.2. Velocity Matching
According to (2.3), for each fixed stopping time $T\in(1/2,1)$, the velocity field $b^{*}$ is the minimizer of following functional
(2.6) ${\mathcal{L}}(b)=\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{(X_{0},X_{1})}\Big{[}\|\dot{\alpha}(t)X_{0}+\dot{\beta}(t)X_{1}-b(t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t,$
where $X_{t}$ is the stochastic interpolant defined as (2.1).
Let $\{X_{0}^{(i)}\}_{i=1}^{n}$ and $\{X_{1}^{(i)}\}_{i=1}^{n}$ be two sets of independent copies of $X_{0}\sim\mu_{0}$ and $X_{1}\sim\mu_{1}$, respectively. Additionally, let $\{t^{(i)}\}_{i=1}^{n}$ be a set of $n$ i.i.d. random variables drawn from the uniform distribution on $[0,T]$. Denote by ${\mathcal{S}}=\{(t^{(i)},X_{0}^{(i)},X_{1}^{(i)})\}_{i=1}^{n}$. Then the empirical risk associated with (2.6) is defined as
(2.7) ^$\widehat{{\mathcal{L}}}_{n}(b)=\frac{1}{n}\sum_{i=1}^{n}\|\dot{\alpha}(t^{(i)})X_{0}^{(i)}+\dot{\beta}(t^{(i)})X_{1}^{(i)}-b(t^{(i)},X_{t}^{(i)})\|_{2}^{2},$
where $X_{t}^{(i)}=\alpha(t^{(i)})X_{0}^{(i)}+\beta(t^{(i)})X_{1}^{(i)}$. This induces the empirical risk minimizer
(2.8) $\widehat{b}\in\operatorname*{arg\,min}_{b\in{\mathscr{B}}}\widehat{{\mathcal{L}}}_{n}(b),$
where ${\mathscr{B}}$ is a vector-valued deep neural network class. The detailed velocity matching algorithm is shown in Algorithm 1. This approach for velocity matching is also used by rectified flow (Liu et al., 2022) and flow matching (Lipman et al., 2023).
Algorithm 1 Velocity matching.
1:Observations $\{X_{1}^{(i)}\}_{i=1}^{n}\sim^{\operatorname{i.i.d.}}\mu_{1}$.
2:Drawn $\{X_{0}^{(i)}\}_{i=1}^{n}\sim^{\operatorname{i.i.d.}}\mu_{0}=N(0,I_{d})$.
3:Drawn $\{t^{(i)}\}_{i=1}^{n}\sim^{\operatorname{i.i.d.}}\operatorname{Unif}[0,T]$.
4:Construct empirical interpolants via (2.1).
5:Initialize the deep neural network $b_{\theta}:{\mathbb{R}}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$.
6:repeat
7:     Compute the empirical risk ^$\widehat{{\mathcal{L}}}_{n}(b_{\theta})$ in (2.7).
8:     Compute the gradient $\nabla_{\theta}\widehat{{\mathcal{L}}}_{n}(b_{\theta})$.
9:     Gradient descent update $\theta\leftarrow\theta-\alpha\nabla_{\theta}\widehat{{\mathcal{L}}}_{n}(b_{\theta})$.
10:until converged
11:An estimator $\widehat{b}=b_{\theta}$ of the velocity field.
Besides the error of velocity matching itself, one is actually interested in the error of profitability flow with estimated velocity, for which
(2.9) $\,\mathrm{d}\widehat{x}(t)=\widehat{b}(t,\widehat{x}(t))\,\mathrm{d}t.$
Denote by $\widehat{\mu}_{T}$ the push-forward distribution of $\mu_{0}$ by ODE (2.9) at time $T$. The following corollary characteristic the 2-Wasserstein distance between $\mu_{T}$ and $\widehat{\mu}_{T}$.
Proposition 2.3.
The 2-Wasserstein distance between the probability flow (2.8) and the estimated flow (2.9) at the stopping time $T$ is bounded as follows:
$W_{2}(\widehat{\mu}_{T},\mu_{T})\leq\exp(\|\widehat{b}\|_{\operatorname{Lip}}T)\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|\widehat{b}(t,X_{t})-b^{*}(t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t,$
where the Lipschitz constant of the velocity estimator $\widehat{b}$ is defined as
$\|\widehat{b}\|_{\operatorname{Lip}}=\sup\Big{\{}\|\nabla\widehat{b}(t,x)\|_{\operatorname{op}}:(t,x)\in(0,1)\times{\mathbb{R}}^{d}\Big{\}}.$
The error bound in Proposition 2.3 depends exponentially on the Lipschitz constant of the estimated velocity, which requires us to constrain the Lipschitz constant of the deep neural network. In practical applications, various techniques for restricting the Lipschitz constant of deep neural networks have been proposed, such as weight clipping (Arjovsky et al., 2017), gradient penalty (Gulrajani et al., 2017), spectral normalization (Miyato et al., 2018), and Lipschitz network (Zhang et al., 2022).
2.3. Euler Sampling
We turn to focus on sampling from the estimated probability flow equation in this section. We replace the velocity $b^{*}$ in probability flow ODE (2.4) by its estimated counterpart $\widehat{b}$ defined as (2.8), and employ the forward Euler method (Iserles, 2008; Butcher, 2016) to discretize this ODE. Denote by $K\in{\mathbb{N}}_{+}$ the total number of steps, then the step size is given as $\tau=T/K$. Define $\{t_{k}=k\tau\}_{k=0}^{K}$ as the set of time points. Applying forward Euler method deduces the following scheme:
(2.10) $x_{k}=x_{k-1}+\tau\widehat{b}(t_{k-1},x_{k-1}),\quad 1\leq k\leq K.$
Similar to the flow $g_{t,s}^{*}$ associated with ODE (2.4), the flow induced by Euler method (2.10) is defined as $\widehat{E}_{k,\ell}^{\tau}:x_{k}\mapsto x_{\ell}$ for each $0\leq k\leq\ell\leq K$. The following proposition demonstrates that the Euler flow $\widehat{E}_{k,\ell}^{\tau}$ inherits the semi-group property of continuous flow $g_{t,s}^{*}$ in Proposition 2.2.
Proposition 2.4 (Semi-group property).
For each $x\in{\mathbb{R}}^{d}$, it holds that
  1. (i)
    $\widehat{E}_{k,k}^{\tau}(x)-x=0$ for each $0\leq k\leq\ell\leq K$, and
  2. (ii)
    $\widehat{E}_{k,\ell}^{\tau}(x)=\widehat{E}_{j,\ell}^{\tau}\circ\widehat{E}_{k,j}^{\tau}(x)$ for each $0\leq k\leq j\leq\ell\leq K$.
It is true that the Euler flow $\widehat{E}_{k,\ell}^{\tau}$ pushes-forward the distribution $\mu_{k\tau}$ approximately onto the distribution $\mu_{\ell\tau}$ for each $1\leq k\leq\ell\leq K$. Hence the Euler flow $\widehat{E}_{k,\ell}^{\tau}$ is an alternative approach for evaluating the flow $g_{t,s}^{*}$. However, it is important to note that the Euler sampling incurs a substantial computational cost, as it necessitates a large number of velocity network evaluations. This makes the Euler sampling encounter challenges in time-sensitive application scenarios. Therefore, there is a pressing need to develop an efficient simulation-free approach for evaluating the flow $g_{t,s}^{*}$.
2.4. Characteristic Fitting and Semi-Group Penalty
In order to estimate the flow $g_{t,s}^{*}$ via a simulation-free approach, we leverage a deep neural network to fit characteristics using data samples obtained by Euler flow. Let ${\mathcal{Z}}=\{\widehat{Z}_{0}^{(i)}\}_{i=1}^{m}$ be a set of $m$ random copies of $\widehat{Z}_{0}\sim\mu_{0}$. Further, one obtains $m$ discrete characteristics $\{(\widehat{Z}_{k}^{(i)})_{k=0}^{K}\}_{i=1}^{m}$ by the Euler method (2.10) with $\widehat{Z}_{k}^{(i)}=\widehat{E}_{0,k}^{\tau}(\widehat{Z}_{0}^{(i)})$. Then the empirical counterpart of (2.5) is defined as
[…]
The characteristic generator can be obtained by the empirical risk minimization
(2.12) $\widehat{g}\in\operatorname*{arg\,min}_{g\in{\mathscr{G}}}\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g),$
where ${\mathscr{G}}$ is a set of vector-valued deep neural networks. It is important to note that the estimator $\widehat{g}_{t,s}$ serves as a neural network approximation for the flow $g_{t,s}^{*}$, eliminating the need to simulate the ODE when evaluating $\widehat{g}_{t,s}$. Therefore, the characteristic generator $\widehat{g}_{t,s}$ is an efficient alternative to the Euler flow $\widehat{E}_{k,\ell}^{\tau}$ in (2.10). Additionally, the idea of fitting characteristics using deep neural network is also used by previous literature, such as DSNO (Zheng et al., 2023a) and CTM (Kim et al., 2024).
Despite that both the continuous flow $g_{t,s}^{*}$ and Euler flow $\widehat{E}_{k,\ell}^{\tau}$ satisfy semi-group property (Propositions 2.2 and 2.4), the characteristic generator $\widehat{g}$ defined as (2.12) does not satisfy the semi-group property in general. In order to ensure the long-term stability of the estimator, we introduce the semi-group constraint, which requires
$\Delta_{kj\ell}(g)=\|g_{k\tau,\ell\tau}(\widehat{Z}_{k}^{(i)})-g_{j\tau,\ell\tau}\circ\widehat{E}_{k,j}^{\tau}(\widehat{Z}_{k}^{(i)})\|_{2}$
to be as small as possible for each $0\leq k\leq j\leq\ell\leq K$. Consequently, we propose the semi-group-penalized risk
(2.13) ^$\widehat{{\mathcal{R}}}_{m,K}^{\operatorname{Euler},\lambda}(g)=\widehat{{\mathcal{R}}}_{m,K}^{\operatorname{Euler}}(g)+\lambda\widehat{{\mathscr{P}}}(g),$
where $\lambda>0$ is the penalty parameter, and the semi-group-penalty is defined as
$\widehat{{\mathscr{P}}}(g)=\frac{2}{mK^{2}}\sum_{i=1}^{m}\Bigg{\{}\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\sum_{\ell=j+1}^{K-1}\Delta_{kj\ell}^{2}(g)\Bigg{\}}.$
The complete training and sampling procedures of the characteristic generator is summarized in Algorithms 23 and 4.
Algorithm 2 Training procedure of characteristic generator.
1:Velocity estimator $\widehat{b}$.
2:# Euler sampling
3:for $i=1,\ldots,m$ do
4:     Sample initial value $\widehat{Z}_{0}^{(i)}\sim\mu_{0}=N(0,I_{d})$.
5:     for $k=1,\ldots,K$ do
6:         $\widehat{Z}_{k}^{(i)}\leftarrow\widehat{E}_{k-1,k}^{\tau}(\widehat{Z}_{k-1}^{(i)})$ by (2.10).
7:     end for
8:end for
9:# Characteristic fitting
10:Initialize the neural network $g_{\phi}:{\mathbb{R}}\times{\mathbb{R}}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$.
11:repeat
12:     Compute the empirical risk ^$\widehat{{\mathcal{R}}}_{m,K,n}(g_{\phi})$ in (2.11) or (2.13).
13:     Compute the gradient $\nabla_{\phi}\widehat{{\mathcal{R}}}_{m,K,n}(g_{\phi})$.
14:     Gradient descent update $\phi\leftarrow\phi-\alpha\nabla_{\phi}\widehat{{\mathcal{R}}}_{m,K,n}(g_{\phi})$.
15:until converged
16:Characteristic generator $\widehat{g}_{t,s}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ for each $0\leq t\leq s<T$.
Algorithm 3 One-step sampling of characteristic generator.
1:Characteristic generator $\widehat{g}_{t,s}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ for each $0\leq t\leq s<T$.
2:Sample initial value $\widehat{Z}_{0}\sim\mu_{0}=N(0,I_{d})$.
3:$\widehat{Z}_{T}\leftarrow\widehat{g}_{0,T}(\widehat{Z}_{0})$.
4:Generated samples $\widehat{Z}_{T}$.
Algorithm 4 Fine-grained sampling of characteristic generator.
1:Characteristic generator $\widehat{g}_{t,s}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ for each $0\leq t\leq s<T$.
2:Sample initial value $\widehat{Z}_{0}\sim\mu_{0}=N(0,I_{d})$.
3:Choose a sequence of time points $0=t_{0}<\ldots<t_{K}=T$.
4:for $k=1,\ldots,K$ do
5:     $\widehat{Z}_{T}\leftarrow\widehat{g}_{t_{k-1},t_{k}}(\widehat{Z}_{k-1})$.
6:end for
7:Generated samples $\widehat{Z}_{T}$.
As shown in Algorithm 3, in the sampling phase, one only needs to evaluate the characteristic generator $\widehat{g}_{0,T}$ once. As a consequence, our generator diminishes the sampling time in comparison to Euler sampling. However, it necessitates a significant number of ODE simulations during the training phase as Algorithm 2. Nevertheless, the benefits outweigh the costs. In practical application scenarios, one can simulate ODE (2.10) and fit the probability flow (2.12) (Algorithm 2) on extensive and high-performance computing platforms. Consequently, the derived estimators $\widehat{g}_{0,T}$ can be deployed in computationally constrained and time-sensitive application scenarios.
The characteristic generator is not restricted to one-step generation as Algorithm 3. On the contrary, it can be employed iteratively to produce refined sampling algorithms, thereby enhancing the quality of generation, albeit with a slight increase in computational cost. Algorithm 4 presents the comprehensive procedure for achieving this fine-grained sampling.
3. Convergence Rates Analysis
In this section, we present a comprehensive convergence rate analysis for the characteristic generator. We begin by illustrating Assumptions 1 and 2 in Section 3.1, and then establish regularity properties of the probability flow ODE (2.4) in Section 3.2. In Section 3.3 and Section 3.4, we propose a non-asymptotic error analysis for velocity matching and Euler sampling, respectively. In Section 3.5, a convergence rate analysis for the characteristic generator is established. Finally, in Section 3.6, we apply the aforementioned analysis to two widely-used types of probability flow ODE: linear interpolant and Föllmer flow.
3.1. Discussions of Assumptions
Assumption 1 is standard and commonly-used in flow-based generative models (Liu et al., 2022; Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023). The following proposition demonstrates that the velocity field $b^{*}(t,x)$ is a spatial linear combination of $x$ and the score function $s^{*}(t,x)=\nabla\log\rho_{t}(x)$. This connection suggests that stochastic interpolants that satisfy Assumption 1 are closely related to diffusion models (Ho et al., 2020; Song et al., 2021c).
Proposition 3.1 (Velocity and score).
Suppose Assumption 1 holds. Then the following equality holds:
$b^{*}(t,x)=\frac{\dot{\beta}_{t}}{\beta_{t}}x+\alpha_{t}^{2}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}s^{*}(t,x),\quad(t,x)\in(0,1)\times{\mathbb{R}}^{d}.$
Assumption 2 necessitates that the target distribution be a mixture of a Gaussian distribution and a distribution of compact support. Previous research (Lee et al., 2023; Oko et al., 2023; Chen et al., 2023a, d) has investigated the assumption of compact support in the target distribution. However, the probability flow that pushes a Gaussian prior distribution towards a target distribution with compact support may lack regularity because the velocity field is not generally Lipschitz. Therefore, it becomes crucial to impose additional intractable regularity properties on the probability flow ODE. In contrast, the probability flow towards a distribution of compact support with Gaussian smoothing exhibits high regularity even without any additional assumptions. The regularity properties of the probability flow ODE are established in Section 3.2. Furthermore, Assumption 2 covers a wide range of target distributions. In essence, our generative model has the capability to learn arbitrarily complex distributions with compact supports after Gaussian smoothing. This guarantees the application of our model to a variety of generative problems.
3.2. Properties of Probability Flow ODE
In this section, we present elementary properties of the probability flow ODE (2.4). To begin with, the following proposition shows that the velocity fields is local bounded.
Proposition 3.2 (Local bounded velocity).
Suppose Assumptions 1 and 2 hold. Let $T\in(1/2,1)$ and $R\in(1,+\infty)$. Then it follows that
$\max_{1\leq k\leq d}|b_{k}^{*}(t,x)|\leq B_{\operatorname{vel}}R,\quad(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty},$
where the constant $B_{\operatorname{vel}}$ only depends on $d$ and $\sigma$.
With the aid of Proposition 3.2, we show that the probability flow and its time derivatives are also local bounded.
Corollary 3.3 (Local bounded flow).
Suppose Assumptions 1 and 2 hold. Let $T\in(1/2,1)$ and $R\in(1,+\infty)$. Then it follows that
$\max_{1\leq k\leq d}|g_{k}^{*}(t,s,x)|\leq B_{\operatorname{flow}}R,\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty},$
where the constant $B_{\operatorname{flow}}$ only depends on $d$ and $\sigma$.
Corollary 3.4 (Local bounded time derivatives of flow).
Suppose Assumptions 1 and 2 hold. Let $T\in(1/2,1)$ and $R\in(1,+\infty)$. Then it follows that
$\max\Big{\{}\|\partial_{t}g^{*}(t,s,x)\|_{2},\|\partial_{s}g^{*}(t,s,x)\|_{2}\Big{\}}\leq B_{\operatorname{vel}}R,\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty}.$
Further, the spatial gradient of the velocity field is uniformly bounded in spectral norm, as demonstrated by the next proposition.
Proposition 3.5 (Bounded spatial gradient of velocity).
Suppose Assumptions 1 and 2 hold. Then there exists a constant $G<\infty$ such that
$\|\nabla b^{*}(t,x)\|_{\operatorname{op}}\leq G,\quad(t,x)\in(0,1)\times{\mathbb{R}}^{d},$
where the constant $G$ only depends on $d$ and $\sigma$.
A direct consequence of this proposition is that, under Assumptions 1 and 2, the velocity field is uniformly Lipschitz continuous with respect to the spatial variable, as shown by the following corollary.
Corollary 3.6 (Lipschitz continuity).
Suppose Assumptions 1 and 2 hold. Then
$\|b^{*}(t,x)-b^{*}(t,x^{\prime})\|_{2}\leq G\|x-x^{\prime}\|_{2},\quad(t,x,x^{\prime})\in(0,1)\times{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}.$
The uniform Lipschitz continuity of the velocity field or score plays a crucial role in controlling the error accumulation along the ODE or Euler method, as detailed in Corollary 3.10 and Theorem 3.11. Previous studies have often made the direct assumption of Lipschitz continuity for the velocity field or score at each time $t\in(0,1)$ (Chen et al., 2023d, a, c, b; Benton et al., 2024b; Gao and Zhu, 2024). However, this assumption appears to be restrictive, as it is difficult to verify. On the contrary, Assumptions 1 and 2 are easily satisfied and cover a large amount of generative tasks.
Another important consequence of Proposition 3.5 is shown as follows.
Corollary 3.7 (Bounded spatial gradient of flow).
Suppose Assumptions 1 and 2 hold. Then
$\|\nabla g^{*}(t,s,x)\|_{\operatorname{op}}\leq\exp(G(s-t)),\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty}.$
Finally, we state the bound of time derivative of the velocity in the following proposition.
Proposition 3.8 (Bounded time derivative of velocity).
Suppose Assumptions 1 and 2 hold. Let $T\in(1/2,1)$ and $R\in(1,+\infty)$. Then it follows that
$\|\partial_{t}b^{*}(t,x)\|_{2}\leq D\kappa(T)R:=D\sup_{t\in[0,T]}\Big{(}\frac{\dot{\alpha}_{t}^{2}}{\alpha_{t}^{2}}+\frac{|\ddot{\alpha}_{t}|}{\alpha_{t}}\Big{)}R,\quad(t,x)\in[0,T]\times{\mathbb{B}}_{R}^{\infty},$
where the constant $D$ only depends on $d$ and $\sigma$.
Proposition 3.8 establishes the Lipschitz continuity of the velocity in time, a crucial requirement for controlling the discretization error of the Euler method. For detailed illustrations, refer to Theorem 3.11. In contrast, previous work (Gao and Zhu, 2024, Assumption 2) simply assumes the Lipschitz continuity of the score in time.
3.3. Analysis for Velocity Matching
In this section, we focus on the time-averaged $L^{2}$-error of the velocity estimator $\widehat{b}$ in (2.8), that is,
(3.1) ${\mathcal{E}}_{T}(\widehat{b})=\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b^{*}(t,X_{t})-\widehat{b}(t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t.$
As shown in Proposition 2.3, the $2$-Wasserstein error is bounded by the averaged $L^{2}$-risk of the velocity estimator. The majority of existing literature on theoretical analysis of diffusion and flow-based generative models commonly assumes that the $L^{2}$-risk of score or velocity matching is sufficiently small (Lee et al., 2022, 2023; Chen et al., 2023d, c; Benton et al., 2024a, b; Gao and Zhu, 2024). However, this line of research lacks the ability to quantitatively characterize the convergence rate of velocity matching with respect to the number of samples. Additionally, no prior theoretical guidance for the selection of neural networks is provided in this literature. To the best of our knowledge, only a limited number of studies have specifically focused on investigating the convergence rates of score matching (Oko et al., 2023; Chen et al., 2023b; Han et al., 2024) and velocity matching (Chang et al., 2024; Gao et al., 2024; Jiao et al., 2024).
In this work, we aim to establish a non-asymptotic error bound for the $L^{2}$-risk of the estimated velocity. The main result is stated as follows.
Theorem 3.9 (Convergence rate for velocity matching).
Suppose Assumptions 1 and 2 hold. Let $T\in(1/2,1)$. Set the hypothesis class ${\mathscr{B}}$ as a deep neural network class, which is defined as
=${\mathscr{B}}=\left\{b\in N(L,S):\begin{aligned} &\|b(t,x)\|_{\infty}\leq B_{\operatorname{vel}}\log^{1/2}n,\\ &\|\partial_{t}b(t,x)\|_{2}\leq 3D\kappa(T)\log^{1/2}n,\\ &\|\nabla b(t,x)\|_{\operatorname{op}}\leq 3G,\leavevmode\nobreak\ (t,x)\in[0,T]\times{\mathbb{R}}^{d}\end{aligned}\right\},$
where the depth and the width of the neural network are given, respectively, as $L=C$ and $S=C\kappa^{\frac{2(d+1)}{d+3}}(T)n^{\frac{d+1}{d+3}}$. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}\big{[}{\mathcal{E}}_{T}(\widehat{b})\big{]}\leq C\kappa^{2}(T)n^{-\frac{2}{d+3}}\log^{2}n,$
where $C$ is a constant only depending on $d$ and $\sigma$.
The rate of velocity matching in Theorem 3.9 is consistent with the minimax optimal rate ${\mathcal{O}}(n^{-\frac{2}{d+3}})$ in nonparametric regression (Stone, 1982; Yang and Barron, 1999; Gyorfi et al., 2002; Tsybakov, 2009) given that the target function is Lipschitz continuous. Moreover, our theoretical findings align with convergence rates of nonparametric regression using deep neural networks (Bauer and Kohler, 2019; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Farrell et al., 2021; Kohler et al., 2022; Jiao et al., 2023a). It is noteworthy that our results improve upon the rate ${\mathcal{O}}(n^{-\frac{2}{d+5}})$ derived by Chen et al. (2023b); Chang et al. (2024).
In Theorem 3.9, the hypothesis class ${\mathscr{B}}$ is set as a deep neural network class with Lipschitz constraints. The assumption of uniformly Lipschitz continuity of velocity estimator plays a crucial role in controlling the discretization error induced by Euler method. For further details, refer to Theorem 3.11. This assumption is standard in the theoretical analysis for flow-based or diffusion models, as considered by Kwon et al. (2022, Assumption A2) and Chen et al. (2023c, Assumption 3). In practical applications, various techniques for restricting the Lipschitz constant of deep neural networks have been reviewed in Section 2.2. In the theoretical perspective, the approximation properties of deep neural network with Lipschitz constraint has been studied by Chen et al. (2022); Huang et al. (2022); Jiao et al. (2023b); Ding et al. (2024). In this work, an approximation error bound for deep neural networks with Lipschitz constraint is established in Section H.
With the aid of Proposition 2.3 and Theorem 3.9, we obtain a 2-Wasserstein error of the estimated flow (2.9) as follows.
Corollary 3.10.
Under the same conditions as Theorem 3.9. The 2-Wasserstein between the probability flow (2.8) and the estimated flow (2.9) at the stopping time $T$ is bounded as follows:
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}(\widehat{\mu}_{T},\mu_{T})\Big{]}\leq C\kappa^{2}(T)n^{-\frac{2}{d+3}}\log^{2}(n),$
where the constant $C$ only depends on $d$ and $\sigma$.
Corollary 3.10 highlights that the 2-Wasserstein error of the estimated flow converges to zero at a rate of ${\mathcal{O}}(n^{-\frac{2}{d+3}})$, omitting some logarithmic factors. In contrast, the $2$-Wasserstein error bounds derived by Benton et al. (2024b, Theorem 1) and Albergo and Vanden-Eijnden (2023, Proposition 3) are under a “black-box” assumption that the velocity matching error is sufficiently small. Therefore, these results can not capture how the error converges to zero as the number of samples increases.
3.4. Analysis for Euler Sampling
The main objective of this section is to estimate the 2-Wasserstein error of the Euler flow (2.10). Despite that there has been numerous studies on the sampling error of SDE-based diffusion models (Lee et al., 2022, 2023; Chen et al., 2023d; Benton et al., 2024a), as well as flow-based methods (Chen et al., 2023c; Gao and Zhu, 2024; Li et al., 2024c, a), most of these works assume that the velocity matching error is sufficiently small. Furthermore, there is limited work that integrates the sampling error with the velocity matching error (Chang et al., 2024; Gao et al., 2024; Jiao et al., 2024). To address this gap, we propose the following theorem.
Theorem 3.11 (Error analysis for Euler flow).
Under the same conditions as Theorem 3.9. Let the number of time steps for Euler method $K$ be a positive integer. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}\Big{(}(\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{T}\Big{)}\Big{]}\leq C\kappa^{2}(T)\Big{\{}n^{-\frac{2}{d+3}}\log^{2}n+\frac{1}{K^{2}}\log n\Big{\}},$
where the constant $C$ only depends on $d$ and $\sigma$. Further, if $K\geq Cn^{\frac{1}{d+3}}$, then it follows that
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}\Big{(}(\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{T}\Big{)}\Big{]}\leq C\kappa^{2}(T)n^{-\frac{2}{d+3}}\log^{2}n.$
The error bound of the Euler flow (2.10), as derived in Theorem 3.11, can be roughly divided into two main components. The first term arises from velocity matching, aligning with the error bound presented in Corollary 3.10. The second term corresponds to the discretization error introduced by Euler method. Moreover, as the number of time steps $K$ increases, the error bound in Theorem 3.11 converges to that in Corollary 3.10.
Based on Theorem 3.11, we can derive a $2$-Wasserstein error bound between the target distribution and the push-forward distribution of $\mu_{0}$ by Euler flow $\widehat{E}_{0,K}^{\tau}$ as follows.
Corollary 3.12.
Under the same conditions as Theorem 3.11. Suppose the number of time steps for Euler method $K$ satisfies $K\geq Cn^{\frac{1}{d+3}}$. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}\Big{(}(\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big{)}\Big{]}\leq C\kappa^{2}(T)n^{-\frac{2}{d+3}}\log^{2}(n)+2\max\{\alpha_{T}^{2},(1-\beta_{T})^{2}\}W_{2}^{2}(\mu_{0},\mu_{1}),$
where the constant $C$ only depends on $d$ and $\sigma$.
Corollary 3.12 presents a bound for the total error of flow-based generative models. The first term in the error bound corresponds to the 2-Wasserstein error of the Euler flow, as derived in Theorem 3.11. The second term captures the convergence of the interpolant distribution $\mu_{T}$ to the target distribution $\mu_{1}$. It is worth noting that as the stopping time $T$ approaches one, the first term tends to infinity, while the second term simultaneously decreases. This trade-off within the error bound highlights the importance of carefully selecting the stopping time $T$ and provides practical guidance for determining it in real-world applications.
3.5. Analysis for Characteristic Generator
Despite the empirical success of simulation-free one-step approaches for the efficient sampling of flow-based generative models (Salimans and Ho, 2022; Song et al., 2023; Zheng et al., 2023a; Kim et al., 2024; Ren et al., 2024), the theoretical analysis for these line of methods still remains unclear. In this section, we establish a thorough analysis for the characteristic generator. To the best of our knowledge, this is the first analysis for one-step sampling method.
To measure the error of the characteristic generator, we focus on the time-average squared 2-Wasserstein distance between the distribution associated to the characteristic generator $\widehat{g}$ and the target distribution
(3.2) ${\mathcal{D}}(\widehat{g})=\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}W_{2}^{2}\Big{(}(\widehat{g}_{t,s})_{\sharp}\mu_{t},\mu_{s}\Big{)}\,\mathrm{d}s\mathrm{d}t.$
The main result is stated as follows.
Theorem 3.13 (Error analysis for characteristic generator).
Under the same conditions as Theorem 3.11. Further, set the hypothesis class ${\mathscr{G}}$ as a deep neural network class, which is defined as
[…]
where the depth and the width of the neural network are given, respectively, as $L=C$ and $S=Cm^{\frac{d+2}{d+4}}$. Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{D}}(\widehat{g})\big{]}\leq C\kappa^{2}(T)n^{-\frac{2}{d+3}}\log^{2}n+Cm^{-\frac{2}{d+4}}\log^{2}m+C\Big{(}\frac{\log m}{K}+\frac{\log n}{K^{2}}\Big{)},$
where the constant $C$ only depends on $d$ and $\sigma$. Furthermore, if the number of time steps $K$ for Euler method and the number of samples $m$ for characteristic fitting satisfy
(3.3) $K\geq\max\big{\{}Cn^{\frac{1}{d+3}},C\kappa^{-2}(T)n^{-\frac{2}{d+3}}\big{\}}\quad\text{and}\quad m\geq C\kappa^{-(d+4)}(T)n^{\frac{d+4}{d+3}},$
respectively, then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{D}}(\widehat{g})\big{]}\leq C\kappa^{2}(T)n^{-\frac{2}{d+3}}\log^{2}n.$
In contrast to the error bound of Euler sampling in Theorem 3.11, the error bound of the characteristic generator in Theorem 3.13 incorporates an additional error term ${\mathcal{O}}(m^{-\frac{2}{d+4}})$. This error term corresponds to the error of the standard nonparametric regression for characteristic fitting, attaining the minimax optimality (Stone, 1982; Yang and Barron, 1999; Gyorfi et al., 2002; Tsybakov, 2009) given that $g^{*}$ is Lipschitz continuous.
It is noteworthy that the number of samples $m$ for characteristic fitting can be arbitrarily large because training samples can be generated from copies of $Z_{0}\sim\mu_{0}$ using Euler sampling (2.10). Without loss of generality, we consider the case where $m\gg n$. In this scenario, the error bound in Theorem 3.13 aligns with the convergence rate in Theorem 3.11.
In the context of distillation, Euler flow (2.10) is commonly referred as the “teacher” model. Theorem 3.13 guarantees that, when the number of teacher samples is sufficiently large, the characteristic generator can generate new samples that are as good as those generated by the teacher model. However, the theorem also highlights that the teacher model serves as a bottleneck for the characteristic generator, as it cannot surpass the generative quality of the teacher model. These theoretical findings align with empirical observations reported in prior studies (Salimans and Ho, 2022; Song et al., 2023; Kim et al., 2024). One potential approach to overcome this bottleneck is by combining these one-step generative models with GANs, as demonstrated by Lu et al. (2023b); Kim et al. (2024).
3.6. Applications
In this section, the theoretical analysis is applied to two types of flow-based method in Table 2: linear interpolant and Föllmer flow. The convergence rates of them are shown in Corollaries 3.14 and 3.15, respectively.
Corollary 3.14 (Convergence rate of linear interpolant).
Under the same conditions as Theorem 3.13. Set the stopping time $T$ as
$T=1-Cn^{-\frac{1}{3(d+3)}}\log^{\frac{1}{2}}n.$
Suppose the number of time step $K$ satisfies $K\geq Cn^{\frac{1}{d+3}}$. Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}\Big{(}(\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big{)}\Big{]}\leq Cn^{-\frac{2}{3(d+3)}}\log n.$
Further, suppose the number of time step $K$ and number of samples $m$ satisfies (3.3). Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{D}}(\widehat{g})\big{]}\leq Cn^{-\frac{2}{3(d+3)}}\log n,$
where the constant $C$ is independent of $n$.
Corollary 3.15 (Convergence rate of Föllmer flow).
Under the same conditions as Theorem 3.13. Set the stopping time $T$ as
$T=1-Cn^{-\frac{2}{5(d+3)}}\log^{\frac{3}{5}}n.$
Suppose the number of time step $K$ satisfies $K\geq Cn^{\frac{1}{d+3}}$. Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}\Big{(}(\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big{)}\Big{]}\leq CCn^{-\frac{2}{5(d+3)}}\log^{\frac{3}{5}}n.$
Further, suppose the number of time step $K$ and number of samples $m$ satisfies (3.3), respectively. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{D}}(\widehat{g})\big{]}\leq Cn^{-\frac{2}{5(d+3)}}\log^{\frac{3}{5}}n,$
where the constant $C$ is independent of $n$.
4. Numerical Studies
In this section, we delve into the numerical performance of characteristic learning. To begin with, we introduce several technical improvements in Section 4.1. Subsequently, the experimental results and discussions are presented in Section 4.2.
4.1. Technical improvements
In this section several technical enhancements to the algorithms in Section 2 are introduced. Empirical evidence indicates that these methods exhibit superior numerical performance. It is noteworthy that improved algorithms in this section are mathematically equivalent to the previous ones in Section 2. Consequently, these improvements remain within the established mathematical framework and theoretical analysis. Specifically, in Section 4.1.1, we adopt a denoiser matching algorithm as a replacement for the velocity matching in Section 2.2. Additionally, in Section 4.1.2, we replace Euler method with the technique of exponential integrator. Finally, the modified characteristic learning algorithm is summarized in Section 4.1.3.
4.1.1. Denoiser matching
To begin with, we define the denoiser denoiser $D^{*}$ as
$D^{*}(t,x)={\mathbb{E}}\big{[}X_{1}\big{|}X_{t}=x\big{]},\quad(t,x)\in(0,1)\times{\mathbb{R}}^{d},$
which recovers $X_{1}$ from noised observation $X_{t}=\alpha_{t}X_{0}+\beta_{t}X_{1}$. It is apparent that $D^{*}$ minimizes the following functional for each $T\in(0,1)$,
(4.1) ${\mathcal{L}}(D)=\frac{1}{T}\int_{0}^{T}{\mathbb{E}}\Big{[}\|X_{1}-D(t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t.$
An estimator $\widehat{D}$ of the denoiser $D^{*}$ can be obtained by the empirical risk minimization similar to (2.8) using data set ${\mathcal{S}}=\{(t^{(i)},X_{0}^{(i)},X_{1}^{(i)})\}_{i=1}^{n}$. In the context of distillation for diffusion models, the denoiser network $\widehat{D}$ is referred as the “teacher” network.
By an argument similar to Proposition 3.1, the velocity field is a spatial linear combination of $x$ and denoiser $D^{*}(t,x)$, that is,
(4.2) $b^{*}(t,x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha_{t}}}{\alpha_{t}}\Big{)}D^{*}(t,x).$
Thus the denoiser matching is equivalent to estimating the velocity field, but the former has better numerical stability (Karras et al., 2022; Kim et al., 2024). Furthermore, the semi-linear form (4.2) enable us to use the exponential integrator, which is more stable than Euler method. See Section 4.1.2 for detailed discussions.
4.1.2. Exponential integrator
With the aid of the denoiser estimator $\widehat{D}$, the estimated probability flow (2.9) is replaced by
(4.3) $\frac{\,\mathrm{d}\widehat{x}(t)}{\,\mathrm{d}t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}}\widehat{x}(t)+\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\widehat{D}(t,\widehat{x}(t)),\quad t\in(0,1).$
Observe that the solution of the semi-linear ODE (4.3) can be exactly formulated by the “variation of constants” formula as
(4.4) $\widehat{x}(s)=\varPhi(t,s)\widehat{x}(t)+\int_{t}^{s}\psi(\tau,s)\widehat{D}(\tau,\widehat{x}(\tau))\,\mathrm{d}\tau,$
where $\varPhi(t,s)$ and $\psi(t,s)$ are defined as
$\varPhi(t,s)=\exp\Big{(}\int_{t}^{s}\frac{\dot{\alpha}_{\tau}}{\alpha_{\tau}}\,\mathrm{d}\tau\Big{)},\quad\psi(t,s)=\varPhi(t,s)\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)},\quad 0\leq t\leq s\leq T.$
By a similar argument to Euler method (2.10), we replace $\widehat{D}(\tau,\widehat{x}(\tau))$ in (4.4) by $\widehat{D}(t,\widehat{x}(t))$ and implies an explicit scheme
(4.5) $\widehat{x}(s)\approx\varPhi(t,s)\widehat{x}(t)+\varPsi(t,s)\widehat{D}(t,\widehat{x}(t)),\quad 0\leq t\leq s\leq T.$
where $\varPsi(t,s)$ is a integral defined as
$\varPsi(t,s)=\int_{t}^{s}\varPhi(\tau,s)\beta_{\tau}\Big{(}\frac{\dot{\beta}_{\tau}}{\beta_{\tau}}-\frac{\dot{\alpha}_{\tau}}{\alpha_{\tau}}\Big{)}\,\mathrm{d}\tau.$
Notice that the integrals $\varPhi$ and $\varPsi$ can be computed analytically given the interpolant coefficients $\alpha_{t}$ and $\beta_{t}$. The integral scheme (4.5) is commonly referred to as the first-order exponential integrator (Hochbruck and Ostermann, 2010), which has been utilized in sampling of diffusion models by Zhang and Chen (2023); Lu et al. (2022, 2023a); Zheng et al. (2023b). We display the generated images and corresponding FID using exponential integrator (4.5) in Figure 1.
Figure 1. Samples generated by probability flow ODE with exponential integrator.
In practice, we find the first-order exponential integrator outperforms other methods, such as Euler and Heun methods. Nevertheless, it is important to note that this integral scheme remains a first-order method, akin to the Euler method, and does not improve the convergence rate of the discretization error. Consequently, our analysis encompasses this integral scheme as well.
4.1.3. Characteristic fitting
In Section 2.4, we directly parameterize the probability flow by a deep neural network. However, in practice training such a neural network is unstable. In this section we present some technical tricks to get a more stable training algorithm by exploiting as much of the structure of the problem as possible without changing its mathematical nature.
Recall the explicit formulation (4.4) of the solution
$x(s)=\varPhi(t,s)x(t)+\varPsi(t,s)\frac{\int_{t}^{s}\psi(\tau,s)D^{*}(\tau,x(\tau))\,\mathrm{d}\tau}{\int_{t}^{s}\psi(\tau,s)\,\mathrm{d}\tau},$
where the fraction in the second term can be viewed as a weighted average of $D^{*}(\tau,x(\tau))$ on $[s,t]$. Our main idea is to approximate this term using a deep neural network $D_{{\mathcal{S}}}$, which is referred to the student model. Then the probability flow is parameterized by
(4.6) $g(t,s,x)=\varPhi(t,s)x+\varPsi(t,s)D_{{\mathcal{S}}}(t,s,x),\quad 0\leq t\leq s\leq T.$
Since the exact denoiser $D^{*}$ is unknown, the student network $D_{{\mathcal{S}}}$ can only be trained from the denoiser estimator $\widehat{D}$. Therefore, $\widehat{D}$ is referred to the teacher model and denoted by $D_{{\mathcal{T}}}=\widehat{D}$ thereafter.
We next design the objective functional for the student model $D_{{\mathcal{S}}}$ to utilize as much of the structure of the problem as possible. First, it is apparent that
$\lim_{s\rightarrow t^{+}}\frac{\int_{t}^{s}\psi(\tau,s)D^{*}(\tau,x(\tau))\,\mathrm{d}\tau}{\int_{t}^{s}\psi(\tau,s)\,\mathrm{d}\tau}=D^{*}(t,x(t)).$
This allows us to reuse the denoiser matching objective functional (4.1) as the local risk to ensure the local consistency of the student model
(4.7) ${\mathcal{R}}_{\operatorname{loc}}(D_{{\mathcal{S}}})=\int_{0}^{T}{\mathbb{E}}\Big{[}\|X_{1}-D_{{\mathcal{S}}}(t,t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t.$
On the other hand, the outputs of generator (4.6) are required to align with the numerical solutions (4.5) and satisfy the semi-group properties, as discussed in Section 2.4. This implies the following global risk
(4.8) ${\mathcal{R}}_{\operatorname{glo}}(D_{{\mathcal{S}}})=\int_{0}^{T}\int_{t}^{T}\int_{s}^{T}{\mathbb{E}}\Big{[}\|g_{s,T}^{\operatorname{off}}\circ g_{u,s}\circ g_{t,u}^{\operatorname{int}}(X_{t})-g_{s,T}^{\operatorname{off}}\circ g_{t,s}^{\operatorname{off}}(X_{t})\|_{2}^{2}\Big{]}{\,\mathrm{d}u\mathrm{d}s\mathrm{d}t},$
where $g$ is induced by the student model $D_{{\mathcal{S}}}$ defined as (4.6), $g^{\operatorname{int}}$ denotes the exponential integrator given by teacher model $D_{{\mathcal{T}}}$, and $g^{\operatorname{off}}$ denotes an offline copy of $g$ for training stability. The population risk (4.8) can be considered as a variant of the original objective functional (2.13), ensuring the long-range consistency of the generator. Combining the short-range denoiser matching risk (4.7) and the long-range characteristic fitting risk (4.8) yields the final training procedure for the characteristic generator. We conclude the practical characteristic learning algorithm in Algorithm 5. The one-step sampling procedure is the same as that in Algorithm 2. For better sampling quality, one can divide the time interval into pieces as Algorithm 4.
Algorithm 5 Practical training procedure of characteristic generator.
1:Observations $X_{1}\sim\mu_{1}$, and pre-trained denoiser $D_{{\mathcal{T}}}$.
2:Initialize the student neural network $D_{{\mathcal{S}},\phi}:{\mathbb{R}}\times{\mathbb{R}}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$.
3:Choose the loss parameter $\lambda$.
4:repeat
5:     # Short-range denoiser matching
6:     Drawn $X_{0}\sim\mu_{0}=N(0,I_{d})$ and $t\sim\operatorname{Unif}[0,T]$.
7:     Construct stochastic interpolant $X_{t}=\alpha_{t}X_{0}+\beta_{t}X_{1}$.
8:     ^$\widehat{{\mathcal{R}}}_{\operatorname{loc}}(D_{{\mathcal{S}},\phi})\leftarrow\|D_{{\mathcal{S}},\phi}(t,t,X_{t})-X_{0}\|_{2}^{2}$.
9:     # Long-range characteristic matching
10:     Drawn $s\sim\operatorname{Unif}[t,T]$ and $u\sim\operatorname{Unif}[s,T]$.
11:     ^$\widehat{{\mathcal{R}}}_{\operatorname{glo}}(D_{{\mathcal{S}},\phi})\leftarrow\|g_{s,T}^{\operatorname{off}}\circ g_{u,s}\circ g_{t,u}^{\operatorname{int}}(X_{t})-g_{s,T}^{\operatorname{off}}\circ g_{t,s}^{\operatorname{off}}(X_{t})\|_{2}^{2}$.
12:     # Combined objective functional
13:     Compute the gradient $\nabla_{\phi}\{\lambda\widehat{{\mathcal{R}}}_{\operatorname{loc}}(D_{{\mathcal{S}},\phi})+\widehat{{\mathcal{R}}}_{\operatorname{glo}}(D_{{\mathcal{S}},\phi})\}$.
14:     Gradient descent update $\phi\leftarrow\phi-\alpha\nabla_{\phi}\{\lambda\widehat{{\mathcal{R}}}_{\operatorname{loc}}(D_{{\mathcal{S}},\phi})+\widehat{{\mathcal{R}}}_{\operatorname{glo}}(D_{{\mathcal{S}},\phi})\}$.
15:until converged
16:Characteristic generator $\widehat{g}(t,s,x)=\varPhi(t,s)x+\varPsi(t,s)D_{{\mathcal{S}},\phi}(t,s,x)$.
Remark 4.1 (Comparison with CTM (Kim et al., 2024)).
Kim et al. (2024) proposed a similar method, but our approach differs from the CTM in two significant aspects. First, the integral scheme $g^{\operatorname{int}}$ employed by CTM is Euler-based. While Euler method coincides with the first-order exponential integrator for VE-ODE (Song et al., 2021c), its numerical stability cannot be guaranteed for general probability flow ODEs. In contrast, the exponential integrator used in our method may ease potential training instability as it fully exploits the semi-linearity of the ODE system. Secondly, CTM relies on GAN training in their approach, borrowing a pre-trained discriminator and treating $g$ as the generator. This reliance on GAN training may cause potential training instability and require extra training of the discriminator. Our method, however, does not require this additional GAN training burden.
4.2. Experiment results and discussions
In this section, we validate the generation quality and sampling efficiency of the characteristic generator using both synthetic and real data through numerical experiments. Föllmer flow is token as the underlying ODE, and all results can be generalized to arbitrary probability flow ODE without loss of generality. We use the Fréchet inception distance (FID) to measure sampling quality on image data. Lower FID means better performance.
4.2.1. Synthetic 2-dimensional dataset
On a 2-dimensional dataset where the target distribution shapes like a Swiss roll, the characteristic generator trained by Algorithm 2 works well. We display the original dataset and samples generated by Euler method (NFE=100) and the characteristic generator (NFE=1) in Figure 2.
Figure 2. Original Swiss roll samples and samples generated by ODE model with Euler method (ODE-Euler) and the characteristic generator (CG).
From Figure 2, it can be observed that the generative quality of the one-step characteristic generator closely resembles that of Euler method with 100 function evaluations (NFE). This indicates that the original training procedure (Algorithm 2) can yield commendable generation outcomes when dealing with relatively simple target distribution.
4.2.2. Real dataset
In this section, we apply the characteristic generator to two real dataset: MNIST and CIFAR-10. The characteristic generators are trained by Algorithm 5.
Generated images by the characteristic generator are displayed in Figure 3 (MNIST on top and CIFAR-10 on bottom). The experimental results illustrate that one-step generation has high generation quality, which can be significantly improved by iterating the characteristic generator by a few steps.
Figure 3. Samples generated by the characteristic generator on CIFAR-10.
We compare the images generated by the numerical sampler and characteristic generator with different numbers of function evaluations (NFE) in Figure 4. The numerical sampler fails to accurately generate images at small NFE values such as 1 and 2. In fact, with 1 NFE, the solution is actually close to the mean of the target distribution. It is necessary to have five or more NFE for the numerical ODE solvers to work properly. In contrast, the characteristic generator is capable of generating high-quality images even with only 1 NFE.
Furthermore, we compare the convergence of FID by NFE for the characteristic generator in Table 3. It is evident that the characteristic generator converges faster than the numerical sampler and achieves better scores at smaller NFE values.
Figure 4. Comparison of samples generated by ODE with exponential integrator (ODE-EI) and characteristic generators (CG) under different NFE on MNIST and CIFAR-10.
Table 3. Comparison of FID by NFE between the exponential integrator (EI) and characteristic generator (CG) in MNIST and CIFAR-10.
Dataset Method NFE=1 NFE=2 NFE=5 NFE=10 NFE=20
MNIST ODE-EI 46.55 46.71 2.69 0.66 0.28
MNIST CG 1.97 1.15 0.28 0.20 0.13
CIFAR-10 ODE-EI 14.06 15.42 5.38 3.16 2.50
CIFAR-10 CG 4.59 3.50 2.90 2.76 2.63
4.2.3. Comparison with other generative models
Table 4 presents the FID on CIFAR-10 achieved by various generative models. Our characteristic generator demonstrates superior generation quality compared to models without GAN, regardless of whether it is one-step or few-step generation. Notably, our proposed method achieves a comparable FID to the state-of-the-art method CTM (Kim et al., 2024), without the requirement of additional GAN training as employed by CTM. Moreover, our proposed characteristic generator with NFE=4 achieves similar or even superior generation performance compared to GAN models.
Table 4. Performance comparisons on CIFAR-10.
Model NFE $\downarrow$ FID $\downarrow$
GAN Models
BigGAN (Brock et al., 2019) 1 8.51
StyleGAN-Ada (Karras et al., 2019) 1 2.92
Diffusion + Sampler
DDPM (Ho et al., 2020) 1000 3.17
DDIM (Song et al., 2021a) 100 4.16
Score SDE (Song et al., 2021c) 2000 2.20
EDM (Karras et al., 2022) 35 2.01
Diffusion + Distillation
KD (Luhman and Luhman, 2021) 1 9.36
DFNO (Zheng et al., 2023a) 1 5.92
Rectified Flow (Liu, 2022) 1 4.85
PD (Salimans and Ho, 2022) 1 9.12
CD (Song et al. (2023), retrained by Kim et al. (2024)) 1 10.53
CTM (without GAN) (Kim et al., 2024, Table 3) 1 5.19
CG (ours) 1 4.59
PD (Salimans and Ho, 2022) 2 4.51
CTM (without GAN) (Kim et al., 2024, Table 3) 18 3.00
CG (ours) 2 3.50
CG (ours) 4 2.83
Diffusion + Distillation + GAN
CD (with GAN) (Lu et al., 2023b) 1 2.65
CTM (with GAN) (Kim et al., 2024) 1 1.98
CTM (with GAN) (Kim et al., 2024) 2 1.87
5. Related Works
5.1. Fast sampling method for diffusion and flow-based models
Diffusion and flow-based models have demonstrated impressive generative performance across various applications. However, their iterative sampling process requires a substantial number of evaluations of the score or velocity neural network, which currently limits their real-time application. In recent years, there has been a surge of fast sampling methods aimed at accelerating the sampling process of diffusion or flow-based models.
The sampling process of the diffusion or flow-based model can be considered as numerically solving SDE or ODE. Therefore, one approach to address this issue is to develop acceleration algorithms for these equations (Zhang and Chen, 2023; Lu et al., 2022, 2023a; Zheng et al., 2023b; Gao and Zhu, 2024; Li et al., 2024a). For instance, Lu et al. (2022) effectively utilizes the semi-linear structure of the probability flow ODE by employing an exponential integrator. Furthermore, this algorithm incorporates adaptive step size schedules and high-order approximations. While these strategies can achieve high-quality generation requiring 10-15 neural network evaluations, generating samples in a single step still poses a significant challenge.
There is another line of recent works that aim to propose a simulation-free one-step sampling method. It is important to note that SDE has probabilistic trajectories, while the trajectory of ODE is deterministic, which is known as “self-consistency” (Song et al., 2023). This line of work is referred to the distillation, which can be divided into two distinct categories.
In the first category, researchers aim to train a deep neural network that maps noise to the endpoint of the probability flow ODE, while disregarding the information at intermediate time points. This class of methods includes knowledge distillation (KD) (Luhman and Luhman, 2021), Euler particle transport (EPT) (Gao et al., 2022), rectified flow (Liu et al., 2022), and diffusion model sampling with neural operator (DSNO) (Zheng et al., 2023a). Unfortunately, these methods are hindered by training instability and low generation quality, as they solely focus on long-range information and are unable to capture the short-range structure at intermediate time points.
The second category of distillation, which is highly relevant to our work, aims to fit the characteristics at each time point using deep neural networks, as demonstrated by Salimans and Ho (2022); Song et al. (2023); Kim et al. (2024); Zhou et al. (2024). These methods take into account both the long-range and short-range structures of the original probability flow, resulting in a high quality of one-step generation. Furthermore, these models exhibit the potential for further enhancement through a few-step iteration process. However, despite their impressive generation quality and training stability, these methods have not yet undergone rigorous theoretical analysis. In contrast, our paper presents a comprehensive framework for generative models utilizing characteristic matching and establishes a rigorous convergence analysis, providing theoretical guarantees for these methods. Additionally, we incorporate the exponential integrator into the characteristic matching procedure. Notably, our characteristic generator surpasses the generation quality achieved by Salimans and Ho (2022); Song et al. (2023); Kim et al. (2024) without the assistance of GANs.
5.2. Error analysis for diffusion and flow-based models
Although a large body of literature has been devoted to the theoretical analysis for diffusion and flow-based generative models, a majority of these works rely on intractable assumptions, such as the regularity of the probability flow SDEs or ODEs. In contrast, our theoretical findings are established under fewer and milder assumptions on the prior and target distribution (Assumptions 1 and 2).
The errors of diffusion-based and flow-based one-step generative models primarily focus on three aspects: velocity matching error, discretization error, and characteristic fitting error. Existing literature on theoretical analysis of these generative models commonly assumes that the $L^{2}$-risk of score or velocity matching is sufficiently small (Lee et al., 2022, 2023; Chen et al., 2023d, c; Benton et al., 2024a, b). However, only a limited number of studies have specifically investigated the convergence rates of score matching (Oko et al., 2023; Chen et al., 2023b; Han et al., 2024) and velocity matching (Chang et al., 2024; Gao et al., 2024; Jiao et al., 2024). The convergence rate of the velocity, as derived in Theorem 3.9, achieves minimax optimality under the assumption of Lipschitz continuity of the target function, which improves upon the rates proposed by Chen et al. (2023b); Chang et al. (2024). The discretization error of the numerical sampler has been explored for both diffusion models (Lee et al., 2022, 2023; Chen et al., 2023d; Benton et al., 2024a), and flow-based methods (Chen et al., 2023c; Gao and Zhu, 2024; Li et al., 2024c, a). To the best of our knowledge, Theorem 3.13 is the first to systematically analyze these three errors, providing theoretical guidance for selecting suitable neural networks and determining the number of numerical discretization steps.
6. Conclusions
In this paper, we propose the characteristic generator, a novel one-step generative model that combines sampling efficiency with high generation quality. In terms of theoretical analysis, we have examined the errors in velocity matching, Euler discretization, and characteristic fitting, enabling us to establish a non-asymptotic convergence rate for the characteristic generator in $2$-Wasserstein distance. This analysis represents the first comprehensive investigation into simulation-free one-step generative models, refining the error analysis of flow-based generative models in prior research. We have validated the effectiveness of our method through experiments on synthetic and real datasets. The results demonstrate that the characteristic generator achieves high generation quality with just a single evaluation of the neural network. This highlights the efficiency and stability of our model in generating high-quality samples.
Finally, we would like to emphasize that our framework of one-step generation is highly versatile and can be extended to conditional generative learning directly. By incorporating the encoder-decoder technique, our approach can be generalized to the latent space, enabling its application in a wide range of practical scenarios, including video generation. The characteristic generator improved by these techniques lays a technical foundation for deploying large-scale generative models on end devices.
On the theoretical front, we intend to exploit the regularity of the velocity field in our error analysis for velocity matching. This will allow us to achieve a faster convergence rate. Additionally, we plan to analyze higher order and more stable numerical schemes, such as the high-order exponential integrator, in order to provide a comprehensive understanding of their effectiveness. Furthermore, we aim to establish a theoretical foundation for the role of semi-group penalties in the characteristic fitting. This will contribute to a deeper understanding of their impact and significance in our framework.
Appendix A Supplemental Definitions and Lemmas
In this section, we introduce some supplemental definitions and lemmas that are used in the proofs.
We first introduce sub-Gaussian random variable [Vershynin, 2018, Wainwright, 2019].
Definition A.1 (Sub-Gaussian).
A random variable $X$ with mean $\mu={\mathbb{E}}[X]$ is sub-Gaussian if there is a positive number $\sigma$ such that
$\log{\mathbb{E}}[\exp(\lambda(X-\mu))]\leq\frac{\sigma^{2}\lambda^{2}}{2},\quad\lambda\in{\mathbb{R}}.$
Here the constant $\sigma$ is referred to as the variance proxy.
The following results (Lemmas A.2 to A.4) shows some important properties of sub-Gaussian variables, whose proofs can be found in Wainwright [2019, Section 2.1].
Lemma A.2 (Chernoff bound).
Let $X$ be a $\sigma^{2}$-sub-Gaussian random variable with zero mean. Then it holds that
$\operatorname{Pr}\big{\{}|X|\geq t\big{\}}\leq 2\exp\Big{(}-\frac{t^{2}}{2\sigma^{2}}\Big{)}.$
Lemma A.3.
Let $X$ be a $\sigma^{2}$-sub-Gaussian random variable with zero mean. Then it holds that
${\mathbb{E}}\Big{[}\exp\Big{(}\frac{\lambda X^{2}}{2\sigma^{2}}\Big{)}\Big{]}\leq\frac{1}{\sqrt{1-\lambda}},\quad\lambda\in[0,1).$
Lemma A.4.
Let $X_{0}$ be a $\sigma_{0}^{2}$-sub-Gaussian, and let $X_{1}$ be a $\sigma_{1}^{2}$-sub-Gaussian independent of $X_{0}$. Then the randon variable $\alpha X_{0}+\beta X_{1}$ is sub-Gaussian with variance proxy $\alpha^{2}\sigma_{0}^{2}+\beta^{2}\sigma_{1}^{2}$.
Lemmas A.5 and A.6 show bounds of the tail probability and the expectation of the maximum of $N$ sub-Gaussian variables.
Lemma A.5.
Let $\{X_{n}\}_{n=1}^{N}$ be a set of $\sigma^{2}$-sub-Gaussian random variables with zero mean, then it follows that
$\operatorname{Pr}\Big{\{}\max_{1\leq n\leq N}|X_{n}|\geq t\Big{\}}\leq 2N\exp\Big{(}-\frac{t^{2}}{2\sigma^{2}}\Big{)}.$
The following lemma states that the expectation of the maximum of the squares of $N$ sub-Gaussian variables is bounded by $\log N$.
Lemma A.6.
Let $\{X_{n}\}_{n=1}^{N}$ be a set of $\sigma^{2}$-sub-Gaussian random variables with zero mean, then it follows that
${\mathbb{E}}\Big{[}\max_{1\leq n\leq N}X_{n}^{2}\Big{]}\leq 4\sigma^{2}(\log N+1).$
Proof of Lemma A.5.
It is straightforward that
$\operatorname{Pr}\Big{\{}\max_{1\leq n\leq N}|X_{n}|\geq t\Big{\}}\leq\sum_{n=1}^{N}\operatorname{Pr}\big{\{}|X_{n}|\geq t\big{\}}\leq 2N\exp\Big{(}-\frac{t^{2}}{2\sigma^{2}}\Big{)},$
where the last inequality holds from Lemma A.2. This completes the proof. ∎
Proof of Lemma A.6.
By Jensen’s inequality, it is straightforward that
$\exp\Big{(}\frac{\lambda}{2\sigma^{2}}{\mathbb{E}}\Big{[}\max_{1\leq n\leq N}\xi_{n}^{2}\Big{]}\Big{)}\leq{\mathbb{E}}\Big{[}\max_{1\leq n\leq N}\exp\Big{(}\frac{\lambda\xi_{n}^{2}}{2\sigma^{2}}\Big{)}\Big{]}\leq N{\mathbb{E}}\Big{[}\exp\Big{(}\frac{\lambda\xi_{1}^{2}}{2\sigma^{2}}\Big{)}\Big{]}\leq\frac{N}{\sqrt{1-\lambda}},$
where the last inequality holds from Lemma A.3 for each $\lambda\in[0,1)$. Letting $\lambda=1/2$ yields the desired inequality. ∎
Lemma A.7 (Fourth moment of standard Gaussian).
Let $\epsilon\sim N(0,I_{d})$. Then ${\mathbb{E}}[\|\epsilon\|_{2}^{4}]=d^{2}+2d$.
Proof of Lemma A.7.
It is straightforward that
${\mathbb{E}}[\|\epsilon\|_{2}^{4}]={\mathbb{E}}\Big{[}\sum_{k=1}^{d}\epsilon_{k}^{4}+\sum_{k\neq\ell}\epsilon_{k}^{2}\epsilon_{\ell}^{2}\Big{]}=\sum_{k=1}^{d}{\mathbb{E}}[\epsilon_{k}^{4}]+\sum_{k\neq\ell}{\mathbb{E}}[\epsilon_{k}^{2}]{\mathbb{E}}[\epsilon_{\ell}^{2}]=d^{2}+2d,$
where we used the fact that ${\mathbb{E}}[X^{4}]=3$ for $X\sim N(0,1)$. ∎
We next introduce the notation of covering number and Vapnik-Chervonenkis dimension (VC-dimension), both of which measure the complexity of a function class [Mohri et al., 2018, Vaart and Wellner, 2023]. They are used in the error analysis for velocity matching (Section D) and characteristic fitting (Section F).
Definition A.8 (Covering number).
Let ${\mathscr{F}}$ be a class of functions mapping from ${\mathbb{R}}^{d}$ to ${\mathbb{R}}$. Suppose ${\mathcal{D}}=\{X^{(i)}\}_{i=1}^{n}$ is a set of samples in ${\mathbb{R}}^{d}$. Define the $L^{\infty}({\mathcal{D}})$-norm of the function $f\in{\mathscr{F}}$ as $\|f\|_{L^{\infty}({\mathcal{D}})}=\max_{1\leq i\leq n}|f(X^{(i)})|$. A function set ${\mathscr{F}}_{\delta}$ is called an $L^{\infty}({\mathcal{D}})$ $\delta$-cover of ${\mathscr{F}}$ if for each $f\in{\mathscr{F}}$, there exits $f_{\delta}\in{\mathscr{F}}_{\delta}$ such that $\|f-f_{\delta}\|_{L^{\infty}({\mathcal{D}})}\leq\delta$. Furthermore,
$N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))=\inf\Big{\{}|{\mathscr{F}}_{\delta}|:{\mathscr{F}}_{\delta}\text{ is a $L^{\infty}({\mathcal{D}})$ $\delta$-cover of ${\mathscr{F}}$}\Big{\}}$
is called the $L^{\infty}({\mathcal{D}})$ $\delta$-covering number of ${\mathscr{F}}$.
Definition A.9 (VC-dimension).
Let ${\mathscr{F}}$ be a class of functions from ${\mathbb{R}}^{d}$ to {±$\{\pm 1\}$. For any non-negative integer $m$, we define the growth function of ${\mathscr{F}}$ as
$\Pi_{{\mathscr{F}}}(m)=\max_{\{X^{(i)}\}_{i=1}^{m}\subseteq{\mathbb{R}}^{d}}\big{|}\{(f(X^{(1)}),\ldots,f(X^{(m)})):f\in{\mathscr{F}}\}\big{|}.$
A set $\{X^{(i)}\}_{i=1}^{m}$ is said to be shattered by ${\mathscr{F}}$ when |$|\{(f(X^{(1)}),\ldots,f(X^{(m)})):f\in{\mathscr{F}}\}|=2^{m}$. The Vapnik-Chervonenkis dimension of ${\mathscr{F}}$, denoted $\operatorname{VCdim}({\mathscr{F}})$, is the size of the largest set that can be shattered by ${\mathscr{F}}$, that is, $\operatorname{VCdim}({\mathscr{F}})=\max\{m:\Pi_{{\mathscr{F}}}(m)=2^{m}\}$. For a class ${\mathscr{F}}$ of real-valued functions, we define $\operatorname{VCdim}({\mathscr{F}})=\operatorname{VCdim}(\operatorname{sign}({\mathscr{F}}))$.
Lemma A.10.
Let ${\mathscr{F}}$ be a class of functions mapping from ${\mathbb{R}}^{d}$ to ${\mathbb{R}}$, and let ${\mathscr{H}}$ be a function class defined as =${\mathscr{H}}=\{(x,f)\mapsto h(f,x)\in{\mathbb{R}}:f\in{\mathscr{F}}\}$. Suppose ${\mathcal{D}}=\{X^{(i)}\}_{i=1}^{n}$ is a set of samples in ${\mathbb{R}}^{d}$. If there exists a constant $L>0$ such that for each $f,f^{\prime}\in{\mathscr{F}}$,
$\max_{1\leq i\leq n}|h(f,X^{(i)})-h(f^{\prime},X^{(i)})|\leq L\max_{1\leq i\leq n}|f(X^{(i)})-f^{\prime}(X^{(i)})|,$
then the following inequality holds for each $\delta>0$,
$N(L\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))\leq N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}})).$
Proof of Lemma A.10.
Let ${\mathscr{F}}_{\delta}$ be an $L^{\infty}({\mathcal{D}})$ $\delta$-cover of ${\mathscr{F}}$ with $|{\mathscr{F}}_{\delta}|=N(L\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))$. Define ${\mathscr{H}}_{\delta}=\{(x,f)\mapsto h(f,x)\in{\mathbb{R}}:f\in{\mathscr{F}}_{\delta}\}$. Then for each $h(f,\cdot)\in{\mathscr{H}}$, there exists $h(f_{\delta},\cdot)\in{\mathscr{H}}_{\delta}$, such that
$\max_{1\leq i\leq n}|h(f,X^{(i)})-h(f_{\delta},X^{(i)})|\leq L\max_{1\leq i\leq n}|f(X^{(i)})-f_{\delta}(X^{(i)})|\leq L\delta.$
Thus ${\mathscr{H}}_{\delta}$ is an $L^{\infty}({\mathcal{D}})$ $(L\delta)$-cover of ${\mathscr{H}}$. This completes the proof. ∎
We then bound the covering number by VC-dimension as following lemma.
Lemma A.11 (Anthony et al. [1999, Theorem 12.2]).
Let ${\mathscr{F}}$ be a class of functions mapping from ${\mathbb{R}}^{d}$ to $[0,B]$. Then it follows that for each $n\geq\operatorname{VCdim}({\mathscr{F}})$,
$\sup_{{\mathcal{D}}\in({\mathbb{R}}^{d})^{n}}\log N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))\leq\operatorname{VCdim}({\mathscr{F}})\log\Big{(}\frac{enB}{\delta\operatorname{VCdim}({\mathscr{F}})}\Big{)}.$
The following lemma provides a VC-dimension bound for neural network classes with piecewise-polynomial activation functions.
Lemma A.12 (Bartlett et al. [2019, Theorem 7]).
The VC-dimension of a neural network class with piecewise-polynomial activation functions is bounded as follows
$\operatorname{VCdim}(N(L,S))\leq CLS\log(S),$
where $C$ is an absolute constant.
Appendix B Proof of Results in Section 2
The proofs of Propositions 2.1 and 2.3 are shown in this section.
The proof of Proposition 2.1 follows from the proof of Albergo et al. [2023b, Theorem 2] and Albergo et al. [2023a, Theorem 2.6].
Proof of Proposition 2.1.
The characteristic function of $X_{t}$ (2.1) is given as
$\varphi_{X_{t}}(\xi)$ $={\mathbb{E}}_{X_{t}}[\exp(i\xi\cdot X_{t})]$
$=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\exp(i\xi\cdot(\alpha_{t}x_{0}+\beta_{t}x_{1}))\rho_{0}(x_{0})\rho_{1}(x_{1})\,\mathrm{d}x_{0}\,\mathrm{d}x_{1}$
$=\int_{{\mathbb{R}}^{d}}\exp(i\xi\cdot(\alpha_{t}x_{0}))\rho_{0}(x_{0})\,\mathrm{d}x_{0}\int_{{\mathbb{R}}^{d}}\exp(i\xi\cdot(\beta_{t}x_{1}))\rho_{1}(x_{1})\,\mathrm{d}x_{1}$
$={\mathbb{E}}_{X_{0}}[\exp(i\alpha_{t}\xi\cdot X_{0})]{\mathbb{E}}_{X_{1}}[\exp(i\beta_{t}\xi\cdot X_{1})]=\varphi_{X_{0}}(\alpha_{t}\xi)\varphi_{X_{1}}(\beta_{t}\xi),$
where $\xi\in{\mathbb{R}}^{d}$, and we used the fact that $X_{0}$ is independent of $X_{1}$. On the other hand,
$\int_{{\mathbb{R}}^{d}}\exp(i\xi\cdot x)\Big{(}\frac{1}{\beta_{t}}\int_{{\mathbb{R}}^{d}}\rho_{0}(x_{0})\rho_{1}\Big{(}\frac{x-\alpha_{t}x_{0}}{\beta_{t}}\Big{)}\,\mathrm{d}x_{0}\Big{)}\,\mathrm{d}x$
$=\frac{1}{\beta_{t}}\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\exp(i\alpha_{t}\xi\cdot x_{0})\exp(i\xi\cdot(x-\alpha_{t}x_{0}))\rho_{0}(x_{0})\rho_{1}\Big{(}\frac{x-\alpha_{t}x_{0}}{\beta_{t}}\Big{)}\,\mathrm{d}x_{0}\,\mathrm{d}x$
$=\int_{{\mathbb{R}}^{d}}\exp(i\alpha_{t}\xi\cdot x_{0})\rho_{0}(x_{0})\Big{\{}\int_{{\mathbb{R}}^{d}}\exp\Big{(}i\xi\cdot(x-\alpha_{t}x_{0})\Big{)}\rho_{1}\Big{(}\frac{x-\alpha_{t}x_{0}}{\beta_{t}}\Big{)}\,\mathrm{d}\Big{(}\frac{x-\alpha_{t}x_{0}}{\beta_{t}}\Big{)}\Big{\}}\,\mathrm{d}x_{0}$
$=\int_{{\mathbb{R}}^{d}}\exp(i\alpha_{t}\xi\cdot x_{0})\rho_{0}(x_{0})\Big{\{}\int_{{\mathbb{R}}^{d}}\exp(i\beta_{t}\xi\cdot x_{1})\rho_{1}(x_{1})\,\mathrm{d}x_{1}\Big{\}}\,\mathrm{d}x_{0}=\varphi_{X_{0}}(\alpha_{t}\xi)\varphi_{X_{1}}(\beta_{t}\xi).$
Combining the above two equality deduces that for each $\xi\in{\mathbb{R}}^{d}$,
$\varphi_{X_{t}}(\xi)=\int_{{\mathbb{R}}^{d}}\exp(i\xi\cdot x)\Big{(}\frac{1}{\beta_{t}}\int_{{\mathbb{R}}^{d}}\rho_{0}(x_{0})\rho_{1}\Big{(}\frac{x-\alpha_{t}x_{0}}{\beta_{t}}\Big{)}\,\mathrm{d}x_{0}\Big{)}\,\mathrm{d}x.$
By using Fourier inversion theorem, we obtain the density function of $X_{t}$ as
$\rho_{t}(x)=\frac{1}{\beta_{t}}\int_{{\mathbb{R}}^{d}}\rho_{0}(x_{0})\rho_{1}\Big{(}\frac{x-\alpha_{t}x_{0}}{\beta_{t}}\Big{)}\,\mathrm{d}x_{0}.$
By a same argument, we have
$\rho_{t}(x)=\frac{1}{\alpha_{t}}\int_{{\mathbb{R}}^{d}}\rho_{0}\Big{(}\frac{x-\beta_{t}x_{1}}{\alpha_{t}}\Big{)}\rho_{1}(x_{1})\,\mathrm{d}x_{1}.$
We next turn to verify that the density function solves the transport equation. Let $\phi$ be an arbitrary smooth testing function. By the definition of the interpolant (2.1), it holds that
${\mathbb{E}}_{X_{t}}\big{[}\phi(X_{t})\big{]}={\mathbb{E}}_{(X_{0},X_{1})}\big{[}\phi(\alpha_{t}X_{0}+\beta_{t}X_{1})\big{]}.$
Taking derivative with respect to $t$ on the left-hand side of the equality yields
(B.1) $\frac{\partial}{\partial t}\Big{(}\int_{{\mathbb{R}}^{d}}\phi(x)\rho_{t}(x)\,\mathrm{d}x\Big{)}=\int_{{\mathbb{R}}^{d}}\phi(x)\partial_{t}\rho_{t}(x)\,\mathrm{d}x.$
Similarly, for the right-hand side of the equality, we have
$\frac{\partial}{\partial t}\Big{(}\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\phi(\alpha_{t}x_{0}+\beta_{t}x_{2})\rho_{0}(x_{0})\rho_{1}(x_{1})\,\mathrm{d}x_{0}\,\mathrm{d}x_{1}\Big{)}$
$=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}(\dot{\alpha}_{t}x_{0}+\dot{\beta}_{t}x_{1})\cdot\nabla\phi(\alpha_{t}x_{0}+\beta_{t}x_{1})\rho_{0}(x_{0})\rho_{1}(x_{1})\,\mathrm{d}x_{0}\,\mathrm{d}x_{1}$
$=\int_{{\mathbb{R}}^{d}}{\mathbb{E}}[\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}X_{1}|X_{t}=x]\cdot\nabla\phi(x)\rho_{t}(x)\,\mathrm{d}x$
(B.2) $=\int_{{\mathbb{R}}^{d}}b_{t}^{*}(x)\cdot\nabla\phi(x)\rho_{t}(x)\,\mathrm{d}x=-\int_{{\mathbb{R}}^{d}}\phi(x)\nabla\cdot(b_{t}^{*}(x)\rho_{t}(x))\,\mathrm{d}x,$
where the first equality follows from the chain rule, the forth equality holds from the definition of $b_{t}^{*}$ (2.3), and the last equality is due to integration by parts and the divergence theorem [Evans, 2010, Theorem 1 in Section C.2]. Combining (B.1) and (B.2) gives the desired transport equation. ∎
Proof of Proposition 2.3.
See Lemma D.3 and its proof for details. ∎
Appendix C Properties of the Probability Flow ODE
In this section, we present some auxiliary properties of the probability flow ODE in Section C.1. Proofs of the propositions in Section 3.2 are given in Section C.2.
C.1. Auxiliary Properties
Lemmas C.1 to C.4 are established under Assumption 1.
Lemma C.1.
Suppose Assumption 1 holds. Then the conditional score function is given as
$\nabla_{x}\log\rho_{t|1}(x|x_{1})=-\frac{1}{\alpha_{t}^{2}}x+\frac{\beta_{t}}{\alpha_{t}^{2}}x_{1},\quad t\in(0,1).$
Proof of Lemma C.1.
Given $X_{1}=x_{1}$, using the definition of stochastic interpolant (2.1) implies
$(X_{t}|X_{1}=x_{1})\sim N(\beta_{t}x_{1},\alpha_{t}^{2}I_{d}),\quad t\in(0,1),$
which implies the desired result immediately. ∎
Lemma C.2.
Suppose Assumption 1 holds. Then the following holds
${\mathbb{E}}[X_{1}|X_{t}=x]=\frac{1}{\beta_{t}}x+\frac{\alpha_{t}^{2}}{\beta_{t}}\nabla_{x}\log\rho_{t}(x).$
Proof of Lemma C.2.
It is straightforward that
$\nabla_{x}\log\rho_{t}(x)$ $=\frac{\nabla_{x}\rho_{t}(x)}{\rho_{t}(x)}=\frac{1}{\rho_{t}(x)}\nabla_{x}\Big{(}\int_{{\mathbb{R}}^{d}}\rho_{t|1}(x|x_{1})\rho_{1}(x_{1})\,\mathrm{d}x_{1}\Big{)}$
$=\int_{{\mathbb{R}}^{d}}\frac{\rho_{t|1}(x|x_{1})\rho_{1}(x_{1})}{\rho_{t}(x)}\nabla_{x}\log\rho_{t|1}(x|x_{1})\,\mathrm{d}x_{1}$
$=\int_{{\mathbb{R}}^{d}}\rho_{1|t}(x_{1}|x)\Big{(}-\frac{1}{\alpha_{t}^{2}}x+\frac{\beta_{t}}{\alpha_{t}^{2}}x_{1}\Big{)}\,\mathrm{d}x_{1}$
$=-\frac{1}{\alpha_{t}^{2}}x+\frac{\beta_{t}}{\alpha_{t}^{2}}{\mathbb{E}}[X_{1}|X_{t}=x],$
where the we used the definition of conditional density and Lemma C.1. This completes the proof. ∎
Lemma C.3.
Suppose Assumptions 1 holds. Then it follows that
$\nabla b^{*}(t,x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}I_{d}+\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\frac{\beta_{t}^{2}}{\alpha_{t}^{2}}\operatorname{Cov}(X_{1}|X_{t}=x).$
Proof of Lemma C.3.
According to the proof of Proposition 3.1, we have
$b^{*}(t,x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}{\mathbb{E}}[X_{1}|X_{t}=x],$
which deduces
(C.1) $\nabla_{x}b^{*}(t,x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}I_{d}+\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\nabla_{x}{\mathbb{E}}[X_{1}|X_{t}=x].$
Hence it suffices to estimate the gradient of the conditional expectation. In fact,
$\nabla_{x}{\mathbb{E}}[X_{1}|X_{t}=x]=\nabla_{x}\Big{(}\frac{1}{{\rho_{t}(x)}}\int_{{\mathbb{R}}^{d}}x_{1}\rho_{t|1}(x|x_{1})\rho_{1}(x_{1})\,\mathrm{d}x_{1}\Big{)}$
$=-\frac{\nabla_{x}\rho_{t}(x)}{{\rho_{t}^{2}(x)}}\int_{{\mathbb{R}}^{d}}x_{1}^{T}\rho_{t|1}(x|x_{1})\rho_{1}(x_{1})\,\mathrm{d}x_{1}+\frac{1}{{\rho_{t}(x)}}\int_{{\mathbb{R}}^{d}}\nabla_{x}\rho_{t|1}(x|x_{1})x_{1}^{T}\rho_{1}(x_{1})\,\mathrm{d}x_{1}$
(C.2) $=-\nabla_{x}\log\rho_{t}(x)\int_{{\mathbb{R}}^{d}}x_{1}^{T}\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1}+\int_{{\mathbb{R}}^{d}}\nabla_{x}\log\rho_{t|1}(x|x_{1})x_{1}^{T}\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1},$
where we used the definition of the conditional density that
$\rho_{1|t}(x_{1}|x)=\frac{\rho_{t|1}(x|x_{1})\rho_{1}(x_{1})}{\rho_{t}(x)}.$
For the first term in the right-hand side of (C.2), applying Lemma C.2 implies
$-\nabla_{x}\log\rho_{t}(x)\int_{{\mathbb{R}}^{d}}x_{1}^{T}\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1}$
(C.3) $=\Big{(}\frac{1}{\alpha_{t}^{2}}x-\frac{\beta_{t}}{\alpha_{t}^{2}}{\mathbb{E}}[X_{1}|X_{t}=x]\Big{)}{\mathbb{E}}[X_{1}|X_{t}=x]^{T}.$
For the second term, it follows from Lemma C.1 that
$\int_{{\mathbb{R}}^{d}}\log\nabla_{x}\rho_{t|1}(x|x_{1})x_{1}^{T}\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1}$
(C.4) $=-\frac{1}{\alpha_{t}^{2}}x{\mathbb{E}}[X_{1}|X_{t}=x]^{T}+\frac{\beta_{t}}{\alpha_{t}^{2}}{\mathbb{E}}[X_{1}X_{1}^{T}|X_{t}=x].$
Plugging (C.3) and (C.4) into (C.2) yields
(C.5) $\nabla{\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\beta_{t}}{\alpha_{t}^{2}}\operatorname{Cov}(X_{1}|X_{t}=x).$
Substituting (C.5) into (C.1) completes the proof. ∎
The following lemma gives an explicit expression of time derivative of the velocity.
Lemma C.4 (Gao et al. [2023, Proposition 63]).
Suppose Assumption 1 holds. Then it follows that for each $(t,x)\in(0,1)\times{\mathbb{R}}^{d}$,
$\partial_{t}b^{*}(t,x)$ $=\Big{(}\frac{\ddot{\alpha}_{t}}{\alpha_{t}}-\frac{\dot{\alpha}_{t}^{2}}{\alpha_{t}^{2}}\Big{)}x+\Big{(}\alpha_{t}^{2}\frac{\ddot{\beta}_{t}}{\beta_{t}}-\dot{\alpha}_{t}\alpha_{t}\frac{\dot{\beta}_{t}}{\beta_{t}}-\ddot{\alpha}_{t}\alpha_{t}+\dot{\alpha}_{t}^{2}\Big{)}\frac{\beta_{t}}{\alpha_{t}^{2}}{\mathbb{E}}[X_{1}|X_{t}=x]$
$\quad+\frac{\beta_{t}^{2}}{\alpha_{t}^{2}}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-2\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\operatorname{Cov}(X_{1}|X_{t}=x)x$
$\quad-\frac{\beta_{t}^{3}}{\alpha_{t}^{2}}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}^{2}\Big{(}{\mathbb{E}}[X_{1}X_{1}^{T}X_{1}|X_{t}=x]-{\mathbb{E}}[X_{1}X_{1}^{T}|X_{t}=x]{\mathbb{E}}[X_{1}|X_{t}=x]\Big{)}.$
Lemmas C.5 to C.8 are established under Assumptions 1 and 2.
Lemma C.5.
Suppose Assumptions 1 and 2 hold. Then it follows that
$(X_{1}|X_{t}=x)\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U+\sqrt{\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x.$
Proof of Lemma C.5.
According to the definition of stochastic interpolant (2.1), we have
(C.6) $\rho_{t|1}(x|x_{1})=\frac{1}{(2\pi)^{d/2}\alpha_{t}}\exp\Big{(}-\frac{1}{\alpha_{t}^{2}}\|\beta_{t}x_{1}-x\|_{2}^{2}\Big{)}.$
Using Assumption 2, the target distribution is given as
(C.7) $\rho_{1}(x_{1})=\frac{1}{(2\pi)^{d/2}\sigma}\int_{{\mathbb{R}}^{d}}\exp\Big{(}-\frac{1}{\sigma^{2}}\|x_{1}-u\|_{2}^{2}\Big{)}\,\mathrm{d}\nu(u).$
Combining (C.6) and (C.7) implies
$\rho_{1|t}(x_{1}|x)=\frac{\rho_{t|1}(x|x_{1})\rho_{1}(x_{1})}{\rho_{t}(x)}$
$=\frac{1}{(2\pi)^{d}\sigma\alpha_{t}}\frac{1}{\rho_{t}(x)}\int_{{\mathbb{R}}^{d}}\exp\Big{(}-\frac{1}{\alpha_{t}^{2}}\|\beta_{t}x_{1}-x\|_{2}^{2}\Big{)}\exp\Big{(}-\frac{1}{\sigma^{2}}\|x_{1}-u\|_{2}^{2}\Big{)}\,\mathrm{d}\nu(u)$
$=\frac{1}{(2\pi)^{d/2}}\sqrt{\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\int_{{\mathbb{R}}^{d}}\exp\Big{(}-\frac{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\sigma^{2}\alpha_{t}^{2}}\Big{\|}x_{1}-\frac{\sigma^{2}\beta_{t}x+\alpha_{t}^{2}u}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{\|}_{2}^{2}\Big{)}\,\mathrm{d}\nu(u),$
which can be viewed as a convolution of $\nu$ and a Gaussian distribution. Therefore, we can construct a random variable which obeys the conditional density $\rho_{1|t}(x_{1}|x)$ as
$(X_{1}|X_{t}=x)\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U+\sqrt{\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
where $U$ and $\epsilon$ are two independent random variables sample from $\nu$ and $N(0,I_{d})$, respectively. This completes the proof. ∎
Lemma C.6 (Conditional expectation).
Suppose Assumptions 1 and 2 hold. Then the following inequalities hold for each $t\in(0,1)$ and $x\in{\mathbb{R}}^{d}$,
$|{\mathbb{E}}[X_{1,k}|X_{t}=x]|\leq M(1+|x_{k}|),\quad 1\leq k\leq d,$
where $M$ is a constant only depending on $d$ and $\sigma$.
Proof of Lemma C.6.
According to Lemma C.5, it holds that
${\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
which implies the desired inequalities directly. ∎
Lemma C.7 (Conditional covariance).
Suppose Assumptions 1 and 2 hold. Then the following inequalities hold for each $t\in(0,1)$ and $x\in{\mathbb{R}}^{d}$,
$\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}\preceq\operatorname{Cov}(X_{1}|X_{t}=x)\preceq\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}+d\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}I_{d}.$
Proof of Lemma C.7.
It is straightforward from Lemma C.5 that
${\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
which implies
${\mathbb{E}}[X_{1}|X_{t}=x]{\mathbb{E}}[X_{1}|X_{t}=x]^{T}$
$=\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}{\mathbb{E}}[U]{\mathbb{E}}[U]^{T}+\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]x^{T}+\Big{(}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}xx^{T}.$
On the other hand, using Lemma C.5 deduces
${\mathbb{E}}[X_{1}X_{1}^{T}|X_{t}=x]$ $=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]x^{T}+\Big{(}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}xx^{T}$
$\quad+\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}{\mathbb{E}}[UU^{T}]+\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
Combining the above two equalities yields
$\operatorname{Cov}(X_{1}|X_{t}=x)$ $={\mathbb{E}}[X_{1}X_{1}^{T}|X_{t}=x]-{\mathbb{E}}[X_{1}|X_{t}=x]{\mathbb{E}}[X_{1}|X_{t}=x]^{T}$
$=\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}\operatorname{Cov}(U)+\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
According to Lemma C.5, the random variable $\|U\|_{\infty}\leq 1$ and thus $\operatorname{Cov}(U)\preceq dI_{d}$. Consequently,
$\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}\preceq\operatorname{Cov}(X_{1}|X_{t}=x)\preceq\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}+d\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}I_{d},$
for each $(t,x)\in(0,1)\times{\mathbb{R}}^{d}$. This completes the proof. ∎
By a same argument as Lemmas C.6 and C.7, we also have the following inequality. See Gao et al. [2024, Lemma A.8] for a detailed proof.
Lemma C.8.
Suppose Assumptions 1 and 2 hold. Then the following inequalities hold for each $t\in(0,1)$ and $x\in{\mathbb{R}}^{d}$,
$\|{\mathbb{E}}[X_{1}X_{1}^{T}X_{1}|X_{t}=x]-{\mathbb{E}}[X_{1}^{T}X_{1}|X_{t}=x]{\mathbb{E}}[X_{1}|X_{t}]\|_{2}\leq M\alpha_{t}^{2}(1+\|x\|_{2}),$
where $M$ is a constant only depending on $d$ and $\sigma$.
C.2. Proof of Results in Section 3.2
Then we show proofs of propositions and corollaries in Section 3.2.
Proof of Proposition 3.1.
Using the definition of velocity (2.3), we have
$b^{*}(t,x)$ $={\mathbb{E}}[\dot{\alpha}_{1}X_{0}+\dot{\beta}_{t}X_{1}|X_{t}=x]={\mathbb{E}}\Big{[}\dot{\beta}_{t}X_{1}+\frac{\dot{\alpha}_{t}}{\alpha_{t}}(X_{t}-\beta_{t}X_{1})\Big{|}X_{t}=x\Big{]}$
$=\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}{\mathbb{E}}[X_{1}|X_{t}=x]+\frac{\dot{\alpha}_{t}}{\alpha_{t}}x=\alpha_{t}^{2}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\nabla_{x}\log\rho_{t}(x)+\frac{\dot{\beta}_{t}}{\beta_{t}}x,$
where the second equality holds from the definition of stochastic interpolant (2.1), and the last equality is due to Lemma C.2. This completes the proof. ∎
Proof of Proposition 3.2.
Using the definition of velocity (2.3), we have
$b^{*}(t,x)$ $={\mathbb{E}}[\dot{\alpha}_{1}X_{0}+\dot{\beta}_{t}X_{1}|X_{t}=x]={\mathbb{E}}\Big{[}\dot{\beta}_{t}X_{1}+\frac{\dot{\alpha}_{t}}{\alpha_{t}}(X_{t}-\beta_{t}X_{1})\Big{|}X_{t}=x\Big{]}$
$=\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}{\mathbb{E}}[X_{1}|X_{t}=x]+\frac{\dot{\alpha}_{t}}{\alpha_{t}}x$
$=\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x\Big{)}+\frac{\dot{\alpha}_{t}}{\alpha_{t}}x$
$=\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x$
where the second equality holds from the definition of stochastic interpolant (2.1), and the fourth equality follows from Lemma C.6. This completes the proof. ∎
Proof of Corollary 3.3.
For each $1\leq k\leq d$, it follows from (2.4) that
$x_{k}(s)=x_{k}(t)+\int_{t}^{s}b_{k}^{*}(\tau,x(\tau))\,\mathrm{d}\tau.$
Using the triangular inequality and Jensen’s inequality, we have
$|x_{k}(s)|$ $\leq|x_{k}(t)|+\int_{t}^{s}|b_{k}^{*}(\tau,x(\tau))|\,\mathrm{d}\tau$
$\leq|x_{k}(t)|+\int_{t}^{s}\Big{|}\frac{\alpha_{\tau}(\alpha_{\tau}\dot{\beta}_{\tau}-\dot{\alpha}_{\tau}\beta_{\tau})}{\alpha_{\tau}^{2}+\sigma^{2}\beta_{\tau}^{2}}\Big{|}\,\mathrm{d}\tau+\int_{t}^{s}\Big{|}\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{|}|x_{k}(\tau)|\,\mathrm{d}\tau,$
where the second inequality follows from Lemma C.6, and $M$ is a constant only depending on $d$ and $\sigma$. Applying Gronwall’s inequality [Evans, 2010, Section B.2] yields
$|x_{k}(s)|\leq C\exp\Big{(}\int_{t}^{s}\Big{|}\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{|}\,\mathrm{d}\tau\Big{)}\Big{(}|x_{k}(t)|+\int_{t}^{s}\Big{|}\frac{\alpha_{\tau}(\alpha_{\tau}\dot{\beta}_{\tau}-\dot{\alpha}_{\tau}\beta_{\tau})}{\alpha_{\tau}^{2}+\sigma^{2}\beta_{\tau}^{2}}\Big{|}\,\mathrm{d}\tau\Big{)},$
where the constant $C$ depends only on $d$ and $\sigma$. This completes the proof. ∎
Proof of Corollary 3.4.
It follows from the definition of the probability flow $g^{*}$ (2.4) that
$\partial_{t}g^{*}(t,s,x)=-b^{*}(t,x)\quad\text{and}\quad\partial_{s}g^{*}(t,s,x)=b^{*}(s,x).$
Then we obtain the desired result according to Proposition 3.2. ∎
Proof of Proposition 3.5.
It sufficient to show that
(C.8) $\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}\preceq\nabla b^{*}(t,x)\preceq\Big{(}\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}+d\frac{\alpha_{t}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}\Big{)}I_{d}.$
For the upper bound, it follows from Lemma C.3 and Lemma C.7 that
$\nabla b^{*}(t,x)$ $\preceq\frac{\dot{\alpha}_{t}}{\alpha_{t}}I_{d}+\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\frac{\beta_{t}^{2}}{\alpha_{t}^{2}}\Big{(}\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}+d\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}I_{d}\Big{)}$
(C.9) $=\Big{(}\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}+d\frac{\alpha_{t}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}\Big{)}I_{d}.$
By a similar argument, we have
(C.10) $\nabla b^{*}(t,x)\succeq\frac{\dot{\alpha}_{t}}{\alpha_{t}}I_{d}+\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\frac{\beta_{t}^{2}}{\alpha_{t}^{2}}\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}=\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
Combining (C.9) and (C.10) yields (C.8). This competes the proof. ∎
With the aid of above auxiliary lemmas, we provide the following proof of Corollary 3.6.
Proof of Corollary 3.6.
We first show the Lipschitz continuity of the velocity. It is straightforward that for each $t\in(0,1)$ and $x,x^{\prime}\in{\mathbb{R}}^{d}$,
$\|b^{*}(t,x)-b^{*}(t,x^{\prime})\|_{2}$ $=\Big{\|}\int_{0}^{1}\frac{\,\mathrm{d}}{\,\mathrm{d}\tau}b^{*}(t,x^{\prime}+\tau(x-x^{\prime}))\,\mathrm{d}\tau\Big{\|}_{2}$
$=\Big{\|}\int_{0}^{1}\nabla b^{*}(t,x^{\prime}+\tau(x-x^{\prime}))\,\mathrm{d}\tau(x-x^{\prime})\Big{\|}_{2}$
$\leq\int_{0}^{1}\|\nabla b^{*}(t,x^{\prime}+\tau(x-x^{\prime}))\|_{\operatorname{op}}\,\mathrm{d}\tau\|x-x^{\prime}\|_{2}$
(C.11) $\leq\Big{(}\int_{0}^{1}G\,\mathrm{d}\tau\Big{)}\|x-x^{\prime}\|_{2}=G\|x-x^{\prime}\|_{2},$
where the first inequality follows from the definition of the operator norm $\|\cdot\|_{\operatorname{op}}$ and Jensen’s inequality, and the second inequality is due to Proposition 3.5. This shows the Lipschitz continuity of the velocity.
We now turn to focus on the Lipschitz continuity of the flow. It suffices to show that the solution at time $s$ depends Lipschitz continuously on the solution $x_{t}$ at time $t$. Let $x(\cdot)$ and $x^{\prime}(\cdot)$ be two continuous vector-valued functions satisfy the ODE (2.4) with different values at $t$. Then it follows that
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}(x_{k}(t)-x_{k}^{\prime}(t))=b_{k}^{*}(t,x(t))-b_{k}^{*}(t,x^{\prime}(t)),\quad 1\leq k\leq d.$
By the definition of $\ell_{2}$-norm, we have
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|x(t)-x^{\prime}(t)\|_{2}$ $=\frac{1}{2\|x(t)-x^{\prime}(t)\|_{2}}\sum_{k=1}^{d}\frac{\,\mathrm{d}}{\,\mathrm{d}t}(x_{k}(t)-x_{k}^{\prime}(t))^{2}$
$=\frac{1}{\|x(t)-x^{\prime}(t)\|_{2}}\sum_{k=1}^{d}(x_{k}(t)-x_{k}^{\prime}(t))(b_{k}^{*}(t,x(t))-b_{k}^{*}(t,x^{\prime}(t)))$
$\leq\|b^{*}(t,x(t))-b^{*}(t,x^{\prime}(t))\|_{2}\leq G\|x(t)-x^{\prime}(t)\|_{2},$
where the first inequality follows from Cauchy-Schwarz inequality, and the second inequality is due to (C.11). Then applying Gronwall’s inequality [Evans, 2010, Section B.2] completes the proof. ∎
Proof of Proposition 3.8.
According to Lemma C.6, we find
(C.12) $\|{\mathbb{E}}[X_{1}|X_{t}=x]\|_{2}\leq M_{1}R,\quad(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty},$
where $M_{2}$ is a constant only depending on $d$ and $\sigma$. For the conditional covariance, using Lemma C.7 implies
(C.13) $\|\operatorname{Cov}(X_{1}|X_{t}=x)\|_{\operatorname{op}}\leq M_{2}\alpha_{t}^{2},\quad(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty},$
where $M_{2}$ is a constant only depending on $d$ and $\sigma$. In addition, applying Lemma C.8 yields
(C.14) $\|{\mathbb{E}}[X_{1}X_{1}^{T}X_{1}|X_{t}=x]-{\mathbb{E}}[X_{1}^{T}X_{1}|X_{t}=x]{\mathbb{E}}[X_{1}|X_{t}]\|_{2}\leq M_{3}\alpha_{t}^{2}R,$
for each $(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty}$. Substituting (C.12), (C.13) and (C.14) to Lemma C.4 achieves the desired result. ∎
Appendix D Proof of Results in Section 3.3
In the section, we present the proofs of Theorem 3.9 and Corollary 3.10.
D.1. Proof of Theorem 3.9
In this section, we prove Theorem 3.9. Specifically, we propose the oracle inequality in Lemma D.1, which decomposes the $L^{2}$-risk into approximation error, the generalization error, and the truncation error. Then we provide an approximation error bound in Lemma D.2. By making a trade-off between three errors, we finally obtain the convergence rate for the velocity estimator, which completes the proof Theorem 3.9.
Recall the weighted $L^{2}$-risk (3.1) of a measurable function $b:{\mathbb{R}}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ as
${\mathcal{E}}_{T}(b)=\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b(t,X_{t})-b^{*}(t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t,$
and for the sake of notation simplicity, define the truncated $L^{2}$-risk with truncation parameter $R>1$ as
${\mathcal{E}}_{T,R}(b)=\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b(t,X_{t})-b^{*}(t,X_{t})\|_{2}^{2}\mathbbm{1}\{\|X_{t}\|_{\infty}\leq R\}\Big{]}\,\mathrm{d}t.$
Lemma D.1 (Oracle inequality for velocity estimation).
Suppose that Assumptions 1 and 2 hold. Let $T\in(1/2,1)$ and $R\in(1,+\infty)$. Further, assume that for each $b\in{\mathscr{B}}$,
(D.1) $\max_{1\leq k\leq d}|b_{k}(t,x)|\leq B_{\operatorname{vel}}R,\quad(t,x)\in[0,T]\times{\mathbb{R}}^{d},$
Then the following inequality holds for each $n\geq\max_{1\leq k\leq d}\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})$,
${\mathbb{E}}_{{\mathcal{S}}}\big{[}{\mathcal{E}}_{T}(\widehat{b})\big{]}\leq 2\inf_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)+C\lambda(T)R^{2}\Big{(}\max_{1\leq k\leq d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}(n)}+\frac{1}{\exp(\theta R^{2})}\Big{)},$
where the constant $\theta$ only depends on $\sigma$, the constant $C$ only depends on $d$ and $\sigma$, and the constant $\lambda(T)$ is defined as $\lambda(T)=\max\{1,\sup_{t\in[0,T]}\dot{\alpha}_{t}^{2}\}$.
Proof of Lemma D.1.
Before proceeding, we introduce some notations, aiming to reformulate the original velocity matching problem to a standard regression model.
Given a pair of random variables $(X_{0},X_{1})$ sampled from $\mu_{0}\times\mu_{1}$, define a stochastic process $Y_{t}=\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}X_{1}$ for each $t\in[0,T]$. Recall the stochastic interpolant $X_{t}=\alpha_{t}X_{0}+\beta_{t}X_{1}$. We define the noise term as $\varepsilon_{t}=Y_{t}-b^{*}(t,X_{t})$. Since $b^{*}(t,x)={\mathbb{E}}[Y_{t}\big{|}X_{t}=x]$, we have ${\mathbb{E}}[\varepsilon_{t}|X_{t}=x]=0$ for each $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$. Therefore, in the rest of this proof, it suffices to consider the following regression model:
(D.2) $Y_{t}=b^{*}(t,X_{t})+\varepsilon_{t},\quad X_{t}\sim\mu_{t},\leavevmode\nobreak\ t\sim\operatorname{Unif}[0,T].$
Recall the data set ${\mathcal{S}}=\{(t^{(i)},X_{0}^{(i)},X_{1}^{(i)})\}_{i=1}^{n}$. Then we define the data set corresponding to the regression model (D.2) as $\{(t^{(i)},X_{t}^{(i)},Y_{t}^{(i)})\}$, for which
$X_{t}^{(i)}=\alpha(t^{(i)})X_{0}^{(i)}+\beta(t^{(i)})X_{1}^{(i)}\quad\text{and}\quad Y_{t}^{(i)}=\dot{\alpha}(t^{(i)})X_{0}^{(i)}+\dot{\beta}(t^{(i)})X_{1}^{(i)}.$
The noise terms $\{\varepsilon_{t}^{(i)}\}_{i=1}^{n}$ can be defined by $\varepsilon_{t}^{(i)}=Y_{t}^{(i)}-b^{*}(t^{(i)},X_{t}^{(i)})$ for each $1\leq i\leq n$.
We divide the proof into four steps.
Step 1. Sub-Gaussian noise.
In this step, we show that the noise term $(\varepsilon_{t}|X_{t}=x)$ in (D.2) is sub-Gaussian for each $t\in[0,T]$ and $x\in{\mathbb{R}}^{d}$, and aim to estimate its variance proxy. According to the definition of stochastic interpolant $X_{t}$ (2.1), we have
$Y_{t}=\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}X_{1}=\frac{\dot{\alpha}_{t}}{\alpha_{t}}X_{t}+\frac{\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t}}{\alpha_{t}}X_{1},\quad t\in[0,T],$
which implies from Lemma C.5 that
$(Y_{t}|X_{t}=x)$ $=(\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}X_{1}|X_{t}=x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\frac{\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t}}{\alpha_{t}}(X_{1}|X_{t}=x)$
=$\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U+\sqrt{\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})}x,$
where $U\in[0,1]^{d}$ and $\epsilon\sim N(0,I_{d})$ are two independent variables. Then taking expectation on both sides of the equality yields
${\mathbb{E}}[Y_{t}|X_{t}=x]=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\sigma^{2}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})}x.$
Therefore, by the definition of the noise term, the following equality holds
$(\varepsilon_{t}|X_{t}=x)$ $=(Y_{t}|X_{t}=x)-{\mathbb{E}}[Y_{t}|X_{t}=x]$
=$\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}(U-{\mathbb{E}}[U])+\sqrt{\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon.$
Since that the random variable $U\in[0,1]^{d}$, using Hoeffding’s lemma [Mohri et al., 2018, Lemma D.1] implies that $U$ is $1$-sub-Gaussian. Further, applying Lemma A.4 deduces that each element of $(\varepsilon_{t}|X_{t}=x)$ is sub-Gaussian for each $t\in[0,T]$ and $x\in{\mathbb{R}}^{d}$ with variance proxy
(D.3) $\sigma_{T}^{2}=\sup_{t\in[0,T]}\Big{\{}\frac{\alpha_{t}^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}+\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{\}}\leq C\lambda(T),$
where $C$ is a constant only depending on $\sigma$.
Step 2. Truncation.
Notice that the velocity fields $b^{*}$ is defined on ${\mathbb{R}}^{d}$. It is necessary to restrict the original problem onto a compact subset of ${\mathbb{R}}^{d}$ by the technique of truncation. To begin with, we define the truncated population and empirical excess risks with radius $R>1$, respectively, as
${\mathcal{E}}_{T,R}(b)$ $=\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b^{*}(t,X_{t})-b(t,X_{t})\|_{2}^{2}\mathbbm{1}\{\|X_{t}\|_{\infty}\leq R\}\Big{]}\,\mathrm{d}t,$
^$\widehat{{\mathcal{E}}}_{T,R,n}(b)$ $=\frac{1}{n}\sum_{i=1}^{n}\|b^{*}(t^{(i)},X_{t}^{(i)})-b(t^{(i)},X_{t}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}.$
The population excess risk of the estimator $\widehat{b}$ can be decomposed by
(D.4) ${\mathbb{E}}_{{\mathcal{S}}}[{\mathcal{E}}_{T}(\widehat{b})]\leq{\mathbb{E}}_{{\mathcal{S}}}\Big{[}{\mathcal{E}}_{T}(\widehat{b})-{\mathcal{E}}_{T,R}(\widehat{b})\Big{]}+{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\sup_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)-2\widehat{{\mathcal{E}}}_{T,R,n}(b)\Big{]}+2{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big{]}.$
The first term in the right-hand side of (D.4) corresponds to the truncation error, which is estimated in the rest of this step. The second term of (D.4) is studied in Step 3. Finally, we bound the last term of (D.4) in Step 4.
For each hypothesis $b\in{\mathscr{B}}$, it follows that
${\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b^{*}(t,X_{t})-b(t,X_{t})\|_{2}^{2}\mathbbm{1}\{\|X_{t}\|_{\infty}>R\}\Big{]}$
$\leq{\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big{[}\|b^{*}(t,X_{t})-b(t,X_{t})\|_{2}^{4}\Big{]}{\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big{[}\mathbbm{1}\{\|X_{t}\|_{\infty}>R\}\Big{]}$
(D.5) $\leq 8\Big{(}{\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big{[}\|b(t,X_{t})\|_{2}^{4}\Big{]}+{\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big{[}\|b^{*}(t,X_{t})\|_{2}^{4}\Big{]}\Big{)}\operatorname{Pr}^{1/2}\{\|X_{t}\|_{\infty}>R\},$
where the first inequality follows from Cauchy-Schwarz inequality, and the second inequality is due to the triangular inequality. The boundedness of the hypothesis (D.1) deduces
(D.6) ${\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big{[}\|b(t,X_{t})\|_{2}^{4}\Big{]}\leq dB_{\operatorname{vel}}^{2}R^{2}.$
Then we consider the fourth moment of $b^{*}(t,X_{t})$ in (D.5). By using Assumption 2, we have
$Y_{t}=\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}X_{1}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}U+\sigma\dot{\beta}_{t}\epsilon,\quad U\sim\nu,\leavevmode\nobreak\ \epsilon\sim N(0,I_{d}),$
which implies
${\mathbb{E}}^{1/2}\Big{[}\|Y_{t}\|_{2}^{4}\Big{]}$ $\leq{\mathbb{E}}^{1/2}\Big{[}(\|\dot{\alpha}_{t}X_{0}\|_{2}+\|\dot{\beta}_{t}U\|_{2}+\|\sigma\dot{\beta}_{t}\epsilon\|_{2})^{4}\Big{]}$
$\leq 27\Big{(}\dot{\alpha}_{t}^{4}{\mathbb{E}}\Big{[}\|X_{0}\|_{2}^{4}\Big{]}+\dot{\beta}_{t}^{4}{\mathbb{E}}\Big{[}\|U\|_{2}^{4}\Big{]}+\sigma^{4}\dot{\beta}_{t}^{4}{\mathbb{E}}\Big{[}\|\epsilon\|_{2}^{4}\Big{]}\Big{)}^{1/2}$
$\leq 81d(\dot{\alpha}_{t}^{2}+\dot{\beta}_{t}^{2}+\sigma^{2}\dot{\beta}_{t}^{2})\leq C\lambda(T),\quad t\in[0,T],$
where the second inequality holds from the triangular inequality, the last inequality follows from Lemma A.7, and the constant $C$ only depends in $d$ and $\sigma$. Consequently,
(D.7) ${\mathbb{E}}_{X_{t}}^{1/2}\Big{[}\|b^{*}(t,X_{t})\|_{2}^{4}\Big{]}={\mathbb{E}}_{X_{t}}^{1/2}\Big{[}\|{\mathbb{E}}[Y_{t}|X_{t}]\|_{2}^{4}\Big{]}\leq{\mathbb{E}}^{1/2}\Big{[}\|Y_{t}\|_{2}^{4}\Big{]}\leq C\lambda(T),$
where the first inequality is due to the definition of velocity, and the second inequality follows from Jensen’s inequality. We next consider the tail probability of $X_{t}$ in (D.5). According to Assumption 2, we find
$X_{t}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\alpha_{t}X_{0}+\beta_{t}U+\sigma\beta_{t}\epsilon,\quad U\sim\nu,\leavevmode\nobreak\ \epsilon\sim N(0,I_{d}),$
which implies from Hoeffding’s lemma [Mohri et al., 2018, Lemma D.1] and Lemma A.4 that $X_{t}$ is a $(\alpha_{t}^{2}+\beta_{t}^{2}+\sigma^{2}\beta_{t}^{2})$-sub-Gaussian random variable. Then it follows from Lemma A.5 that
(D.8) $\sup_{t\in(0,1)}\operatorname{Pr}\big{\{}\|X_{t}\|_{\infty}>R\big{\}}\leq 2d\sup_{t\in(0,1)}\exp\Big{(}-\frac{R^{2}}{2(\alpha_{t}^{2}+\beta_{t}^{2}+\sigma^{2}\beta_{t}^{2})}\Big{)}\leq\frac{2d}{\exp(\theta R^{2})},$
where $\theta$ is a constant only depending on $\sigma$. Substituting (D.6), (D.7), and (D.8) into (D.5) yields
${\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b^{*}(t,X_{t})-b(t,X_{t})\|_{2}^{2}\mathbbm{1}\{\|X_{t}\|_{\infty}>R\}\Big{]}\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})},\quad t\in[0,T],$
where $C$ is a constant only depending on $d$ and $\sigma$. As a consequence, for each hypothesis $b\in{\mathscr{B}}$, it follows that
(D.9) ${\mathcal{E}}_{T}(b)-{\mathcal{E}}_{T,R}(b)\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})}.$
Hence the first term in the right-hand side of (D.4) can be bounded by
(D.10) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}{\mathcal{E}}_{T}(\widehat{b})-{\mathcal{E}}_{T,R}(\widehat{b})\Big{]}\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})}.$
We next bound another truncation term by a similar argument, which will be used in Step 4. It follows from Cauchy-Schwarz inequality and (D.8) that
(D.11) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}>R\}\Big{]}\leq{\mathbb{E}}_{{\mathcal{S}}}^{1/2}\Big{[}(\varepsilon_{t,k}^{(i)})^{4}\Big{]}\operatorname{Pr}^{1/2}\{\|X_{t}^{(i)}\|_{\infty}>R\}\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})},$
where $C$ is a constant only depending on $d$ and $\sigma$, and the last inequality follows from the fact that $\varepsilon_{t,k}^{(i)}$ is sub-Gaussian with variance proxy in (D.3).
Step 3. Relate the truncated population excess risk of the estimator with its empirical counterpart.
In this step, we prove the following inequality:
(D.12) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\sup_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)-2\widehat{{\mathcal{E}}}_{T,R,n}(b)\Big{]}\leq CR^{2}\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}(n)},$
where $C$ is a constant only depending on $d$ and $\sigma$, and $n\geq\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})$ for each $1\leq k\leq d$. For simplicity of notation, we define the $k$-th term of excess risks as
${\mathcal{E}}_{T,R}^{k}(b)$ $=\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}(b_{k}^{*}(t,X_{t})-b_{k}(t,X_{t}))^{2}\mathbbm{1}\{\|X_{t}\|_{\infty}\leq R\}\Big{]}\,\mathrm{d}t,$
^$\widehat{{\mathcal{E}}}_{T,R,n}^{k}(b)$ $=\frac{1}{n}\sum_{i=1}^{n}(b_{k}^{*}(t^{(i)},X_{t}^{(i)})-b_{k}(t^{(i)},X_{t}^{(i)}))^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}.$
Applying Proposition 3.2, (D.1), and Lemma G.1 yields the following inequality
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\sup_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}^{k}(b)-2\widehat{{\mathcal{E}}}_{T,R,n}^{k}(b)\Big{]}\leq CR^{2}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}(n)},\quad 1\leq k\leq d,$
which implies (D.12) by summing with respect to $1\leq k\leq d$.
Step 4. Estimate the empirical excess risk.
In this section, we shown the following bound for the empirical excess risk of the estimator
(D.13) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big{]}\leq 2\inf_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)+C\lambda(T)R^{2}\Big{(}\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}n}+\frac{1}{\exp(\theta R^{2})}\Big{)},$
where $C$ is a constant only depending on $d$ and $\sigma$, and $n\geq\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})$ for each $1\leq k\leq d$.
It is straightforward that
^$\widehat{{\mathcal{L}}}_{T,R,n}(\widehat{b})$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)}+b_{k}^{*}(t^{(i)},X_{t}^{(i)})-\widehat{b}_{k}(t^{(i)},X_{t}^{(i)}))^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}$
=^$=\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})+\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}$
$\quad+\frac{2}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}\varepsilon_{t,k}^{(i)}(b_{k}^{*}(t^{(i)},X_{t}^{(i)})-\widehat{b}_{k}(t^{(i)},X_{t}^{(i)}))\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}.$
Since $\widehat{b}$ is a minimizer of ^$\widehat{{\mathcal{L}}}_{T,n}(\cdot)$ over the hypothesis class ${\mathscr{B}}$, it holds that ^$\widehat{{\mathcal{L}}}_{T,R,n}(\widehat{b})\leq\widehat{{\mathcal{L}}}_{T,n}(\widehat{b})\leq\widehat{{\mathcal{L}}}_{T,n}(b)$ for each $b\in{\mathscr{B}}$. Consequently,
(D.14) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big{]}$ $\leq{\mathcal{L}}_{T}(b)-{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}\Big{]}$
$\quad+2{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}\varepsilon_{t,k}^{(i)}\widehat{b}_{k}(t^{(i)},X_{t}^{(i)})\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}\Big{]},$
where we used the fact that ${\mathbb{E}}[\widehat{{\mathcal{L}}}_{T,n}(b)]={\mathcal{L}}_{T}(b)$ and ${\mathbb{E}}[\varepsilon_{t,k}^{(i)}b^{*}(t^{(i)},X_{t}^{(i)})]=0$. For the first term in the right-hand side of (D.14), we have
${\mathcal{L}}_{T}(b)-{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}\Big{]}$
(D.15) $={\mathcal{E}}_{T}(b)+{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}>R\}\Big{]}\leq{\mathcal{E}}_{T,R}(b)+2\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})},$
where the inequality follows from (D.9) and (D.11). It remains to bound the second term in the right-hand side of (D.14). Plugging Proposition 3.2 and (D.3) into Lemma G.2, we have
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}\varepsilon_{t,k}^{(i)}\widehat{b}_{k}(t^{(i)},X_{t}^{(i)})\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}\Big{]}$
(D.16) $\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big{]}+C\lambda(T)R^{2}\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}n},$
where $C$ is a constant only depending $d$ and $\sigma$. Substituting (D.15) and (D.16) into (D.14) yields (D.13).
Finally, plugging (D.10), (D.12), and (D.13) into (D.4) completes the proof. ∎
Lemma D.2 (Approximation error).
Let $T\in(1/2,1)$ and $R\in(1,+\infty)$. Set the hypothesis class ${\mathscr{B}}$ as a deep neural network class, which is defined as
=${\mathscr{B}}=\left\{b\in N(L,S):\begin{aligned} &\|b(t,x)\|_{\infty}\leq B_{\operatorname{vel}}R,\leavevmode\nobreak\ \|\partial_{t}b(t,x)\|_{2}\leq 3D\kappa(T)R,\\ &\|\nabla b(t,x)\|_{\operatorname{op}}\leq 3G,\leavevmode\nobreak\ (t,x)\in[0,T]\times{\mathbb{R}}^{d}\end{aligned}\right\},$
where the depth and the width of the neural network are given, respectively, as $L=C$ and $S=CN^{d+1}$. Then the following inequality holds for each $N\in{\mathbb{N}}_{+}$,
$\inf_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)\leq C\kappa^{2}(T)R^{2}N^{-2},$
where $C$ is a constant only depending on $d$ and $\sigma$.
Proof of Lemma D.2.
Denote by ${\mathbb{B}}_{R,T}^{\infty}=[0,T]\times{\mathbb{B}}_{R}^{\infty}$. According to Corollary H.6, for each element $1\leq k\leq d$, there exists a real-valued deep neural network $b_{k}$ with depth $L_{k}=\lceil\log_{2}(d+1)\rceil+3$ and number of parameters $S_{k}=(22(d+1)+6)(N+1)^{d+1}$, such that
$\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}(b_{k}(t,X_{t})-b_{k}^{*}(t,X_{t}))^{2}\mathbbm{1}\{\|X_{t}\|_{\infty}\leq R\}\Big{]}\,\mathrm{d}t$
$\leq\|b_{k}-b_{k}^{*}\|_{L^{\infty}({\mathbb{B}}_{R,T}^{\infty})}^{2}\leq C\Big{(}T^{2}\|\partial_{t}b_{k}^{*}\|_{L^{\infty}({\mathbb{B}}_{R,T}^{\infty})}^{2}+R^{2}\sum_{\ell=1}^{d}\|\partial_{\ell}b_{k}^{*}\|_{L^{\infty}({\mathbb{B}}_{R,T}^{\infty})}^{2}\Big{)}N^{-2},$
where $C$ is a constant only depending on $d$ and $\sigma$, and the first inequality follows from Hölder’s inequality. Then we construct a vector-valued deep neural network $b(t,x)=(b_{k}(t,x))_{k=1}^{d}$ with depth $L=\max_{1\leq k\leq d}L_{k}$ and number of parameters $S=\sum_{k=1}^{d}S_{k}$, such that
$\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b_{k}(t,X_{t})-b_{k}^{*}(t,X_{t})\|_{2}^{2}\mathbbm{1}\{\|X_{t}\|_{\infty}\leq R\}\Big{]}\,\mathrm{d}t$
$=\sum_{k=1}^{d}\frac{1}{T}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}(b_{k}(t,X_{t})-b_{k}^{*}(t,X_{t}))^{2}\mathbbm{1}\{\|X_{t}\|_{\infty}\leq R\}\Big{]}\,\mathrm{d}t$
$\leq C\Big{(}T^{2}\sum_{k=1}^{d}\|\partial_{t}b_{k}^{*}\|_{L^{\infty}({\mathbb{B}}_{R,T}^{\infty})}^{2}+R^{2}\sum_{\ell=1}^{d}\sum_{k=1}^{d}\|\partial_{\ell}b_{k}^{*}\|_{L^{\infty}({\mathbb{B}}_{R,T}^{\infty})}^{2}\Big{)}N^{-2}\leq C^{\prime}\kappa^{2}(T)R^{2}N^{-2},$
where $C$ and $C^{\prime}$ are two constants only depending on $d$ and $\sigma$, and the last inequality holds from Propositions 3.5 and 3.8. This completes the proof. ∎
Proof of Theorem 3.9.
Set the hypothesis class as that in Lemma D.2. According to Lemma D.2, there exists a vector-valued deep neural network $b\in{\mathscr{B}}$ such that
(D.17) $\inf_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)\leq C\kappa^{2}(T)R^{2}N^{-2},$
where $C$ is a constant only depending on $d$ and $\sigma$. On the other hand, by applying Lemma A.12, the VC-dimension of this deep neural network class ${\mathscr{B}}$ is given as
(D.18) $\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})\leq CN^{d+1}\log N,$
where $C$ is an absolute constant. Plugging (D.17) and (D.18) into Lemma D.1 yields
${\mathbb{E}}_{{\mathcal{S}}}\big{[}{\mathcal{E}}_{T}(\widehat{b})\big{]}$ $\leq CR^{2}\Big{(}\frac{\kappa^{2}(T)}{N^{2}}+\frac{\lambda(T)N^{d+1}\log N}{n\log^{-1}n}+\frac{\lambda(T)}{\exp(\theta R^{2})}\Big{)}$
$\leq C\kappa^{2}(T)R^{2}\Big{(}\frac{1}{N^{2}}+\frac{N^{d+1}\log N}{n\log^{-1}n}+\frac{1}{\exp(\theta R^{2})}\Big{)},$
where $\theta$ is a constant only depending on $\sigma$, and $C$ is a constant only depending on $d$ and $\sigma$. Here the last inequality follows from the fact that
$\lambda(T)=\sup_{t\in[0,T]}\dot{\alpha}_{t}^{2}\leq C^{\prime}\sup_{t\in[0,T]}\Big{(}\frac{\dot{\alpha}_{t}^{2}}{\alpha_{t}^{2}}+\frac{|\ddot{\alpha}_{t}|}{\alpha_{t}}\Big{)}=C^{\prime}\kappa^{2}(T),$
where $C^{\prime}$ is a constant only depending on $\sigma$. By setting $N=Cn^{\frac{1}{d+3}}$, we obtain that
(D.19) ${\mathbb{E}}_{{\mathcal{S}}}\big{[}{\mathcal{E}}_{T}(\widehat{b})\big{]}\leq C\kappa^{2}(T)R^{2}\Big{(}n^{-\frac{2}{d+3}}\log(n)+\exp(-\theta R^{2})\Big{)}$
Then by substituting $R^{2}=\log(n)\theta^{-1}$, we obtain the desired result. ∎
D.2. Proof of Corollary 3.10
Before proceeding, recall the probability flow ODE with exact velocity field (2.4) and estimated velocity field (2.9), respectively, as
(D.20) $\,\mathrm{d}Z(t)$ $=b^{*}(t,Z(t))\,\mathrm{d}t,\quad t\in(0,T),$
$Z(0)$ $=Z_{0},$
and
(D.21) $\,\mathrm{d}\widehat{Z}(t)$ $=\widehat{b}(t,\widehat{Z}(t))\,\mathrm{d}t,\quad t\in(0,T),$
$\widehat{Z}(0)$ $=Z_{0}.$
The following lemma bounds the particle error by the velocity error.
Lemma D.3.
Let $\|{\mathscr{B}}\|_{\operatorname{Lip}}$ be the uniform Lipschitz constant of $b\in{\mathscr{B}}$. Then it follows that
$\|Z(T)-\widehat{Z}(T)\|_{2}\leq\exp(\|{\mathscr{B}}\|_{\operatorname{Lip}}T)\int_{0}^{T}\|b^{*}(t,Z(t))-\widehat{b}(t,Z(t))\|_{2}\,\mathrm{d}t.$
Proof of Lemma D.3.
It is straightforward that
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}^{2}$ $=\sum_{k=1}^{d}\frac{\,\mathrm{d}}{\,\mathrm{d}t}(Z_{k}(t)-\widehat{Z}_{k}(t))^{2}$
$=2\sum_{k=1}^{d}(Z_{k}(t)-\widehat{Z}_{k}(t))(b_{k}^{*}(t,Z(t))-\widehat{b}_{k}(t,\widehat{Z}(t)))$
(D.22) $\leq 2\|Z(t)-\widehat{Z}(t)\|_{2}\|b^{*}(t,Z(t))-\widehat{b}(t,\widehat{Z}(t))\|_{2},$
where the second inequality follows from (D.20) and (D.21), and the inequality holds from Cauchy-Schwarz inequality. On the other hand, we find
(D.23) $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}^{2}=2\|Z(t)-\widehat{Z}(t)\|_{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}.$
Combining (D.22) and (D.23) implies
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}\leq\|b^{*}(t,Z(t))-\widehat{b}(t,\widehat{Z}(t))\|_{2}.$
Then using the triangular inequality implies
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}$ $\leq\|b^{*}(t,Z(t))-\widehat{b}(t,Z(t))\|_{2}+\|\widehat{b}(t,Z(t))-\widehat{b}(t,\widehat{Z}(t))\|_{2}$
$\leq\|b^{*}(t,Z(t))-\widehat{b}(t,Z(t))\|_{2}+\|\widehat{b}\|_{\operatorname{Lip}}\|Z(t)-\widehat{Z}(t)\|_{2}.$
By using Gronwall’s inequality [Evans, 2010, Section B.2], we have
$\|Z(T)-\widehat{Z}(T)\|_{2}\leq\exp(\|\widehat{b}\|_{\operatorname{Lip}}T)\int_{0}^{T}\|b^{*}(t,Z(t))-\widehat{b}(t,Z(t))\|_{2}\,\mathrm{d}t.$
This completes the proof. ∎
Then we turn to estimate the distribution error using the particle error bound derived in Lemma D.3.
Proof of Corollary 3.10.
According to Lemma D.3, we have
$\|Z(T)-\widehat{Z}(T)\|_{2}\leq\exp(\|{\mathscr{B}}\|_{\operatorname{Lip}}T)\int_{0}^{T}\|b^{*}(t,Z(t))-\widehat{b}(t,Z(t))\|_{2}\,\mathrm{d}t.$
Taking expectation with respect to $Z_{0}\sim\mu_{0}$ implies
$W_{2}^{2}(\mu_{T},\widehat{\mu}_{T})$ $\leq{\mathbb{E}}_{Z_{0}\sim\mu_{0}}\big{[}\|Z(T)-\widehat{Z}(T)\|_{2}^{2}\big{]}$
$\leq\exp(2\|{\mathscr{B}}\|_{\operatorname{Lip}}T)\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b^{*}(t,X_{t})-\widehat{b}(t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t,$
where the first inequality follows from the definition of $2$-Wasserstein distance and Jensen’s inequality. Substituting Theorem 3.9 into the above inequality completes the proof. ∎
Appendix E Proofs of Results in Section 3.4
In this section, we present the proof of Theorem 3.11 and Corollary 3.12. Recall Euler scheme (2.10) as
(E.1) $\widehat{Z}_{k}$ $=\widehat{Z}_{k-1}+\widehat{b}(t_{k-1},\widehat{Z}_{k-1})\tau,\quad 1\leq k\leq K,$
$\widehat{Z}_{0}$ $=Z_{0}.$
The following lemma states the discretization error of Euler method, which uses some standard techniques in the numerical analysis for the forward Euler method [Iserles, 2008, Theorem 1.1].
Lemma E.1 (Discretization error of Euler method).
Let $\|{\mathscr{B}}\|_{\operatorname{Lip}}$ be the uniform Lipschitz constant of $b\in{\mathscr{B}}$. Then it follows that
$\|\widehat{Z}(T)-\widehat{Z}_{K}\|_{2}\leq\frac{T}{K}\frac{\exp(\|{\mathscr{B}}\|_{\operatorname{Lip}}T)-1}{\|{\mathscr{B}}\|_{\operatorname{Lip}}}D\kappa(T)R.$
Proof.
The proof is divided into two steps.
Step 1. Local truncation error estimate.
Consider the Taylor expansion of $\widehat{Z}(t_{k+1})$ around $t=t_{k}$,
$\widehat{Z}(t_{k+1})$ $=\widehat{Z}(t_{k})+\frac{\,\mathrm{d}}{\,\mathrm{d}t}\widehat{Z}(t_{k})\tau+\frac{\,\mathrm{d}^{2}}{\,\mathrm{d}t^{2}}\widehat{Z}(\theta)\tau^{2}$
(E.2) $=\widehat{Z}(t_{k})+\widehat{b}(t_{k},\widehat{Z}(t_{k}))\tau+\partial_{t}\widehat{b}(\theta,\widehat{Z}(\theta))\tau^{2},$
where $\theta\in[t_{k},t_{k+1}]$, and the second equality holds from the ODE (2.4). Recalling the forward Euler method (2.10)
(E.3) $\widehat{Z}_{k+1}=\widehat{Z}_{k}+\widehat{b}(t_{k},\widehat{Z}_{k})\tau.$
Subtracting (E.3) from (E.2) yields
$\|\widehat{Z}(t_{k+1})-\widehat{Z}_{k+1}\|_{2}$ $\leq\|\widehat{Z}(t_{k})-\widehat{Z}_{k}\|_{2}+\|\widehat{b}(t_{k},\widehat{Z}(t_{k}))-\widehat{b}(t_{k},\widehat{Z}_{k})\|_{2}\tau+D\kappa(T)R\tau^{2}$
(E.4) $\leq(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)\|\widehat{Z}(t_{k})-\widehat{Z}_{k}\|_{2}+D\kappa(T)R\tau^{2},$
where the first inequality holds from the triangular inequality and Proposition 3.8, and $\|{\mathscr{B}}\|_{\operatorname{Lip}}$ denotes the Lipschitz constant of $b$ with respect to the second variable. According to the definition of the hypothesis class ${\mathscr{B}}$ in Theorem 3.9, we have $\|{\mathscr{B}}\|_{\operatorname{Lip}}\leq 3G$.
Step 2. Global truncation error estimate.
We now show that the following inequality holds
(E.5) $\|\widehat{Z}(t_{k})-\widehat{Z}_{k}\|_{2}\leq\frac{(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)^{k}-1}{\|{\mathscr{B}}\|_{\operatorname{Lip}}}D\kappa(T)R\tau,\quad 1\leq k\leq K.$
We prove (E.5) by induction. When $k=1$, since that $\widehat{Z}(0)=\widehat{Z}_{0}$, it follows from (E.4) that
$\|\widehat{Z}(t_{1})-\widehat{Z}_{1}\|_{2}\leq\delta(t_{0},\widehat{Z}(t_{0}))\tau+D\kappa(T)R\tau^{2},$
which satisfies (E.5). For general $k\geq 2$ we assume that (E.5) holds up to $k-1$. Then applying (E.4) implies that
$\|\widehat{Z}(t_{k})-\widehat{Z}_{k}\|_{2}$ $\leq(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)\|\widehat{Z}(t_{k-1})-\widehat{Z}_{k-1}\|_{2}+D\kappa(T)R\tau^{2}$
$\leq\frac{(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)^{k}-1}{\|{\mathscr{B}}\|_{\operatorname{Lip}}}D\kappa(T)R\tau,$
which proves that (E.5) is true for $k$. Therefore, we have verified the inequality (E.5). Substituting $K\tau=T$ and $(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)^{k}\leq\exp(\|{\mathscr{B}}\|_{\operatorname{Lip}}k\tau)$into (E.5) completes the proof. ∎
Proof of Theorem 3.11.
Combining Lemmas D.3 and E.1, we have
$\|Z(T)-\widehat{Z}_{K}\|_{2}$ $\leq\|Z(T)-\widehat{Z}(T)\|_{2}+\|\widehat{Z}(T)-\widehat{Z}_{K}\|_{2}$
$\leq\exp(\|{\mathscr{B}}\|_{\operatorname{Lip}}T)\Big{(}\int_{0}^{T}\|b^{*}(t,Z(t))-\widehat{b}(t,Z(t))\|_{2}\,\mathrm{d}t+\frac{TD\kappa(T)}{K}R\Big{)}.$
Taking expectation with respect to $Z_{0}\sim\mu_{0}$ implies
$W_{2}^{2}(\mu_{T},\widehat{\mu}_{K})\leq{\mathbb{E}}_{Z_{0}\sim\mu_{0}}\big{[}\|Z(T)-\widehat{Z}_{K}\|_{2}^{2}\big{]}$
$\leq 2\exp(2\|{\mathscr{B}}\|_{\operatorname{Lip}}T)\Big{(}\int_{0}^{T}{\mathbb{E}}_{X_{t}\sim\mu_{t}}\Big{[}\|b^{*}(t,X_{t})-\widehat{b}(t,X_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}t+\frac{T^{2}D^{2}\kappa^{2}(T)}{K^{2}}R^{2}\Big{)}.$
Combining this inequality with Theorem 3.9 competes the proof. ∎
Proof of Corollary 3.12.
We first show that
(E.6) $W_{2}(\mu_{T},\mu_{1})\leq\max\{\alpha_{T},1-\beta_{T}\}W_{2}(\mu_{0},\mu_{1}).$
Indeed, let $X_{0}\sim\mu_{0}$ and $X_{1}\sim\mu_{1}$ be two independent random variables. Then $X_{T}=\alpha_{T}X_{0}+\beta_{T}X_{1}$ is a random variable obeying $\mu_{T}$. It follows that
$\|X_{T}-X_{1}\|_{2}=\|\alpha_{T}X_{0}-(1-\beta_{T})X_{1}\|_{2}\leq\max\{\alpha_{T},1-\beta_{T}\}\|X_{0}-X_{1}\|_{2}.$
Taking expectation on both sides of the inequality with respect to $X_{0}$ and $X_{1}$ and recalling the definition of 2-Wasserstein distance (1.2) implies (E.6).
According to the triangular inequality of the Wasserstein distance [Villani, 2009, Chapter 6], we have
$W_{2}(\widehat{\mu}_{K},\mu_{1})$ $\leq W_{2}(\widehat{\mu}_{K},\mu_{T})+W_{2}(\mu_{T},\mu_{1})$
$\leq W_{2}(\widehat{\mu}_{K},\mu_{T})+2\max\{\alpha_{t},1-\beta_{t}\}W_{2}(\mu_{0},\mu_{1}),$
where we used Lemma E.6. Combining this with Theorem 3.11 implies the desired result. ∎
Appendix F Proof of Results in Section 3.5
In this section, we aim to prove Theorem 3.13. Towards this end, we first relate the averaged $2$-Wasserstein distance of the characteristic generator to its $L^{2}$-risk by Lemma F.1. Then an oracle inequality of $L^{2}$-risk are proposed in Lemma F.2. Finally, by substituting approximation and generalization error bounds into the oracle inequality and using Theorem 3.11 completes the proof.
Lemma F.1.
Let $\widehat{g}_{t,s}$ be the estimator defined as (2.12). Then it follows that
$\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}W_{2}^{2}\Big{(}(\widehat{g}_{t,s})_{\sharp}\mu_{t},\mu_{s}\Big{)}\,\mathrm{d}s\mathrm{d}t$
$\leq{\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t})-\widehat{g}(t,s,Z_{t})\|_{2}^{2}\,\mathrm{d}s\mathrm{d}t\Big{]}.$
Proof of Lemma F.1.
According to the definition of 2-Wasserstein distance as Definition 1.2, it follows that
$W_{2}^{2}\Big{(}(\widehat{g}_{t,s})_{\sharp}\mu_{t},\mu_{s}\Big{)}\leq{\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\|\widehat{g}(t,s,Z_{t})-Z_{s}\|_{2}^{2}\Big{]}.$
Integrating both sides of the inequality with respect to $0\leq t\leq s\leq T$ deduces
$\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}W_{2}^{2}\Big{(}(\widehat{g}_{t,s})_{\sharp}\mu_{t},\mu_{s}\Big{)}\,\mathrm{d}s\mathrm{d}t$
$\leq\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}{\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\|Z_{s}-\widehat{g}(t,s,Z_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}s\mathrm{d}t$
$=\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}{\mathbb{E}}_{Z_{t}\sim\mu_{t}}\Big{[}\|g^{*}(t,s,Z_{t})-\widehat{g}(t,s,Z_{t})\|_{2}^{2}\Big{]}\,\mathrm{d}s\mathrm{d}t$
$={\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t})-\widehat{g}(t,s,Z_{t})\|_{2}^{2}\,\mathrm{d}s\mathrm{d}t\Big{]},$
which completes the proof. ∎
Lemma F.2 (Oracle inequality for characteristic fitting).
Suppose Assumptions 1 and 2 hold. Let $T\in(1/2,1)$ and $R\in(1,+\infty)$. Further, assume the hypothesis class ${\mathscr{G}}$ satisfies the following conditions for each $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$:
  1. (i)
    $\sup_{1\leq k\leq d}|g_{k}(t,s,x)|\leq B_{\operatorname{flow}}R$,
  2. (ii)
    $\|\partial_{t}g(t,s,x)\|_{2},\|\partial_{s}g(t,s,x)\|_{2}\leq 3B_{\operatorname{vel}}R$, and
  3. (iii)
    $\|\nabla g(t,s,x)\|_{\operatorname{op}}\leq 3\exp(\|{\mathscr{B}}\|_{\operatorname{Lip}}T)$.
Then the following inequality holds for each $m\geq\max_{1\leq k\leq d}\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})$,
${\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{R}}_{T}(\widehat{g})\big{]}$ $\leq\inf_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)+CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T})$
$\quad+CR^{2}\max_{1\leq k\leq d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})}{m\log^{-1}(m)}+\frac{CR^{2}}{K}+\frac{CR^{2}}{\exp(\theta R^{2})},$
where the constant $\theta$ only depends on $\sigma$, and the constant $C$ only depends on $d$ and $\sigma$.
Proof of Lemma F.2.
Recall the set of $m$ random variables ${\mathcal{Z}}=\{Z_{0}^{(i)}\}_{i=1}^{m}$ i.i.d. sampled from $\mu_{0}$. Further, let $Z_{t}^{(i)}$ denote the solution of the ODE (2.4) at time $t\in[0,1]$ given initial value $Z_{0}^{(i)}$ for each $1\leq i\leq m$, and let $\widehat{Z}_{k}^{(i)}$ denote the solution of Euler method (2.10) given the same initial value $Z_{0}^{(i)}$ at time $k\tau$ for $1\leq k\leq d$.
We first recall the empirical risk of the characteristic fitting:
[…]
Then replacing Euler solutions $\{\widehat{Z}_{k}^{(i)}:1\leq k\leq d\}_{i=1}^{m}$ in the empirical risk by exact solutions $\{Z_{k}^{(i)}:1\leq k\leq d\}_{i=1}^{m}$ of the ODE (2.4) yields an auxiliary empirical risk
[…]
Next we introduce a spatial truncation to (F.2), which implies the following risk:
[…]
We then define the following semi-discretized risk, which replaces the empirical average with respect to first two variables in (F.3) by its population :
(F.4) ^$\widehat{{\mathcal{R}}}_{T,R,m}(g)=\frac{1}{m}\sum_{i=1}^{m}\Big{\{}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t}^{(i)})-g(t,s,Z_{t}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|Z_{t}^{(i)}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big{\}}.$
Finally, recall the population risk of the characteristic fitting
(F.5) ${\mathcal{R}}_{T}(g)={\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t})-g(t,s,Z_{t})\|_{2}^{2}\,\mathrm{d}s\mathrm{d}t\Big{]},$
of which the spatial truncated counterpart is given as
(F.6) ${\mathcal{R}}_{T,R}(g)={\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}\|g^{*}(t,s,Z_{t})-g(t,s,Z_{t})\|_{2}^{2}\mathbbm{1}\{\|Z_{t}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big{]}.$
According to definitions (F.1) to (F.6), it is straightforward that for each $g\in{\mathscr{G}}$,
${\mathcal{R}}_{T}(\widehat{g})$ $\leq\Big{(}{\mathcal{R}}_{T}(\widehat{g})-{\mathcal{R}}_{T,R}(\widehat{g})\Big{)}+\Big{(}{\mathcal{R}}_{T,R}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,R,m}(\widehat{g})\Big{)}+2\Big{(}\widehat{{\mathcal{R}}}_{T,R,m}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})\Big{)}$
$\quad+2\Big{(}\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big{)}+\Big{(}2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})\Big{)}+\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g),$
where the inequality follows from the fact that $\widehat{g}$ is the minimizer of the empirical risk minimizer ^$\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}$ over the hypothesis class ${\mathscr{G}}$. Taking expectation on both sides of the inequality with respect to ${\mathcal{Z}}\sim\mu_{0}^{m}$ yields
(F.7) ${\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{R}}_{T}(\widehat{g})\big{]}$ $\leq\Big{(}{\mathcal{R}}_{T}(\widehat{g})-{\mathcal{R}}_{T,R}(\widehat{g})\Big{)}+{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\sup_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)-2\widehat{{\mathcal{R}}}_{T,R,m}(g)\Big{]}$
$\quad+2{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,R,m}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})\Big{]}+2{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big{]}$
$\quad+{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})\Big{]}+\inf_{g\in{\mathscr{G}}}{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g)\Big{]},$
where we used the inequality $\sup(a+b)\leq\sup(a)+\sup(b)$.
Up to now, the $L^{2}$-risk of the estimator $\widehat{g}$ is divided into six terms by (F.7). In the remainder of the proof, we bound six these error terms one by one.
  1. (i)
    The first term in the right-hand side of (F.7) measures the error caused by the spatial truncation, for which
    (F.8) ${\mathcal{R}}_{T}(\widehat{g})-{\mathcal{R}}_{T,R}(\widehat{g})\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
    where the constant $\theta$ only depends on $\sigma$, and $C$ only depends on $d$ and $\sigma$.
  2. (ii)
    The second term in the right-hand side of (F.7) is known as the generalization error, for which
    (F.9) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\sup_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)-2\widehat{{\mathcal{R}}}_{T,R,m}(g)\Big{]}\leq CR^{2}\max_{1\leq k\leq d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})}{m\log^{-1}(m)},$
    where the constant $\theta$ only depends on $\sigma$, and $C$ only depends on $d$ and $\sigma$.
  3. (iii)
    The third term in the right-hand side of (F.7) is led by the time discretization. It holds for each $g\in{\mathscr{G}}$ that
    (F.10) ^$\widehat{{\mathcal{R}}}_{T,R,m}(g)-\widehat{{\mathcal{R}}}_{T,R,m,K}(g)\leq\frac{CR^{2}}{K},$
    where the constant $\theta$ only depends on $\sigma$, and $C$ only depends on $d$ and $\sigma$.
  4. (iv)
    The fourth term in the right-hand side of (F.7) is also a truncation error. By an argument similar to (F.8), it holds that
    (F.11) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,R,m,K}(\widehat{g})-\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big{]}\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
    where the constant $\theta$ only depends on $\sigma$, and $C$ only depends on $d$ and $\sigma$.
  5. (v)
    The fifth term in the right-hand side of (F.7) is caused by the error of Euler method, for which
    (F.12) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})\Big{]}\leq CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T}),$
    where the constant $\theta$ only depends on $\sigma$, and $C$ only depends on $d$ and $\sigma$.
  6. (vi)
    The sixth term in the right-hand side of (F.7) is the empirical risk of the estimator. Using the definition of the empirical risk minimizer, we deduce
    (F.13) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g)\Big{]}$ $\leq\inf_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)+CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T})+\frac{CR^{2}}{K}+\frac{CR^{2}}{\exp(\theta R^{2})},$
    where the constant $\theta$ only depends on $\sigma$, and $C$ only depends on $d$ and $\sigma$.
Plugging (F.8) to (F.13) into (F.7) obtains the desired result.
Step 1. Estimate the first term in the right-hand side of (F.7).
For each hypothesis $g\in{\mathscr{G}}$, by an argument similar to (D.5), we have
${\mathbb{E}}_{Z_{t}\sim\mu_{t}}\Big{[}\|g^{*}(t,s,Z_{t})-g(t,s,Z_{t})\|_{2}^{2}\mathbbm{1}\{\|Z_{t}\|_{\infty}>R\}\Big{]}$
$\leq{\mathbb{E}}_{Z_{t}\sim\mu_{t}}^{1/2}\Big{[}\|g^{*}(t,s,Z_{t})-g(t,s,Z_{t})\|_{2}^{4}\Big{]}{\mathbb{E}}_{Z_{t}\sim\mu_{t}}^{1/2}\Big{[}\mathbbm{1}\{\|Z_{t}\|_{\infty}>R\}\Big{]}$
(F.14) $\leq 8\Big{(}{\mathbb{E}}_{Z_{t}\sim\mu_{t}}^{1/2}\Big{[}\|Z_{s}\|_{2}^{4}\Big{]}+{\mathbb{E}}_{Z_{t}\sim\mu_{t}}^{1/2}\Big{[}\|g(t,s,Z_{t})\|_{2}^{4}\Big{]}\Big{)}\operatorname{Pr}^{1/2}\{\|Z_{t}\|_{\infty}>R\},$
where the first inequality holds from Cauchy-Schwarz inequality, and the second inequality is due to the triangular inequality. By using Assumption 2, we have
$Z_{s}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}X_{s}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\alpha_{s}X_{0}+\beta_{s}U+\sigma\beta_{s}\epsilon,\quad U\sim\nu,\leavevmode\nobreak\ \epsilon\sim N(0,I_{d}),$
which implies by an argument similar to (D.7) that
${\mathbb{E}}_{Z_{s}}^{1/2}\Big{[}\|Z_{s}\|_{2}^{4}\Big{]}$ $={\mathbb{E}}_{X_{s}}^{1/2}\Big{[}\|X_{s}\|_{2}^{4}\Big{]}\leq{\mathbb{E}}_{(X_{0},U,\epsilon)}^{1/2}\Big{[}(\|\alpha_{s}X_{0}\|_{2}+\|\beta_{s}U\|_{2}+\|\sigma\beta_{s}\epsilon\|_{2})^{4}\Big{]}$
$\leq 27\Big{(}\alpha_{s}^{4}{\mathbb{E}}_{X_{0}}\Big{[}\|X_{0}\|_{2}^{4}\Big{]}+\beta_{s}^{4}{\mathbb{E}}_{U}\Big{[}\|U\|_{2}^{4}\Big{]}+\sigma^{4}\beta_{s}^{4}{\mathbb{E}}_{\epsilon}\Big{[}\|\epsilon\|_{2}^{4}\Big{]}\Big{)}^{1/2}$
$\leq 81d(\alpha_{s}^{2}+\beta_{s}^{2}+\sigma^{2}\beta_{s}^{2}),$
where the last inequality follows from Lemma A.7. As a consequence,
(F.15) ${\mathbb{E}}_{Z_{t}}^{1/2}\Big{[}\|g^{*}(t,s,Z_{t})\|_{2}^{4}\Big{]}={\mathbb{E}}_{Z_{s}}^{1/2}\Big{[}\|Z_{s}\|_{2}^{4}\Big{]}\leq C,$
where the constant $C$ only depends on $d$ and $\sigma$. Additionally, by using the boundedness of $g\in{\mathscr{G}}$, we have
(F.16) ${\mathbb{E}}_{Z_{t}}^{1/2}\Big{[}\|g(t,s,Z_{t})\|_{2}^{4}\Big{]}\leq dB_{\operatorname{flow}}^{2}R^{2}.$
Further, using (D.8) yields
(F.17) $\sup_{t\in(0,1)}\operatorname{Pr}\big{\{}\|Z_{t}\|_{\infty}>R\big{\}}=\sup_{t\in(0,1)}\operatorname{Pr}\big{\{}\|X_{t}\|_{\infty}>R\big{\}}\leq\frac{2d}{\exp(\theta R^{2})},$
where $\theta$ is a constant only depending on $\sigma$. Substituting inequalities (F.15), (F.16) and (F.17) into (F.14) deduces
${\mathbb{E}}_{Z_{t}\sim\mu_{t}}\Big{[}\|g^{*}(t,s,Z_{t})-\widehat{g}(t,s,Z_{t})\|_{2}^{2}\mathbbm{1}\{\|Z_{t}\|_{\infty}>R\}\Big{]}\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
where $C$ is a constant only depending to $d$ and $\sigma$. Finally, combining the above inequality with definitions (F.5) and (F.6) completes the proof of (F.8).
Step 2. Estimate the second term in the right-hand side of (F.7).
For simplicity of notation, we define the $k$-th term of ${\mathcal{R}}_{T,R}$ (F.6) and ^$\widehat{{\mathcal{R}}}_{T,R,m}$ (F.4), respectively, as
${\mathcal{R}}_{T,R}^{k}(b)$ $={\mathbb{E}}_{Z_{0}\sim\mu_{0}}\Big{[}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}(g_{k}^{*}(t,s,Z_{t})-g_{k}(t,s,Z_{t}))^{2}\mathbbm{1}\{\|Z_{t}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big{]},$
^$\widehat{{\mathcal{R}}}_{T,R,m}^{k}(b)$ $=\frac{1}{m}\sum_{i=1}^{m}\Big{\{}\frac{2}{T^{2}}\int_{0}^{T}\int_{t}^{T}(g_{k}^{*}(t,s,Z_{t}^{(i)})-g_{k}(t,s,Z_{t}^{(i)}))^{2}\mathbbm{1}\{\|Z_{t}^{(i)}\|_{\infty}\leq R\}\,\mathrm{d}s\mathrm{d}t\Big{\}}.$
Applying the boundedness of $g\in{\mathscr{G}}$, Proposition 3.3, and Lemma G.1 yields
${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\sup_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}^{k}(g)-2\widehat{{\mathcal{R}}}_{T,R,m}^{k}(g)\Big{]}\leq CR^{2}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})}{m\log^{-1}(m)},\quad 1\leq k\leq d.$
Summing the above inequalities with respect to $1\leq k\leq d$ completes the proof of (F.9).
Step 3. Estimate the third term in the right-hand side of (F.7).
For each fixed $x\in{\mathbb{B}}_{R}^{\infty}$, we define an auxiliary function
$u(t,s,x)=\|g^{*}(t,s,x)-\widehat{g}(t,s,x)\|_{2}^{2},\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty}.$
It is apparent that the following inequality holds for each $0\leq t\leq s\leq T$ and $x\in{\mathbb{B}}_{R}^{\infty}$,
(F.18) $|\partial_{t}u(t,s,x)|\leq 2\|g^{*}(t,s,x)-g(t,s,x)\|_{2}\|\partial_{t}g^{*}(t,s,x)-\partial_{t}g(t,s,x)\|_{2}\leq CR^{2},$
where the first inequality follows from Cauchy-Schwarz inequality, and the last inequality used Corollaries 3.3 and 3.4, and the definition of hypothesis class ${\mathscr{G}}$. Here the constant $C$ only depends on $d$ and $\sigma$. By the same argument, we have
(F.19) $|\partial_{s}u(t,s,x)|\leq CR^{2},\quad 0\leq t\leq s\leq T,\leavevmode\nobreak\ x\in{\mathbb{B}}_{R}^{\infty}.$
Substituting (F.18) and (F.19) into Lemma F.3 yields (F.10).
Step 4. Estimate the forth term in the right-hand side of (F.7).
We use an argument similar to Step 1. For each $10\leq k\leq\ell\leq K-$ 1, it follows that
${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|Z_{k\tau}^{(i)}\|_{\infty}>R\}\Big{]}$
$\leq 8\Big{(}{\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big{[}\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{4}\Big{]}+{\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big{[}\|\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{4}\Big{]}\Big{)}{\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big{[}\mathbbm{1}\big{\{}\|Z_{k\tau}^{(i)}\|_{\infty}>R\big{\}}\Big{]}$
(F.20) $=8\Big{(}{\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big{[}\|Z_{\ell\tau}^{(i)}\|_{2}^{4}\Big{]}+{\mathbb{E}}_{{\mathcal{Z}}}^{1/2}\Big{[}\|\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{4}\Big{]}\Big{)}\operatorname{Pr}^{1/2}\big{\{}\|Z_{k\tau}^{(i)}\|_{\infty}>R\big{\}},$
where we used Cauchy-Schwarz inequality and the triangular inequality. Substituting inequalities (F.15), (F.16) and (F.17) into (F.20) yields that for $10\leq k\leq\ell\leq K-$ 1,
(F.21) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|Z_{k\tau}^{(i)}\|_{\infty}>R\}\Big{]}\leq\frac{CR^{2}}{\exp(\theta R^{2})},$
where $C$ is a constant only depending on $d$ and $\sigma$. Combining (F.21) with definitions (F.2) and (F.3) deduces (F.11).
Step 5. Estimate the fifth term in the right-hand side of (F.7).
For each $1\leq i\leq m$ and $10\leq k\leq\ell\leq K-$ 1, it follows that
$\|\widehat{Z}_{\ell}^{(i)}-\widehat{g}(k\tau,\ell\tau,\widehat{Z}_{k}^{(i)})\|_{2}$
$\leq\|\widehat{Z}_{\ell}^{(i)}-Z_{\ell\tau}^{(i)}\|_{2}+\|Z_{\ell\tau}^{(i)}-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}+\|\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,\widehat{Z}_{k}^{(i)})\|_{2}$
$\leq\|\widehat{Z}_{\ell}^{(i)}-Z_{\ell\tau}^{(i)}\|_{2}+\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}+\|\widehat{g}\|_{\operatorname{Lip}}\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2},$
where we used the triangular inequality. Squaring both sides of the inequality yields
$\|\widehat{Z}_{\ell}^{(i)}-\widehat{g}(k\tau,\ell\tau,\widehat{Z}_{k}^{(i)})\|_{2}^{2}$
(F.22) $\leq 4\|\widehat{Z}_{\ell}^{(i)}-Z_{\ell\tau}^{(i)}\|_{2}^{2}+4\|\widehat{g}\|_{\operatorname{Lip}}^{2}\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}+2\|g^{*}(k\tau,\ell\tau,Z_{k\tau}^{(i)})-\widehat{g}(k\tau,\ell\tau,Z_{k\tau}^{(i)})\|_{2}^{2}.$
Substituting (F.22) into (F.1) deduces
^$\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})$ $\leq\frac{2}{mK^{2}}\sum_{i=1}^{m}\sum_{k=0}^{K-1}\Big{\{}\frac{1}{2}\Big{(}4\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}+4\|\widehat{g}\|_{\operatorname{Lip}}^{2}\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}\Big{)}$
$\quad+\sum_{\ell=k+1}^{K-1}\Big{(}4\|\widehat{Z}_{\ell}^{(i)}-Z_{\ell\tau}^{(i)}\|_{2}^{2}+4\|\widehat{g}\|_{\operatorname{Lip}}^{2}\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}^{2}\Big{)}\Big{\}}+2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})$
$\leq 4(1+\|\widehat{g}\|_{\operatorname{Lip}}^{2})\frac{1}{m}\sum_{i=1}^{m}\|\widehat{Z}_{K}^{(i)}-Z_{T}^{(i)}\|_{2}^{2}+2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g}),$
where the last inequality holds from the fact that
$\|\widehat{Z}_{k}^{(i)}-Z_{k\tau}^{(i)}\|_{2}\leq\|\widehat{Z}_{K}^{(i)}-Z_{T}^{(i)}\|_{2},\quad 0\leq k\leq K-1,$
which has been shown in the proof of Theorem 3.11. Consequently,
^$\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\leq 4(1+\|\widehat{g}\|_{\operatorname{Lip}}^{2})\frac{1}{m}\sum_{i=1}^{m}\|\widehat{Z}_{K}^{(i)}-Z_{T}^{(i)}\|_{2}^{2}.$
Taking expectation on both sides of the above inequality with respect to ${\mathcal{Z}}$ and plugging $\|\widehat{g}\|_{\operatorname{Lip}}\leq 3\exp(GT)$ imply
(F.23) ${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(\widehat{g})-2\widehat{{\mathcal{R}}}_{T,m,K}(\widehat{g})\Big{]}\leq CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T}),$
where $C$ is a constant only depending on $d$ and $\sigma$. By the same argument as inequality (F.23), we can obtain (F.12) immediately.
Step 6. Estimate the sixth term in the right-hand side of (F.7).
For each fixed $g\in{\mathscr{G}}$ independent of ${\mathcal{Z}}$, it follows that
${\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g)\Big{]}$ $={\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g)-2\widehat{{\mathcal{R}}}_{T,m,K}(g)\Big{]}+2{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,m,K}(g)-\widehat{{\mathcal{R}}}_{T,R,m,K}(g)\Big{]}$
$\quad+2{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,R,m,K}(g)-\widehat{{\mathcal{R}}}_{T,R,m}(g)\Big{]}+2{\mathbb{E}}_{{\mathcal{Z}}}\Big{[}\widehat{{\mathcal{R}}}_{T,R,m}(g)\Big{]},$
where the first term can be estimated by (F.23), the second and third terms can be bounded by an argument similar to (F.11) and (F.10), respectively. For the last term, we have ${\mathbb{E}}_{{\mathcal{Z}}}[\widehat{{\mathcal{R}}}_{T,R,m}(g)]={\mathcal{R}}_{T,R}(g)$. Combining above results yields (F.13). ∎
Proof of Theorem 3.13.
According to Lemma D.2, the following inequality holds
(F.24) $\inf_{g\in{\mathscr{G}}}{\mathcal{R}}_{T,R}(g)\leq\frac{CR^{2}}{N^{2}},$
where $C$ is a constant only depending on $d$ and $\sigma$. On the other hand, by applying Lemma A.12, the VC-dimension of this deep neural network class ${\mathscr{G}}$ is given as
(F.25) $\operatorname{VCdim}(\Pi_{k}{\mathscr{G}})\leq CN^{d+2}\log N,$
where $C$ is an absolute constant. Plugging (F.24) and (F.25) into Lemma F.2 yields
${\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{R}}_{T}(\widehat{g})\big{]}\leq\frac{CR^{2}}{N^{2}}+CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T})+CR^{2}\max_{1\leq k\leq d}\frac{N^{d+2}\log N}{m\log^{-1}(m)}+\frac{CR^{2}}{K}+\frac{CR^{2}}{\exp(\theta R^{2})},$
where $C$ is a constant only depending to $d$ and $\sigma$. By setting $N=Cm^{\frac{1}{d+4}}$ and $R^{2}=\log(m)\theta^{-1}$, we have
${\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{R}}_{T}(\widehat{g})\big{]}\leq Cm^{-\frac{2}{d+4}}\log^{2}(m)+CW_{2}^{2}(\widehat{\mu}_{K},\mu_{T})+\frac{C\log(m)}{K},$
Finally, we relate ${\mathcal{R}}_{T}(\widehat{g})$ to ${\mathcal{D}}(\widehat{g})$ by Lemma F.1. Finally, using Theorem 3.11 completes the proof. ∎
We conclude this section by giving an error bound for 2-dimensional numerical integral.
Lemma F.3.
Let $T>0$ and $K\in{\mathbb{N}}_{+}$. Assume that $u\in W^{1,\infty}([0,T]^{2})$. Define the step size as $\tau=T/K$, and define $\{t_{\ell}=\ell\tau\}_{\ell=0}^{K}$ as the set of time points. Then it follows that
$\frac{T^{2}}{K^{2}}\sum_{k=1}^{K}\Big{\{}\frac{1}{2}u(t_{k-1},t_{k-1})+\sum_{\ell=k+1}^{K}u(t_{k-1},t_{\ell-1})\Big{\}}-\int_{0}^{T}\int_{t}^{T}u(s,t)\,\mathrm{d}s\mathrm{d}t$
$\leq\big{(}\|\partial_{t}u\|_{L^{\infty}([0,T]^{2})}+\|\partial_{s}u\|_{L^{\infty}([0,T]^{2})}\big{)}\frac{T}{K}.$
Proof of Lemma F.3.
According to the Taylor expansion of $u(s,t)$ around $(t_{k-1},t_{\ell-1})$ with $1\leq k\leq\ell\leq K$, it follows that
(F.26) $u(t,s)=u(t_{k-1},t_{\ell-1})+\partial_{t}u(\theta_{k-1}^{t},t_{\ell-1})(t-t_{k-1})+\partial_{s}u(t_{k-1},\theta_{\ell-1}^{s})(s-t_{\ell-1}),$
where $\theta_{k-1}^{t}\in[t_{k-1},t]$ and $\theta_{\ell-1}^{s}\in[t_{\ell-1},s]$.
For $1\leq\ell=k\leq K$, integrating both sides of (F.26) on $(t,s)\in[t_{k-1},t_{k}]\times[t,t_{k}]$ yields
$\int_{t_{k-1}}^{t_{k}}\int_{t}^{t_{k}}u(t,s)\,\mathrm{d}s\mathrm{d}t-u(t_{k-1},t_{k-1})\frac{\tau^{2}}{2}$
$=\partial_{t}u(\theta_{k-1}^{t},t_{k-1})\int_{t_{k-1}}^{t_{k}}(t_{k}-t)(t-t_{k-1})\,\mathrm{d}t+\partial_{s}u(t_{k-1},\theta_{k-1}^{s})\int_{t_{k-1}}^{t_{k}}\int_{t}^{t_{k}}(s-t_{k-1})\,\mathrm{d}s\mathrm{d}t$
$=\partial_{t}u(\theta_{k-1}^{t},t_{k-1})\frac{\tau^{3}}{6}+\partial_{s}u(t_{k-1},\theta_{k-1}^{s})\frac{\tau^{3}}{3},$
where we used the fact that $t_{k}-t_{k-1}=\tau$. By summing both sides of the above equality with respect to $k=1,\ldots,K$, we obtain
$\frac{T^{2}}{2K^{2}}\sum_{k=1}^{K}u(t_{k-1},t_{k-1})-\sum_{k=1}^{K}\int_{t_{k-1}}^{t_{k}}\int_{t}^{t_{k}}u(t,s)\,\mathrm{d}s\mathrm{d}t$
(F.27) $\leq\sup_{(t,s)\in[0,T]^{2}}\big{\{}|\partial_{t}u(t,s)|+|\partial_{s}u(t,s)|\big{\}}\frac{T^{3}}{3K^{2}}.$
For $1\leq\ell<k\leq K$, integrating both sides of (F.26) on $(t,s)\in[t_{k-1},t_{k}]\times[t_{\ell-1},t_{\ell}]$ yields
$\int_{t_{k-1}}^{t_{k}}\int_{t_{\ell-1}}^{t_{\ell}}u(t,s)\,\mathrm{d}s\mathrm{d}t-u(t_{k-1},t_{\ell-1})\tau^{2}$
$=\partial_{t}u(\theta_{k-1}^{t},t_{\ell-1})\tau\int_{t_{k-1}}^{t_{k}}(t-t_{k-1})\,\mathrm{d}t+\partial_{s}u(t_{k-1},\theta_{\ell-1}^{s})\tau\int_{t_{\ell-1}}^{t_{\ell}}(s-t_{\ell-1})\,\mathrm{d}s$
$=\partial_{t}u(\theta_{k-1}^{t},t_{\ell-1})\frac{\tau^{3}}{2}+\partial_{s}u(t_{k-1},\theta_{\ell-1}^{s})\frac{\tau^{3}}{2}.$
By a similar argument to (F.27), it follows that
$\frac{T^{2}}{K^{2}}\sum_{k=1}^{K}\sum_{\ell=k+1}^{K}u(t_{k-1},t_{\ell-1})-\sum_{k=1}^{K}\sum_{\ell=k+1}^{K}\int_{t_{k-1}}^{t_{k}}\int_{t_{\ell-1}}^{t_{\ell}}u(t,s)\,\mathrm{d}s\mathrm{d}t$
(F.28) $\leq\sup_{(t,s)\in[0,T]^{2}}\big{\{}|\partial_{t}u(t,s)|+|\partial_{s}u(t,s)|\big{\}}\frac{T^{3}}{4K}.$
Summing (F.27) and (F.28) completes the proof. ∎
Appendix G Generalization Error Analysis for Least-Squares Regression
In this section, we provide a generalization error analysis for nonparametric regression, which is used in establishing oracle inequalities for velocity matching (Lemma D.1) and characteristic fitting (Lemma F.2).
Let ${\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ be a bounded domain, and let $X\in{\mathcal{X}}$ be a random variable obeying the probability distribution $\mu$. Let $f^{*}:{\mathcal{X}}\rightarrow{\mathbb{R}}$ be a measurable function. Define the population $L^{2}(\mu)$-risk for a function $f:{\mathcal{X}}\rightarrow{\mathbb{R}}$ as
$R(f)=\|f-f^{*}\|_{L^{2}(\mu)}^{2}={\mathbb{E}}_{X\sim\mu}\big{[}(f(X)-f^{*}(X))^{2}\big{]}.$
Let ${\mathcal{D}}=\{X^{(i)}\}_{i=1}^{n}$ be a set of i.i.d. copies of $X\sim\mu$. Then define the empirical risk of $f$ by
$\widehat{R}_{n}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X^{(i)})-f^{*}(X^{(i)}))^{2}.$
The following lemma relates the population risk to its empirical counterpart.
Lemma G.1.
Suppose that $|f^{*}(x)|\leq B$ for each $x\in{\mathcal{X}}$. Let ${\mathscr{F}}$ be a set of functions mapping from ${\mathcal{X}}$ to $[-B,B]$. Then it follows that for each $n\geq\operatorname{VCdim}({\mathscr{F}})$,
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in{\mathscr{F}}}R(f)-2\widehat{R}_{n}(f)\Big{]}\leq CB^{2}\frac{\operatorname{VCdim}({\mathscr{F}})}{n\log^{-1}n},$
where $C$ is an absolute constant.
This lemma provides a generalization error bound with fast rate via the technique of the offset Rademacher complexity, which was first proposed by Liang et al. [2015]. In recent years, this technique has been applied to the convergence rate analysis for deep nonparametric regression, such as Duan et al. [2023, Lemma 14] and Ding et al. [2024, Lemma 4.1].
Proof of Lemma G.1.
We define an auxiliary function class
=${\mathscr{H}}=\Big{\{}x\mapsto h(x)=(f(x)-f^{*}(x))^{2}:f\in{\mathscr{F}}\Big{\}}.$
It is apparent that $0\leq h(x)\leq 4B^{2}$ for each $x\in{\mathcal{X}}$ and $h\in{\mathscr{H}}$. Then it is easy to show that
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in{\mathscr{F}}}R(\widehat{f})-2\widehat{R}_{n}(\widehat{f})\Big{]}\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{h\in{\mathscr{H}}}{\mathbb{E}}[h(X)]-\frac{2}{n}\sum_{i=1}^{n}h(X^{(i)})\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{h\in{\mathscr{H}}}{\mathbb{E}}\Big{[}\frac{3}{2}h(X)-\frac{1}{8B^{2}}h^{2}(X)\Big{]}-\frac{1}{n}\sum_{i=1}^{n}\Big{(}\frac{3}{2}h(X^{(i)})+\frac{1}{8B^{2}}h^{2}(X^{(i)})\Big{)}\Big{]},$
where we used the fact that $h^{2}(x)\leq 4B^{2}h(x)$ for each $x\in{\mathcal{X}}$ and $h\in{\mathscr{H}}$.
Let us introduce a ghost sample ${\mathcal{D}}^{\prime}=\{X^{(i),\prime}\}_{i=1}^{n}$, which is a set of $n$ i.i.d. random copies of $X\sim\mu$. Here the ghost sample ${\mathcal{D}}^{\prime}$ is independent of ${\mathcal{D}}=\{X^{(i)}\}_{i=1}^{n}$. Let $\xi=\{\xi^{(i)}\}_{i=1}^{n}$ be a set of i.i.d. Rademacher variables. Then replacing the expectation by the empirical mean based on the ghost sample ${\mathcal{D}}^{\prime}$ yields
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{h\in{\mathscr{H}}}{\mathbb{E}}_{X}\Big{[}\frac{3}{2}h(X)-\frac{1}{8B^{2}}h^{2}(X)\Big{]}-\frac{1}{n}\sum_{i=1}^{n}\Big{(}\frac{3}{2}h(X^{(i)})+\frac{1}{8B^{2}}h^{2}(X^{(i)})\Big{)}\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{h\in{\mathscr{H}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\frac{3}{2}h(X^{(i),\prime})-\frac{1}{8B^{2}}h^{2}(X^{(i),\prime})\Big{]}-\frac{1}{n}\sum_{i=1}^{n}\Big{(}\frac{3}{2}h(X^{(i)})+\frac{1}{8B^{2}}h^{2}(X^{(i)})\Big{)}\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\sup_{h\in{\mathscr{H}}}\frac{3}{2n}\sum_{i=1}^{n}(h(X^{(i),\prime})-h(X^{(i)}))-\frac{1}{8B^{2}n}\sum_{i=1}^{n}(h^{2}(X^{(i),\prime})+h^{2}(X^{(i)}))\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}{\mathbb{E}}_{\xi}\Big{[}\sup_{h\in{\mathscr{H}}}\frac{3}{2n}\sum_{i=1}^{n}\xi^{(i)}(h(X^{(i),\prime})-h(X^{(i)}))-\frac{1}{8B^{2}n}\sum_{i=1}^{n}(h^{2}(X^{(i),\prime})+h^{2}(X^{(i)}))\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\xi}\Big{[}\sup_{h\in{\mathscr{H}}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h^{2}(X^{(i)})\Big{]},$
where the inequality holds from Jensen’s inequality. Combining the above results, we have
(G.1) ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in{\mathscr{F}}}R(\widehat{f})-2\widehat{R}_{n}(\widehat{f})\Big{]}\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\xi}\Big{[}\sup_{h\in{\mathscr{H}}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h^{2}(X^{(i)})\Big{]}.$
We next estimate the expectation in the right-hand side of (G.1). Let $\delta\in(0,4B^{2})$ and ${\mathscr{H}}_{\delta}$ be a $L^{\infty}({\mathcal{D}})$ $\delta$-cover of ${\mathscr{H}}$ satisfying $|{\mathscr{H}}_{\delta}|=N(\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))$. Then for each $h\in{\mathscr{H}}$, there exists $h_{\delta}\in{\mathscr{H}}_{\delta}$ such that
$\max_{1\leq i\leq n}|h(X^{(i)})-h_{\delta}(X^{(i)})|\leq\delta.$
Without loss of generality, we assume $|h_{\delta}(x)|\leq 4B^{2}$ for each $h_{\delta}\in{\mathscr{H}}_{\delta}$. Consequently, it follows from Hölder’s inequality that
$\frac{1}{n}\sum_{i=1}^{n}\xi^{(i)}h(X^{(i)})-\frac{1}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})\leq\frac{1}{n}\sum_{i=1}^{n}|\xi^{(i)}||h(X^{(i)})-h_{\delta}(X^{(i)})|\leq\delta.$
By the same argument, it holds that
$-\frac{1}{n}\sum_{i=1}^{n}h^{2}(X^{(i)})+\frac{1}{n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\leq 8B^{2}\delta.$
With the help of the above two inequalities, we have
${\mathbb{E}}_{\xi}\Big{[}\sup_{h\in{\mathscr{H}}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h^{2}(X^{(i)})\Big{]}$
(G.2) $\leq{\mathbb{E}}_{\xi}\Big{[}\max_{h_{\delta}\in{\mathscr{H}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\Big{]}+5\delta.$
Observe that $\{\xi^{(i)}h_{\delta}(X^{(i)})\}_{i=1}^{n}$ is a set of $n$ i.i.d. random variables with
$-h_{\delta}(X^{(i)})\leq\xi^{(i)}h_{\delta}(X^{(i)})\leq h_{\delta}(X^{(i)}),\quad 1\leq i\leq n.$
Then it follows Hoeffding’s inequality [Mohri et al., 2018, Theorem D.2] that
$\operatorname{Pr}_{\xi}\Big{\{}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})>t+\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\Big{\}}$
$=\operatorname{Pr}_{\xi}\Big{\{}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})>\frac{nt}{3}+\frac{1}{12B^{2}}\sum_{i=1}^{n}h_{\delta}^{2}(t^{(i)},X_{t}^{(i)})\Big{\}}$
(G.3) $\leq\exp\Big{(}-\frac{(\frac{nt}{3}+\frac{1}{12B^{2}}\sum_{i=1}^{n}h_{\delta}^{2}(t^{(i)},X_{t}^{(i)}))^{2}}{2\sum_{i=1}^{n}h_{\delta}^{2}(t^{(i)},X_{t}^{(i)})}\Big{)}\leq\exp\Big{(}-\frac{nt}{18B^{2}}\Big{)},$
where the first inequality follows from Hoeffding’s inequality [Mohri et al., 2018, Theorem D.2], and the second inequality is due to $(a+b)^{2}/b\leq 4a$ for each $a>0$ and $b\in{\mathbb{R}}$. As a consequence, for each $A>0$,
${\mathbb{E}}_{\xi}\Big{[}\max_{h_{\delta}\in{\mathscr{H}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\Big{]}$
$=\int_{0}^{\infty}\operatorname{Pr}_{\xi}\Big{\{}\max_{h_{\delta}\in{\mathscr{H}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})>t\Big{\}}\,\mathrm{d}t$
$\leq A+|{\mathscr{H}}_{\delta}|\int_{T}^{\infty}\exp\Big{(}-\frac{nt}{18B^{2}}\Big{)}\,\mathrm{d}t=A+\frac{18B^{2}}{n}|{\mathscr{H}}_{\delta}|\exp\Big{(}-\frac{nA}{18B^{2}}\Big{)},$
where the inequality is owing to (G.3). Letting $A=18B^{2}\log|{\mathscr{H}}_{\delta}|n^{-1}$ gives that
(G.4) ${\mathbb{E}}_{\xi}\Big{[}\max_{h_{\delta}\in{\mathscr{H}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\Big{]}\leq 18B^{2}\frac{\log|{\mathscr{H}}_{\delta}|+1}{n}.$
It remains to estimate the covering number $|{\mathscr{H}}_{\delta}|=N(\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))$. Noticing that
$|h(x)-h^{\prime}(x)|=|(f(x)-f^{*}(x))^{2}-(f^{\prime}(x)-f^{*}(x))^{2}|\leq 4B|f(x)-f^{\prime}(x)|,$
we obtain that for $n\geq\operatorname{VCdim}({\mathscr{F}})$,
(G.5) $\log N(\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))\leq\log N\Big{(}\frac{\delta}{4B},{\mathscr{F}},L^{\infty}({\mathcal{D}})\Big{)}\leq\operatorname{VCdim}({\mathscr{F}})\log\Big{(}\frac{4eB^{2}n}{\delta}\Big{)},$
where the first and last inequalities follows from Lemmas A.10 and A.11, respectively. Combining (G.1), (G.2), (G.4) and (G.5), we have
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in{\mathscr{F}}}R(\widehat{f})-2\widehat{R}_{n}(\widehat{f})\Big{]}$ $\leq\inf_{\delta>0}\Big{\{}36B^{2}\frac{\operatorname{VCdim}({\mathscr{F}})}{n}\log\Big{(}\frac{4eB^{2}n}{\delta}\Big{)}+5\delta\Big{\}}.$
Substituting $\delta=4B^{2}/n$ into the above inequality completes the proof. ∎
Lemma G.2.
Suppose that $|f^{*}(x)|\leq B$ for each $x\in{\mathcal{X}}$. Let ${\mathscr{F}}$ be a set of functions mapping from ${\mathcal{X}}$ to $[-B,B]$. Let $\{\varepsilon^{(i)}\}_{i=1}^{n}$ be a set of independent $\sigma^{2}$-sub-Gaussian random variables. Then it follows that for each $n\geq\operatorname{VCdim}({\mathscr{F}})$,
${\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\varepsilon^{(i)}\widehat{f}(X^{(i)})\Big{]}\leq\frac{1}{4}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\big{[}\widehat{R}(\widehat{f})\big{]}+C(B^{2}+\sigma^{2})\frac{\operatorname{VCdim}({\mathscr{F}})}{n\log^{-1}n},$
where $C$ is an absolute constant.
This proof uses a technique similar to the proof of Schmidt-Hieber [2020, Lemma 4].
Proof of Lemma G.2.
Let $\delta\in(0,B)$ and let ${\mathscr{F}}_{\delta}$ be a $L^{\infty}({\mathcal{D}})$ $\delta$-cover of ${\mathscr{F}}$ with $|{\mathscr{F}}_{\delta}|=N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))$. Then there exists $\widehat{f}_{\delta}\in{\mathscr{F}}_{\delta}$, such that
$\max_{1\leq i\leq n}|\widehat{f}(X^{(i)})-\widehat{f}_{\delta}(X^{(i)})|\leq\delta.$
Then it follows from Hölder’s inequality that
(G.6) ${\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\varepsilon^{(i)}(\widehat{f}(X^{(i)})-\widehat{f}_{\delta}(X^{(i)}))\Big{]}\leq\delta{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\frac{1}{n}\sum_{i=1}^{n}|\varepsilon^{(i)}|\Big{]}\leq\delta\sigma,$
where we used the fact that $\{\varepsilon^{(i)}\}_{i=1}^{n}$ are a set of $\sigma^{2}$-sub-Gaussian random variables. Additionally, according to the triangular inequality, we have
(G.7) $\widehat{R}_{n}^{1/2}(\widehat{f}_{\delta})-\widehat{R}_{n}^{1/2}(\widehat{f})\leq\Big{(}\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{\delta}(X^{(i)})-\widehat{f}(X^{(i)}))^{2}\Big{)}^{1/2}\leq\delta.$
Consequently, we have
${\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\varepsilon^{(i)}\widehat{f}(X^{(i)})\Big{]}$ $={\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\varepsilon^{(i)}(\widehat{f}(X^{(i)})-\widehat{f}_{\delta}(X^{(i)})+\widehat{f}_{\delta}(X^{(i)})-f^{*}(X^{(i)}))\Big{]}$
$\leq{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\varepsilon^{(i)}(\widehat{f}_{\delta}(X^{(i)})-f^{*}(X^{(i)}))\Big{]}+\delta\sigma$
$\leq\frac{1}{\sqrt{n}}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\Big{(}\widehat{R}_{n}^{1/2}(\widehat{f})+\delta\Big{)}\sum_{i=1}^{n}\frac{\varepsilon^{(i)}(\widehat{f}_{\delta}(X^{(i)})-f^{*}(X^{(i)}))}{\sqrt{n}\widehat{R}_{n}^{1/2}(\widehat{f}_{\delta})}\Big{]}+\delta\sigma$
$\leq\frac{1}{\sqrt{n}}\Big{(}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}^{1/2}\Big{[}\widehat{R}_{n}(\widehat{f})\Big{]}+\delta\Big{)}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}^{1/2}\Big{[}\psi^{2}(\widehat{f}_{\delta})\Big{]}+\delta\sigma$
(G.8) $\leq\frac{1}{4}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\widehat{R}_{n}(\widehat{f})\Big{]}+\frac{2}{n}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\big{[}\psi^{2}(\widehat{f}_{\delta})\big{]}+\frac{1}{4}\delta^{2}+\delta\sigma,$
where $\psi(\widehat{f}_{\delta})$ is defined as
$\psi(\widehat{f}_{\delta})=\sum_{i=1}^{n}\frac{\widehat{f}_{\delta}(X^{(i)})-f^{*}(X^{(i)})}{\sqrt{n}\widehat{R}_{n}^{1/2}(\widehat{f}_{\delta})}\varepsilon^{(i)}.$
Here the first inequality follows from (G.6), the second inequality holds from (G.7), the third inequality is due to Cauchy-Schwarz inequality, and the last inequality is owing to the weighted AM-GM inequality $ab\leq a/4+b$ for each $a,b\in{\mathbb{R}}$.
Observe that for each fixed function $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$, the random variable $\psi(f)$ is sub-Gaussian with variance proxy $\sigma^{2}$ conditioning on ${\mathcal{D}}=\{X^{(i)}\}_{i=1}^{n}$. Then it follows that
(G.9) ${\mathbb{E}}_{\varepsilon}\big{[}\psi^{2}(\widehat{f}_{\delta})\big{]}\leq{\mathbb{E}}_{\varepsilon}\Big{[}\max_{f_{\delta}\in{\mathscr{F}}_{\delta}}\psi^{2}(f_{\delta})\Big{]}\leq 4\sigma^{2}(\log|{\mathscr{F}}_{\delta}|+1).$
We now estimate the covering number $|{\mathscr{F}}_{\delta}|=N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))$. It follows from Lemma A.11 that for $n\geq\operatorname{VCdim}({\mathscr{F}})$,
(G.10) $\log N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))\leq\operatorname{VCdim}({\mathscr{F}})\log\Big{(}\frac{eBn}{\delta}\Big{)}.$
Combining (G.8), (G.9) and (G.10), and setting $\delta=B/n$ complete the proof. ∎
Appendix H Approximation by Deep Neural Networks with Lipschitz Constraint
The approximation error analysis for deep neural networks has been investigated by Yarotsky [2017, 2018], Yarotsky and Zhevnerchuk [2020], Shen et al. [2019], Shen [2020], Lu et al. [2021], Petersen and Voigtlaender [2018], Jiao et al. [2023a], Duan et al. [2022]. However, limited work has been done for deep neural networks with Lipschitz constraint [Huang et al., 2022, Chen et al., 2022, Jiao et al., 2023b, Ding et al., 2024].
This proof is based on the proof of Yarotsky [2017, Theorem 1]. The ReLU activation function is defined as $\operatorname{ReLU}(x)=\max\{0,x\}$, and the ReQU activation function is defined as the squared ReLU function $\operatorname{ReQU}(x)=(\max\{0,x\})^{2}$.
Lemma H.1.
The maximum or minimum of two inputs can be implemented by a ReLU neural network with $1$ hidden layer and $7$ non-zero parameters.
Proof of Lemma H.1.
According to the equality $a=\operatorname{ReLU}(a)-\operatorname{ReLU}(-a)$, the identity mapping can be implemented by a ReLU neural network with $1$ hidden layer and $4$ non-zero parameters. We also notice that
$\max\{a,b\}=a+\operatorname{ReLU}(b-a)=\operatorname{ReLU}(a)-\operatorname{ReLU}(-a)+\operatorname{ReLU}(b-a),$
which means that the maximum of two inputs can be implemented by $7$ non-zero parameters. By a same argument, with the aid of equality $\min\{a,b\}=a-\operatorname{ReLU}(a-b)$, we can obtain a same result for the minimum of two inputs. This completes the proof. ∎
Lemma H.2.
The product of two inputs can be implement by a ReQU neural network with $1$ hidden layer and $12$ non-zero parameters.
Proof of Lemma H.2.
According to Li et al. [2019, Lemma 2.1], the following identities hold:
$x_{1}x_{2}=\frac{1}{4}w_{3}^{T}\operatorname{ReQU}(w_{1}x_{1}+w_{2}x_{2}),$
where $w_{1}=(1,-1,1,-1)^{T}$, $w_{2}=(1,-1,-1,1)^{T}$ and $w_{3}=(1,1,-1,-1)^{T}$, which completes the proof. ∎
Lemma H.3.
The product of $p$ inputs can be implement by a ReQU neural network with $\lceil\log_{2}p\rceil$ hidden layers and $6(p+1)$ non-zero parameters.
Proof of Lemma H.3.
Define the augmented input vector $(x_{1},\ldots,x_{p},x_{p+1},\ldots,x_{n})$ where $n=2^{\lceil\log_{2}p\rceil}$ and $x_{i}=1$ for $p+1\leq i\leq n$. Observe that $\prod_{i=1}^{p}x_{i}=\prod_{i=1}^{n}x_{i}$. According to Lemma H.2, the mapping $(x_{1},\ldots,x_{n})\rightarrow(x_{1}x_{2},\ldots,x_{n-1}x_{n})\in{\mathbb{R}}^{n/2}$ can be implemented by a ReQU neural network with $1$ hidden layer and $6n$ non-zero parameters. By a same argument, we can construct a ReQU neural network with $1$ hidden layer and $3n$ non-zero parameters, which maps $(x_{1}x_{2},\ldots,x_{n-1}x_{n})$ to $(x_{1}x_{2}x_{3}x_{4},\ldots,x_{n-3}x_{n-2}x_{n-1}x_{n})\in{\mathbb{R}}^{n/4}$. By induction on $n$, the number of layers is given as $\lceil\log_{2}p\rceil$, and the total number of non-zero parameters is given by
$12\times(1+2+2^{2}+\ldots+2^{\lceil\log_{2}p\rceil-1})\leq 6(p+1).$
This completes the proof. ∎
Lemma H.4.
The univariate trapezoid function
(H.1) $\psi(z)=\left\{\begin{aligned} &1,&&|z|<1,\\ &2-|z|,&&1\leq|z|\leq 2,\\ &0,&&2<|z|,\end{aligned}\right.$
can be implement by a ReLU neural network with $3$ hidden layers and $14$ non-zero parameters.
Proof of Lemma H.4.
We first implement the following hat-function by ReLU neural network
$\widetilde{\psi}(z)=\left\{\begin{aligned} &2-|z|,&&|z|\leq 2,\\ &0,&&2<|z|.\end{aligned}\right.$
Noticing that $\widetilde{\psi}(z)=\min\{\operatorname{ReLU}(z+2),\operatorname{ReLU}(-z+2)\}$, by applying Lemma H.1, we find that $\widetilde{\psi}$ can be implemented by a ReLU neural network with $2$ hidden layers and $11$ non-zero parameters. Further, according to the equality $\psi(z)=\min\{1,\widetilde{\psi}(z)\}$, the univariate trapezoid function $\psi$ can be implemented by a ReLU neural network with $3$ hidden layers and $14$ non-zero parameters. This completes the proof. ∎
Lemma H.5 (Approximation error).
Let $p\in{\mathbb{N}}_{+}$, and let $\{N_{k}\}_{k=1}^{p}$ be a set of positive integer. Set the deep neural network class $N(L,S)$ as $L=\lceil\log_{2}p\rceil+3$ and $S=(22p+6)\prod_{k=1}^{p}(N_{k}+1)$. Then for each $u^{*}\in W^{1,\infty}([0,1]^{p})$, there exists a deep neural network $u\in N(L,S)$ such that
$\|u-u^{*}\|_{L^{\infty}([0,1]^{p})}\leq 2^{p}\sum_{k=1}^{p}\frac{1}{N_{k}}\|\partial_{k}u^{*}\|_{L^{\infty}({\mathcal{X}})}.$
Further, it holds that the following holds for each $1\leq k\leq p$:
$\|u\|_{L^{\infty}([0,1]^{p})}=\|u^{*}\|_{L^{\infty}([0,1]^{p})}\quad\text{and}\quad\|\partial_{k}u\|_{L^{\infty}([0,1]^{p})}\leq 3\|\partial_{k}u^{*}\|_{L^{\infty}([0,1]^{p})}.$
Proof of Lemma H.5.
The proof is divided into three steps. In the first step, we approximate the target function based on a partition of unity and the degree-00 Taylor expansion. Then we implement this piece-wise linear function using deep neural network exactly in the second step. Finally, in the last step, we estimate the Lipschitz constant of the deep neural network.
Step 1. Approximate the target function by a piecewise linear function.
Consider a partition of unity formed by a grid of $\prod_{k=1}^{p}(N_{k}+1)$ functions $\phi_{m}$ on the domain $[0,1]^{p}$:
(H.2) $\sum_{m}\phi_{m}(x)\equiv 1,\quad x\in[0,1]^{p},$
where the multi-index $m$ is defined as $m=(m_{1},\ldots,m_{p})^{T}$ with $m_{k}\in\{0,\ldots,N_{k}\}$, and the function $\phi_{m}$ is defined as the product
(H.3) $\phi_{m}(x)=\prod_{k=1}^{p}\psi\Big{(}3N_{k}\Big{(}x_{k}-\frac{m_{k}}{N_{k}}\Big{)}\Big{)}.$
Here $\psi$ is the univariate trapezoid function defined as (H.1). It is noticeable that for each $m$,
(H.4) $\sup_{z\in[0,1]}|\psi(z)|=1,\quad\sup_{x\in[0,1]^{p}}|\phi_{m}(x)|=1,$
and
[…]
Now define a piecewise linear approximation to $u^{*}$ by
(H.6) $u(x)=\sum_{m}\phi_{m}(x)u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}.$
Then it follows that
$|u^{*}(x)-u(x)|\leq\sum_{m}\Big{|}\phi_{m}(x)\Big{(}u^{*}(x)-u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}\Big{)}\Big{|}$
$\leq\sum_{m}\Big{|}u^{*}(x)-u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}\Big{|}\mathbbm{1}\Big{\{}m:\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\leq\frac{2}{3N_{k}},\leavevmode\nobreak\ k\in[p]\Big{\}}$
$\leq 2^{p}\max_{m}\Big{|}u^{*}(x)-u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}\Big{|}\mathbbm{1}\Big{\{}m:\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\leq\frac{2}{3N_{k}},\leavevmode\nobreak\ k\in[p]\Big{\}}$
$\leq 2^{p}\max_{m}\sum_{k=1}^{p}\mathop{\operatorname{ess}\,\sup}_{x\in[0,1]^{p}}|\partial_{k}u^{*}(x)|\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\mathbbm{1}\Big{\{}m:\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\leq\frac{2}{3N_{k}},\leavevmode\nobreak\ k\in[p]\Big{\}}$
$\leq 2^{p}\sum_{k=1}^{p}\mathop{\operatorname{ess}\,\sup}_{x\in[0,1]^{p}}|\partial_{k}u^{*}(x)|\frac{2}{3N_{k}},$
where the first inequality follows from the triangular inequality, the second inequality is due to (H.4) and (H.5), the third inequality used the observation that any $x\in[0,1]^{p}$ belongs to the support of at most $2^{d}$ functions $\phi_{m}$, the forth inequality used Taylor’s theorem of degree-00, and the last inequality holds from Hölder’s inequality. Consequently, we obtain the following inequality
$\|u-u^{*}\|_{L^{\infty}([0,1]^{p})}\leq 2^{p}\sum_{k=1}^{p}\frac{1}{N_{k}}\|\partial_{k}u^{*}\|_{L^{\infty}([0,1]^{p})}.$
Step 2. Implement the piecewise linear function by a deep neural network.
In this step, we implement the piececwise linear approximation (H.6) by a deep neural network. Using Lemmas H.3 and H.4, for each $m$, the function $\phi_{m}$ defined as (H.3) can be implemented by a deep neural network with $\lceil\log_{2}p\rceil+3$ layers and $616p+6(p+1)=22p+$ 6 non-zero parameters. Since $u$ defined in (H.6) is a linear combination of $\prod_{k=1}^{p}(N_{k}+1)$ functions $\phi_{m}$, it can be implemented by a deep neural network with $\lceil\log_{2}p\rceil+3$ layers and $(22p+6)\prod_{k=1}^{p}(N_{k}+1)$ non-zero parameters.
Step 3. Compute the Lipschitz constant of the deep neural network.
According to Step 2, the piecewise linear approximation $u$ can be implemented by a deep neural network with no error. Therefore, it suffices to compute the Lipschitz constant of $u$ in (H.6). Taking derivative on both sides of (H.6) with respect to $x_{k}$ yields
$\partial_{k}u(x)$ $=\sum_{m}u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}\partial_{k}\phi_{m}(x)$
$=\sum_{m}u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}A_{k}(m)\partial_{k}\psi\Big{(}3N_{k}\Big{(}x_{k}-\frac{m_{k}}{N_{k}}\Big{)}\Big{)}$
(H.7) $=\sum_{m}u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}A_{k}(m)\partial_{k}\psi\Big{(}3N_{k}\Big{(}x_{k}-\frac{m_{k}}{N_{k}}\Big{)}\Big{)}\mathbbm{1}\Big{\{}m:\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\leq\frac{2}{3N_{k}}\Big{\}},$
where the constant is given as
$A_{k}(m)=\prod_{\ell\neq k}\psi\Big{(}3N_{\ell}\Big{(}x_{\ell}-\frac{m_{\ell}}{N_{\ell}}\Big{)}\Big{)}\mathbbm{1}\Big{\{}m:\Big{|}x_{\ell}-\frac{m_{\ell}}{N_{\ell}}\Big{|}\leq\frac{2}{3N_{\ell}},\leavevmode\nobreak\ \ell\neq k\Big{\}}.$
It is evident that $0\leq A_{k}(m)\leq 1$ for each $1\leq k\leq p$ and $m$.
We next estimate (H.7) in the following cases.
  1. (i)
    If there exists $m_{k}^{*}\in\{1,\ldots,N_{k}\}$ such that $|x_{k}-\frac{m_{k}^{*}}{N_{k}}|\leq\frac{1}{3N_{k}}$, then
    $\partial_{k}u(x)=u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}A_{k}(m^{*})\partial_{k}\psi\Big{(}3N_{k}\Big{(}x_{k}-\frac{m_{k}^{*}}{N_{k}}\Big{)}\Big{)}=0,$
    where $m^{*}=(m_{1},\ldots,m_{k}^{*},\ldots,m_{p})^{T}$.
  2. (ii)
    If there exists $m_{k}^{*}\in\{1,\ldots,N_{k}\}$ such that $\frac{m_{k}^{*}}{N_{k}}+\frac{1}{3N_{k}}\leq x_{k}\leq\frac{m_{k}^{*}}{N_{k}}+\frac{2}{3N_{k}}$, then
    $\partial_{k}u(x)$ $=u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}A_{k}(m^{*})\partial_{k}\psi\Big{(}3N_{k}\Big{(}x_{k}-\frac{m_{k}^{*}}{N_{k}}\Big{)}\Big{)}$
    $\quad+u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}+1}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}A_{k}(m_{+}^{*})\partial_{k}\psi\Big{(}3N_{k}\Big{(}x_{k}-\frac{m_{k}^{*}+1}{N_{k}}\Big{)}\Big{)}$
    $=-3u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}A_{k}(m^{*})N_{k}$
    $\quad+3u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}+1}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}A_{k}(m_{+}^{*})N_{k}$
    $\leq 3N_{k}\Big{|}u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}+1}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}-u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}\Big{|}$
    $\leq 3\mathop{\operatorname{ess}\,\sup}_{x\in[0,1]^{p}}|\partial_{k}u^{*}(x)|,$
    where $m_{+}^{*}=(m_{1},\ldots,m_{k}^{*}+1,\ldots,m_{p})^{T}$, the first inequality follows from the fact that $|A_{k}(m)|\leq 1$, and the last inequality is due to Taylor’s theorem.
  3. (iii)
    If there exists $m_{k}^{*}\in\{1,\ldots,N_{k}\}$ such that $\frac{m_{k}^{*}}{N_{k}}-\frac{2}{3N_{k}}\leq x_{k}\leq\frac{m_{k}^{*}}{N_{k}}-\frac{1}{3N_{k}}$, then by a same argument, we have
    $\partial_{k}u(x)$ $\leq 3N_{k}\Big{|}u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}-u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{k}^{*}-1}{N_{k}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}\Big{|}$
    $\leq 3\mathop{\operatorname{ess}\,\sup}_{x\in[0,1]^{p}}|\partial_{k}u^{*}(x)|.$
Combining the three cases above, we obtain the following inequality
$\|\partial_{k}u\|_{L^{\infty}([0,1]^{p})}\leq 3\|\partial_{k}u^{*}\|_{L^{\infty}([0,1]^{p})},\quad 1\leq k\leq p.$
This completes the proof. ∎
We have investigated the approximation error of a target function on the hypercube $[0,1]^{p}$ in Lemma H.5. In the following corollary, we extend our analysis to target functions on general bounded domain $[0,T]\times[-R,R]^{d}$.
Corollary H.6.
Let $p\in{\mathbb{N}}_{+}$, ${\mathcal{X}}=\prod_{k=1}^{p}[a_{i},b_{i}]$, and let $\{N_{k}\}_{k=1}^{p}$ be a set of positive integer. Set the deep neural network class $N(L,S)$ as $L=\lceil\log_{2}p\rceil+3$ and $S=(22p+6)\prod_{k=1}^{p}(N_{k}+1)$. Then for each $u^{*}\in W^{1,\infty}({\mathcal{X}})$, there exists a deep neural network $u\in N(L,S)$ such that
$\|u-u^{*}\|_{L^{\infty}({\mathcal{X}})}\leq 2^{p}\sum_{k=1}^{p}\frac{b_{k}-a_{k}}{N_{k}}\|\partial_{k}u^{*}\|_{L^{\infty}({\mathcal{X}})}.$
Further, it holds that for each $1\leq k\leq p$:
$\|u\|_{L^{\infty}({\mathcal{X}})}=\|u^{*}\|_{L^{\infty}({\mathcal{X}})}\quad\text{and}\quad\|\partial_{k}u\|_{L^{\infty}({\mathcal{X}})}\leq 3\|\partial_{k}u^{*}\|_{L^{\infty}({\mathcal{X}})}.$
Proof of Corollary H.6.
We first define a variable transformation on $u^{*}\in W^{1,\infty}({\mathcal{X}})$ as
$\phi:W^{1,\infty}({\mathcal{X}})$ $\rightarrow W^{1,\infty}([0,1]^{p})$
$u^{*}(x)$ $\mapsto(\phi\circ u^{*})(x^{\prime})=u^{*}(a+(b-a)x^{\prime}),$
where $(b-a)x^{\prime}=((b_{k}-a_{k})x_{k}^{\prime})_{k=1}^{p}\in{\mathbb{R}}^{p}$. Then it is noticeable that
$\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big{|}\frac{\partial(\phi\circ u^{*})}{\partial x_{k}^{\prime}}(x^{\prime})\Big{|}=(b_{k}-a_{k})\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big{|}\frac{\partial u^{*}}{\partial x_{k}}(x)\Big{|}.$
Set the deep neural network class $N(L,S)$ as $L=\lceil\log_{2}p\rceil+3$ and $S=(22p+6)\prod_{k=1}^{p}(N_{k}+1)$. According to Lemma H.5, there exists a neural network $u\in N(L,S)$ such that
$|u(x^{\prime})-(\phi\circ u^{*})(x^{\prime})|$ $\leq 2^{p}\sum_{k=1}^{p}\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big{|}\frac{\partial(\phi\circ u^{*})}{\partial x_{k}^{\prime}}(x^{\prime})\Big{|}\frac{1}{N_{k}}$
(H.8) $\leq 2^{p}\sum_{k=1}^{p}\frac{b_{k}-a_{k}}{N_{k}}\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big{|}\frac{\partial u^{*}}{\partial x_{k}}(x)\Big{|},$
and for each $1\leq k\leq p$, the following inequality holds:
(H.9) $\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big{|}\frac{\partial u}{\partial x_{k}^{\prime}}(x^{\prime})\Big{|}\leq 3\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big{|}\frac{\partial(\phi\circ u^{*})}{\partial x_{k}^{\prime}}(x^{\prime})\Big{|}=3(b_{k}-a_{k})\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big{|}\frac{\partial u^{*}}{\partial x_{k}}(x)\Big{|}.$
We next define the inverse transform on $u\in W^{1,\infty}([0,1]^{p})$ as
$\psi:W^{1,\infty}([0,1]^{p})$ $\rightarrow W^{1,\infty}({\mathcal{X}})$
$u(x^{\prime})$ $\mapsto(\psi\circ u)(x)=u\Big{(}\frac{x^{\prime}-a}{b-a}\Big{)},$
where $(x^{\prime}-a)/(b-a)=((x_{k}^{\prime}-a_{k})/(b_{k}-a_{k}))_{k=1}^{p}\in{\mathbb{R}}^{p}$. It follows from (H.9) that
$\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big{|}\frac{\partial(\psi\circ u)}{\partial x_{k}}(x)\Big{|}=\frac{1}{b_{k}-a_{k}}\mathop{\operatorname{ess}\,\sup}_{x^{\prime}\in[0,1]^{p}}\Big{|}\frac{\partial u}{\partial x_{k}^{\prime}}(x^{\prime})\Big{|}\leq 3\mathop{\operatorname{ess}\,\sup}_{x\in{\mathcal{X}}}\Big{|}\frac{\partial u^{*}}{\partial x_{k}}(x)\Big{|},$
for each $1\leq k\leq p$. Then composing $\psi$ on both sides of (H.8) yields the desired inequality. ∎
Appendix I Denoiser parameterization
In practice, we parameterize the network $D_{\theta}(t,x)$ following Karras et al. [2022]:
(I.1) $D_{\theta}(t,x)=c_{\text{skip}}(t)x+c_{\text{out}}(t)F_{\theta}\left(c_{\text{noise}}(t),c_{\text{in}}(t)x\right),$
where $F_{\theta}$ is the neural network to be trained, $c_{\text{skip}}(t)$ scale the skip connection, $c_{\text{in}}(t)$ and $c_{\text{out}}(t)$ scale the input and output of $F_{\theta}$, and $c_{\text{noise}}(t)$ scales time $t$.
Now (4.1) becomes
(I.2) ${\mathcal{L}}(F)=$ $\int_{0}^{1}{\mathbb{E}}_{X_{0}}{\mathbb{E}}_{X_{1}}\left[\omega(t)\|c_{\text{skip}}(t)X_{t}+c_{\text{out}}(t)F\left(c_{\text{noise}}(t),c_{\text{in}}(t)X_{t}\right)-x_{1}\|^{2}\right]\,\mathrm{d}t$
$=$ $\int_{0}^{1}{\mathbb{E}}_{X_{0}}{\mathbb{E}}_{X_{1}}\left[\omega(t)c_{\text{out}}^{2}(t)\|F\left(c_{\text{noise}}(t),c_{\text{in}}(t)X_{t}\right)-\frac{X_{1}-c_{\text{skip}}(t)X_{t}}{c_{\text{out}}(t)}\|^{2}\right]\,\mathrm{d}t$
$=$ $\int_{0}^{1}{\mathbb{E}}_{X_{0}}{\mathbb{E}}_{z}\left[\lambda(t)\|F_{\text{pred}}-F_{\text{target}}\|^{2}\right]\,\mathrm{d}t,$
where $\lambda(t)=\omega(t)c_{\text{out}}^{2}(t),F_{\text{pred}}=F\left(c_{\text{noise}}(t),c_{\text{in}}(t)X_{t}\right)$ and $F_{\text{target}}=\frac{X_{1}-c_{\text{skip}}(t)X_{t}}{c_{\text{out}}(t)}$.
Let $\sigma$ denote the standard deviation of $\mu_{1}$. We now design $c_{\text{in}}$ so that the spatial inputs of $F$ has unit variance.
$\operatorname{Var}[c_{\text{in}}(t)x_{t}]=1,$
$\Leftrightarrow$ $c_{\text{in}}(t)=\sqrt{\frac{1}{\operatorname{Var}[X_{t}]}}=\sqrt{\frac{1}{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}}.$
We then design $c_{\text{out}}$ so that the $F_{\text{target}}$ has unit variance.
$\operatorname{Var}[\frac{X_{1}-c_{\text{skip}}(t)X_{t}}{c_{\text{out}}(t)}]=1,$
$\Leftrightarrow$ $c_{\text{out}}(t)=\sqrt{\operatorname{Var}[(1-\alpha_{t}c_{\text{skip}}(t))X_{1}-c_{\text{skip}}(t)\beta_{t}z]},$
(I.3) $\Leftrightarrow$ $c_{\text{out}}(t)=\sqrt{(1-\alpha_{t}c_{\text{skip}}(t))^{2}\sigma^{2}+c_{\text{skip}}(t)^{2}\beta_{t}^{2}}.$
We then design $c_{\text{skip}}$ that minimizes $c_{\text{out}}$ so that the errors of $F$ are amplified as little as possible. Let $\frac{\partial c_{\text{out}}^{2}(t)}{\partial c_{\text{skip}}(t)}=0$, we obtain
$-\alpha_{t}(1-\alpha_{t}c_{\text{skip}}(t))\sigma^{2}+\beta_{t}^{2}c_{\text{skip}}(t)=0,$
(I.4) $\Leftrightarrow$ $c_{\text{skip}}(t)=\frac{\alpha_{t}\sigma^{2}}{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}.$
We can check that (I.4) is indeed the minima of $c_{\text{out}}$. Meanwhile, (I.3) yields
$c_{\text{out}}(t)=\frac{\beta_{t}\sigma}{\sqrt{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}}.$
We can then design $\omega(t)$ so that $\lambda(t)=1$ uniformly on $[0,1]$.
$\omega(t)=\frac{1}{c_{\text{out}}^{2}(t)}=\frac{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}{\beta_{t}^{2}\sigma^{2}}.$
We conclude the form of coefficients in Table 5.
Table 5. Denoiser parameterization.
function requirements form
$c_{\text{noise}}(t)$ free choice
$c_{\text{in}}(t)$ $\operatorname{Var}[c_{\text{in}}(t)X_{t}]=1$ $\sqrt{\frac{1}{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}}$
$c_{\text{out}}(t)$ $\operatorname{Var}[\frac{X_{1}-c_{\text{skip}}(t)X_{t}}{c_{\text{out}}(t)}]=1$ $\frac{\beta_{t}\sigma}{\sqrt{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}}$
$c_{\text{skip}}(t)$ $\frac{\partial c_{\text{out}}^{2}(t)}{\partial c_{\text{skip}}(t)}=0$ $\frac{\alpha_{t}\sigma^{2}}{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}$
$\omega(t)$ $\lambda(t)=1$ $\frac{\alpha_{t}^{2}\sigma^{2}+\beta_{t}^{2}}{\beta_{t}^{2}\sigma^{2}}$
Now (I.2) can be used as the working denoiser matching loss.
Appendix J Extra experiment details
We use DDPM++ in as the denoiser network architecture, and embed another temporal input into it to construct the characteristic generator as Kim et al. [2024] did. We use Rectified Adam (RAdam) as the optimizer and set learning rate to 0.001 on MNIST and 0.0001 on CIFAR-10. In practice, we initialize $D_{{\mathcal{S}}}$ from the pre-trained $D_{{\mathcal{T}}}$ for fast convergences. Inspired by Esser et al. [2021], Kim et al. [2024], we adaptively balance the the characteristic matching loss and the denoiser matching loss. For characteristic matching on image data, we use the Learned Perceptual Image Patch Similarity (LPIPS) metric instead of $L^{2}$-norm to measure distance [Zhang et al., 2018]. We use exponential moving average (EMA) with rate 0.999 on MNIST and 0.9999 on CIFAR-10. For the choice of $u$ in (4.8), we apply a weighted random strategy which is more likely to choose a long range interval $[t,u]$ to ensure the precision of the teacher solver, as Kim et al. [2024] did, For each batch, we run the teacher solver with at most 19 NFE. On MNIST, we calculate FID in original pixel space. On CIFAR-10, we use Inceptionv3 to extract features (dimension=2048) of images first and then calculate FID in the feature space.